### v1

# Deviation Inequalities for Rényi Divergence Estimators via Variational Expression

Abstract.
Rényi divergences play a pivotal role in information theory, statistics, and machine learning. While several estimators of these divergences have been proposed in the literature with their consistency properties established and minimax convergence rates quantified, existing accounts of probabilistic bounds governing the estimation error are premature. Here, we make progress in this regard by establishing exponential deviation inequalities for smoothed plug-in estimators and neural estimators by relating the error to an appropriate empirical process and leveraging tools from empirical process theory. In particular, our approach does not require the underlying distributions to be compactly supported or have densities bounded away from zero, an assumption prevalent in existing results. The deviation inequality also leads to a one-sided concentration bound from the expectation, which is useful in random-coding arguments over continuous alphabets in information theory with potential applications to physical-layer security. As another concrete application, we consider a hypothesis testing framework for auditing Rényi differential privacy using the neural estimator as a test statistic and obtain non-asymptotic performance guarantees for such a test.
Key words and phrases:
Rényi divergence, deviation inequality, variational estimator, smoothed plug-in estimator, differential privacy
1. Introduction
Rényi divergences play a fundamental role in information theory, statistics, machine learning, and related fields. These divergences as well as information measures derived from them, such as mutual information and entropy[54], characterize fundamental limits in information-theoretic coding problems. For instance, they characterize the error-exponents in hypothesis testing, the capacity of noisy channels, ultimate limits in data compression and rate-distortion theory, to name a few (see e.g., [11, 16]). Further, they appear as discrepancy measures or as optimization metrics in applications. Rényi divergence is a functional of the underlying probability distributions, which are often unknown in practice. In such scenarios, the problem of estimating Rényi divergences from samples arises, for which various estimators have been proposed in the literature, such as those based on plug-in, kernel, and variational methods (see Section 1.3 below). While consistency and minimax error rates for several of these estimators have been established, an understanding of the probability law governing the estimation error, which is pivotal for a complete statistical account of performance, is limited.
In this work, we make progress in this regard by establishing exponential deviation inequalities for two classes of Rényi divergence estimators, namely Gaussian-smoothed plug-in estimator and neural estimator. To this end, we utilize known variational expressions for these divergences (see (1) and (2) below)[19, 50, 10]. Our approach relies on relating the deviation to the suprema of empirical processes indexed by suitable function classes, and using concentration inequalities for the suprema to arrive at the desired conclusion. Our results go beyond the framework of discrete alphabets and compactly supported distributions with densities bounded away from zero, as required in existing works (see e.g., [58, 59, 65]). Specifically, we obtain deviation inequalities for Gaussian-smoothed empirical measures (based on samples) which satisfy neither of these criteria, under the assumption that the unsmoothed distribution is compactly supported or sub-Gaussian. Further, under an additional assumption that the densities have bounded ratios and belong to an appropriate Lipschitz class, we are also able to treat the deviation of the estimate from the unsmoothed value. We also obtain exponential inequalities for neural estimators of Rényi divergences and discuss its application to obtaining the first complete non-asymptotic performance guarantee for auditing Rényi differential privacy (DP) [33, 18] using the neural estimator as a test statistic.
1.1. Estimation of Rényi divergences
For $\alpha\in(0,1)\cup(1,\infty)$, the Rényi divergence of order $\alpha$ between Borel probability measures $\mu$ and $\nu$ on $\mathbb{R}^{d}$ is
$\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right):=\begin{cases}\frac{1}{\alpha-1}\log\big{(}\int_{\mathbb{R}^{d}}p_{\mu}^{\alpha}p_{\nu}^{1-\alpha}d\rho\big{)},&\mu\ll\nu\mbox{ or }\mu\not\perp\nu\mbox{ for }\alpha\in(0,1),\\ \infty,&\mbox{otherwise},\end{cases}$
where $\rho$ is a sigma-finite positive measure such that $\mu,\nu\ll\rho$ and $\mu\not\perp\nu$ denotes that $\mu$ and $\nu$ are non-singular. The special cases of Rényi divergences of order $\alpha\in\{0,1,\infty\}$ are defined as the respective limits of the above expression. The case $\alpha=1$ corresponds to the Kullback-Leibler (KL) divergence and $\alpha=\infty$ is the max-divergence, given by
$\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right):=\begin{cases}\int_{\mathbb{R}^{d}}p_{\mu}\log\frac{p_{\mu}}{p_{\nu}}d\rho,&\mu\ll\nu,\\ \infty,&\mbox{otherwise},\end{cases}$
and
$\mathsf{D}_{\max}\left(\mu\middle\|\nu\right):=\begin{cases}\log\left\|\frac{p_{\mu}}{p_{\nu}}\right\|_{\infty,\rho}&\mu\ll\nu,\\ \infty,&\mbox{otherwise},\end{cases}$
[…]
Rényi divergences can be expressed in a variational form as the supremum over a function class of a functional that involves expectations with respect to (w.r.t.) the underlying distributions. The following variational (or dual) expressions are well-known [^1] The expression in (1a) is known as the Donsker-Varadhan variational expression. The second expression given in (1a) which will be useful for our purposes can be shown via an application of the Legendre-Fenchel duality applied to the convex map $x\mapsto x\log x$ (see [50]). for KL divergence [19, 50]:
$\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right)$ $=\sup_{f}\mathbb{E}_{\mu}[f]-\log\mathbb{E}_{\nu}[e^{f}]$ (1a)
$=\sup_{f}\mathbb{E}_{\mu}[f]-\mathbb{E}_{\nu}[e^{f}]+1,$ (1b)
where the suprema above are over all measurable functions such that the second expectations in (1a) and (1b) are finite [^2] Alternatively, one may consider the supremum to be over all measurable functions with the interpretation, $\infty-\infty=$-$\infty$ and $\infty+\infty=-\infty$, or even restrict the supremum to be over all bounded measurable functions without loss of generality.. The suprema in (1) are achieved when $\mu\ll\nu$ by $f_{\mu,\nu}^{\star}=\log\frac{d\mu}{d\nu}$. For Rényi divergences of order $\alpha\in(0,1)\cup(1,\infty)$, the following variational expression was shown in [10]:
$\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)=\sup_{f}\frac{\alpha}{\alpha-1}\log\left(\int_{\mathbb{R}^{d}}e^{(\alpha-1)f}d\mu\right)-\log\left(\int_{\mathbb{R}^{d}}e^{\alpha f}d\nu\right),$ (2)
where the supremum is over all measurable functions such that the second integral is finite. The supremum is achieved when $\mu\ll\nu$ and $\left(\frac{d\mu}{d\nu}\right)^{\alpha}\in L_{1}(\nu)$ by $f_{\mu,\nu}^{\star}$ [10]. It is known that the value of the expressions on the right hand side (RHS) of (1a), (1b), and (2) are unchanged if $f$ is restricted to the set of all bounded continuous functions (see e.g., [10]). An alternative expression similar in spirit to (1b) also holds for $\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)$ (see Lemma 7 in Appendix).
The aforementioned variational expressions have been leveraged recently to construct an estimator of Rényi divergences and $f$-divergences [7, 65, 10]. The estimate is obtained simply by approximating the expectations in the dual form by that w.r.t. the empirical measures based on samples and maximizing over an appropriate class of functions, e.g., the output of a neural net. Such estimators have excellent adaptability and scalability properties in practice, and have been shown to be consistent under mild assumptions on the underlying distributions. In fact, such estimators are minimax optimal achieving a parametric rate of convergence under sufficient smoothness of densities, provided the neural network is scaled appropriately according to sample size[65]. Additionally, the specific form of the variational formula as well as the resulting estimator lends itself naturally to an analysis via empirical process theory. Here, we utilize these aspects to derive exponential deviation inequalities for Gaussian-smoothed plug-in estimator and neural estimator via powerful concentration inequalities for empirical process suprema.
Our approach to deriving the deviation inequality for Gaussian smoothed plug-in estimator relies on identifying primitive conditions on the distributions which allow the estimation error to be controlled by the suprema of empirical processes with good concentration properties. In this regard, we show that when $\mu$ and $\nu$ are compactly supported or sub-Gaussian, it is possible to relate the desired deviation to suprema of empirical processes indexed by a smooth function class with an integrable envelope [^3] An envelope of a class $\mathcal{F}$ of functions on $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a function $F:\mathcal{X}\rightarrow[0,\infty]$ such that $\sup_{f\in\mathcal{F}}\left|f(x)\right|\leq F(x)$ for all $x\in\mathcal{X}$.(in a suitable Orlicz sense). The integrability of the envelope then translates to a concentration property for the suprema as quantified by a Fuk-Nagaev type inequality[24, 2]. In order to apply the latter, we require an upper bound on the covering entropy (see Definition 3) of a Gaussian-smoothed function class with an envelope $F(x)$ that scales at most as some power of $\left\|x\right\|$. As for neural estimators, we leverage a version of Talagrand’s concentration inequality given in [12] to obtain the desired exponential inequality.
1.2. Contributions
Below, we summarize our main contributions. Let $\hat{\mu}_{n}$ and $\hat{\nu}_{n}$ denote the standard empirical probability measures constructed from $n$ independent and identically distributed (i.i.d.) samples generated according to $\mu$ and $\nu$, respectively. Let * denote the convolution operation and $\gamma_{\sigma}$ be the centered Gaussian measure (or distribution) on $\mathbb{R}^{d}$ with covariance matrix $\sigma^{2}I_{d}$. Then, we show the following [^4] Here, $g=O_{z}(f)$ and $g=\Omega_{z}(f)$, respectively, signifies that $g\leq c_{z}f$ and $g\geq c_{z}f$ for some constant $c_{z}>0$ that depends on $z$. deviation inequalities including explicitly bounding the constant pre-factors appearing in these expressions up to universal constants:
  1. (i)
    For all $\mu,\nu$ whose support is contained in a $d$-dimensional ball of radius $r$, we prove in Theorem 1 that
[…]
    Combining this with an approximation result (Proposition 1) for Rényi divergences, we obtain a deviation inequality (see Corollary 1) for $\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right)\right|$, under a boundedness and Lipschitz continuity assumption on the densities of $\mu,\nu$. An extension to the case of sub-Gaussian $\mu$ and $\nu$ in the one-sample case, i.e. $\hat{\nu}_{n}=\nu$ for all $n$, is also obtained.
  2. (ii)
    For Rényi divergences of order $\alpha\in(0,1)\cup(1,\infty)$, we establish in Theorem 3 that for $\mu,\nu$ as above
    $\mathbb{P}\left(\left|\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\geq\Omega_{\alpha,r,d,\sigma}\big{(}n^{-\frac{1}{2}}+z\big{)}\right)\leq 2e^{-\left(nz^{2}\wedge nz\right)}.$ (4)
  3. (iii)
    For neural estimators of Rényi divergences with a suitably chosen bounded neural net class $\mathcal{G}$, which can approximate $f^{\star}_{\mu,\nu}$ up to an error $\delta$ uniformly on the support of $\rho$ and whose uniform $L^{2}$-covering entropy numbers satisfies an integrability condition, we show in Theorem 4 that
    $\mathbb{P}\left(\left|\hat{\mathsf{D}}_{\alpha,\mathcal{G}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)\right|\geq\big{(}M^{2\alpha}+M^{2\left|\alpha-1\right|}\big{)}\delta+\Omega_{d,M,\alpha}(n^{-\frac{1}{2}}+z)\right)\leq 2e^{-(nz^{2}\wedge nz)}.$ (5)
    Here, $M$ is a parameter which upper bounds the ratio of densities of $\mu$ and $\nu$.
Exponential concentration inequalities combined with random coding- or expurgation-based arguments are useful in information-theoretic security to show existence of coding schemes satisfying secrecy criteria quantified in terms of a divergence measure (see e.g., [44, 25, 63, 64]). The deviation inequalities (3)-(4) are continuous alphabet analogues of such inequalities for Rényi divergences, which, to the best of our knowledge, are known currently only for finite alphabets. Hence, we expect that these inequalities will have potential applications to secure and covert communications over a Gaussian wiretap channel[43]. As a concrete application of (5), we consider a hypothesis testing based framework for auditing Rényi DP and obtain non-asymptotic performance guarantees for this test. While this application has been considered before in [18] and [67], the auditing guarantees were either partial or only valid asymptotically. Here, we address these shortcomings thus making the auditing framework more appealing in practice.
1.3. Related work
The literature on estimation of information and divergence measures is extensive (see e.g., [77] for a survey). Several estimators of such measures have been proposed and their statistical properties analyzed. In this regard, consistency, convergence rates, and minimax rates of various estimators based on plug-in, kernel, k-nearest neighbour, and transform methods have been studied in [39, 72, 4, 52, 40, 78, 15, 42, 53, 50, 73, 69, 36, 34, 60, 80, 74, 51, 49, 13, 14, 9, 8, 28, 1, 26, 31, 46, 30, 82, 56, 70, 71, 32] (see also references therein). Further statistical properties such as the asymptotic distribution of the estimation error in the flavor of the classical central limit theorem have also been characterized, see e.g. [57, 36, 48, 38, 27, 66, 61, 67, 62]. Consistency of neural estimators of mutual information and Rényi divergences have been explored in [7] and [10], respectively, while minimax convergence rate for neural estimation of $f$-divergences over a sufficiently regular class of distributions has been established in [68, 65].
Probabilistic tail inequalities for kernel-based estimators of Rényi divergences and neural estimators of KL divergence for compactly supported distributions have been derived in [58, 59, 65]. For plug-in estimators of KL divergence, Chernoff-type concentration bounds have been studied in [29] for the finite alphabet setting, improving upon earlier works [3, 45]. However, all these results are limited to distributions with compact support and smooth densities bounded away from zero, which we aim to generalize upon.
2. Preliminaries
2.1. Notation
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a sufficiently rich probability space on which all random variables are defined. Let $\rho$ be a sigma-finite positive measure on the Borel sigma algebra $\mathcal{B}(\mathbb{R}^{d})$ and $\mathcal{P}(\mathbb{R}^{d})$ be the set of Borel probability measures. For $\mu\in\mathcal{P}(\mathbb{R}^{d})$ such that $\mu\ll\rho$, we use $p_{\mu}=\frac{d\mu}{d\rho}$ to denote its density (Radon-Nikodym derivative), w.r.t. $\rho$. $\mu^{\otimes n}$ stands for the $n$-fold product measure, and $\delta_{x}$ represents the Dirac measure at $x$. $\mathbbm{1}_{\mathcal{E}}$ denotes the indicator of set $\mathcal{E}$. $\mu*\nu$ denotes the convolution of $\mu$ and $\nu$; likewise, fgf*g represents convolution of two measurable functions $f,g$. $X\sim\mu$ stands for random variable $X$ distributed according to $\mu$. The expectation of $X\sim\mu$, whenever it exists, will be denoted by $\mathbb{E}_{\mu}[X]$. We write $\gamma_{\sigma}=N(0,\sigma^{2}I_{d})$ for the centered Gaussian distribution on $\mathbb{R}^{d}$ with covariance matrix $\sigma^{2}I_{d}$, and use $\varphi_{\sigma}(x)=(2\pi\sigma^{2})^{-d/2}e^{-\left\|x\right\|^{2}/(2\sigma^{2})}$ for the corresponding density. Inequalities between random variables are interpreted to hold almost surely or surely, which will be evident from the context.
For $1\leq r\leq\infty$, let $L^{r}(\rho)$ be the space of all real-valued measurable functions $f$ on $\mathbb{R}^{d}$ such that $\left\|f\right\|_{r,\rho}:=\big{(}\int_{\mathfrak{S}}\left|f\right|^{r}d\rho\big{)}^{1/r}<\infty$, with the usual identification of functions that are equal $\rho$-almost everywhere (a.e.). For $1\leq r<\infty$, the space $(L^{r},\left\|\cdot\right\|_{r,\rho})$ is a separable Banach (and hence, Polish) space. When $\rho$ is the Lebesgue measure on $\mathbb{R}^{d}$, we use $\|\cdot\|_{r}$ and $L^{r}(\mathbb{R}^{d})$ to denote the relevant norm and space of functions. The corresponding restrictions to a measurable subset $\mathcal{X}\subseteq\mathbb{R}^{d}$ are denoted by $\|\cdot\|_{r,\mathcal{X}}$ and $L^{r}(\mathcal{X})$, respectively. The support of $\rho$ is denoted by $\mathrm{spt}(\rho)$. We use $a\lesssim_{x}b$ to denote that $a\leq c_{x}b$ for some constant $c_{x}>0$ which depends only on some parameter $x$, and omit the subscript $x$ for universal constants. For $x,y\in\mathbb{R}$, $x\wedge y=\min\{x,y\}$ and $x\vee y=\max\{x,y\}$.
2.2. Definitions
The following function classes will play a role in our results.
Definition 1 (Hölder class, see e.g. [75]).
For a $d$-dimensional vector $k$ with non-negative integers as elements, consider the differential operator
$D^{k}=\frac{\partial^{|k|}}{\partial x_{1}^{k_{1}}\ldots\partial x_{d}^{k_{d}}}.$
For a function $f:\mathcal{X}\rightarrow\mathbb{R}$ defined on a set $\mathcal{X}\subseteq\mathbb{R}^{d}$ with nonempty interior, the Hölder norm of order $\beta>0$ is
$\left\|f\right\|_{\beta,\mathcal{X}}:=\max_{0\leq|k|\leq\underaccent{\bar}{\beta}}\sup_{x}\left|D^{k}f(x)\right|+\max_{|k|=\underaccent{\bar}{\beta}}\sup_{x,y}\frac{\left|D^{k}f(x)-D^{k}f(y)\right|}{\left\|x-y\right\|^{\beta-\underaccent{\bar}{\beta}}},$ (6)
where $\underaccent{\bar}{\beta}$ is the largest integer strictly smaller than $\beta$, and the supremums are taken on the interior of $\mathcal{X}$ with $x\neq y$. The Hölder class $C_{M}^{\beta}(\mathcal{X})$ of smoothness parameter $\beta$ and norm parameter $M$ is the set of all continuous functions $f:\mathcal{X}\rightarrow\mathbb{R}$ such that $\left\|f\right\|_{\beta,\mathcal{X}}\leq M$.
Definition 2 (Lipschitz class, see [17]).
For $r\in(0,\infty]$, $m\in\mathbb{N}$, and $f\in L^{r}\big{(}\rho\big{)}$, the $m$-th modulus of smoothness of $f$ is
$\chi_{m,r}(f,t,\rho):=\sup_{y\in\mathbb{R}^{d},\left\|y\right\|\leq t}\left\|\Delta_{y}^{m}f\right\|_{r,\rho},$
where $\Delta_{y}^{m}f(x)=\sum_{j=0}^{m}(-1)^{m-j}f(x+jy)$. For $0<s\leq 1$ and $\mathcal{X}\subseteq\mathbb{R}^{d}$, the Lipschitz class with smoothness parameter $s$ and norm parameter $M$ is $\mathsf{Lip}_{s,r,M}(\rho):=\big{\{}f\in L^{r}\big{(}\rho\big{)}:\left\|f\right\|_{\mathrm{Lip}(s,r,\rho)}\leq M\big{\}}$ where
$\left\|f\right\|_{\mathrm{Lip}(s,r,\rho)}:=\inf_{\begin{subarray}{c}g:\\ g=f,\rho\text{ a.e.}\end{subarray}}\left\|g\right\|_{r}+\sup_{t>0}t^{-s}\chi_{1,r}(g,t)$
is the Lipschitz seminorm.
The above definition of Lipschitz class w.r.t. a general $\rho$ is a slight generalization of the one given in [17] w.r.t. the Lebesgue measure. This is required since we also want to treat discrete probability measures on $\mathbb{R}^{d}$. It is known that $\mathsf{Lip}_{s,r,M}(\rho)$ grows as the smoothness parameter $s$ decreases. Also, the class of functions with bounded variation (for $d=1$) is contained in $\mathsf{Lip}_{1,r,M}(\rho)$ for some appropriately chosen $M$ and $\rho$ being the Lebesgue measure. Hence, the Lipschitz class of smoothness $0<s\leq 1$ contains most probability densities of practical interest.
We will also require the notion of covering entropy as stated next.
Definition 3 (Covering entropy, see e.g.[75]).
Let $(\Theta,\mathsf{d})$ be a metric space. A set $\Theta^{\prime}\subseteq\Theta$ is an $\epsilon$-covering of $(\Theta,\mathsf{d})$ if for every $\theta\in\Theta$, there exists $\tilde{\theta}\in\Theta^{\prime}$ such that $\mathsf{d}(\theta,\tilde{\theta})\leq\epsilon$. The $\epsilon$-covering number is
$N(\epsilon,\Theta,\mathsf{d}):=\inf\left\{|\Theta^{\prime}|:\,\Theta^{\prime}\mbox{ is an }\epsilon\mbox{-covering of }\Theta\right\},$
and the $\epsilon$-covering entropy is $\log N(\epsilon,\Theta,\mathsf{d})$.
The Orlicz norm on the space of random variables that specify their tail behaviour will play a role in our proofs.
Definition 4 (Orlicz space, see e.g., [75]).
An increasing convex function $\psi:[0,\infty)\rightarrow[0,\infty)$ with $\psi(0)=0$ and $\lim_{x\rightarrow\infty}\psi(x)=\infty$ is called an Orlicz function. The associated Orlicz norm of a random variable $X$ taking values in a Banach space $\mathfrak{B}$ is
$\left\|X\right\|_{\psi}=\inf\left\{c>0:\mathbb{E}\left[\psi\left(\frac{\left\|X\right\|_{\mathfrak{B}}}{c}\right)\right]\leq 1\right\}.$ (7)
Examples of Orlicz functions include $\hat{\psi}_{q}(z)=z^{q}$ and $\psi_{q}(z)=e^{z^{q}}-1$, $z\in\mathbb{R}$, for $q\geq 1$; in particular, $\psi_{2}$ corresponds to the sub-Gaussian class defined next.
Definition 5 (Sub-Gaussian distribution, see e.g., [35]).
A distribution $\mu\in\mathcal{P}\big{(}\mathbb{R}^{d}\big{)}$ is $\sigma^{2}$-sub-Gaussian for $\sigma>0$ if $X\sim\mu$ satisfies
$\mathbb{E}\left[e^{u\cdot(X-\mathbb{E}[X])}\right]\leq e^{\frac{\sigma^{2}\left\|u\right\|^{2}}{2}},\quad\forall~u\in\mathbb{R}^{d}.$
For $M\geq 0$, let $\mathsf{SG}(M)$ be the set of all $\sigma^{2}$-sub-Gaussian distributions with $\sigma^{2}\vee\left\|\mathbb{E}[X]\right\|\leq M$.
We will also consider $\psi_{q}$ for $0<q<1$. In this case, $\left\|X\right\|_{\psi_{q}}$ as given in (7), only defines a quasi-norm on the space of random variables as opposed to a norm (since $\psi_{q}$ is not convex). However, we will allow this abuse of notation.
We also require the following technical assumption on certain function classes of interest.
Assumption 1 (Pointwise separability and existence of envelope).
A class of measurable functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is pointwise separable if it is separable w.r.t. the topology of pointwise convergence on $\mathbb{R}^{d}$. Further, a class $\mathcal{F}$ has a measurable envelope if there exists a measurable function $F:\mathbb{R}^{d}\rightarrow[0,\infty)$ such that $\sup_{f\in\mathcal{F}}\left|f(x)\right|\leq F(x)$ for all $x\in\mathbb{R}^{d}$. Here, $\sup_{f\in\mathcal{F}}\left|f(x)\right|$ is the minimal envelope, but other choices can also be considered. Henceforth, whenever we consider an envelope, measurability will be implicitly assumed.
Assumption 1 effectively allows empirical process suprema over $\mathcal{F}$ to be replaced by those over a countable sub-class of functions and avoids measurability issues that could arise when dealing with such quantities.
3. Main Results
For $X^{n}\sim\mu^{\otimes n}$ and $Y^{n}\sim\nu^{\otimes n}$ where $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$, let
$\hat{\mu}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}\quad\text{and}\quad\hat{\nu}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{Y_{i}}$
denote the empirical distributions of $X^{n}$ and $Y^{n}$, respectively. Let $\mathbb{B}_{d}(r):=\{x\in\mathbb{R}^{d}:\left\|x\right\|_{2}\leq r\}$ denote the Euclidean ball of radius $r$ and $\mathcal{P}\big{(}\mathbb{B}_{d}(r)\big{)}:=\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\mathrm{spt}(\mu)\subseteq\mathbb{B}_{d}(r)\}$ be the set of Borel probability measures whose support is contained in $\mathbb{B}_{d}(r)$.
3.1. Deviation Inequalities for Smoothed Plug-in Estimators
We first state the deviation inequality for Gaussian-smoothed plug-in estimator of KL divergence.
Theorem 1 (Compactly supported and sub-Gaussian distributions).
Let $0<\sigma\leq 1$, $r\geq 0$, $p\geq 2$, $L\geq 1$ and $0\leq\tau\leq 1$. Then, the following hold:
  1. (i)
[…]
    where $\xi_{r,d,\sigma}$ and $\tilde{\xi}_{r,d,\sigma}$ are defined in (35) and (36), respectively.
  2. (ii)
    $\sup_{\begin{subarray}{c}\mu,\nu\in\mathsf{SG}(L):\\ X\sim\mu,~\left\|X\right\|_{\psi_{p}}\leq L\end{subarray}}\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\gtrsim\check{\xi}_{d,L,\sigma}\left(\hat{\xi}_{d,L,\sigma}n^{-\frac{1}{2}(1-\tau)}+z\right)\right)$
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq 6e^{-\hat{\zeta}_{n,d,p,\tau}(z)},$ (9)
    where $\hat{\xi}_{d,L,\sigma}$, $\check{\xi}_{d,L,\sigma}$ and $\hat{\zeta}_{n,d,p,\tau}(z)$ are given in (42b), (42c) and (43), respectively.
The proof of Theorem 1 is given in Section 4.1 and relies on a more general result (Lemma 2) stated below. At a high level, the proof relies on relating the desired deviation to suprema of empirical processes indexed by appropriate function classes and leveraging powerful concentration inequalities from empirical process theory. More specifically, (8) utilizes Talagrand’s concentration inequality for bounded function classes given in [12, Theorem 2.3] (see Theorem 6 below), while (9) relies on a Fuk-Nagaev type inequality given in [2, Theorem 4] (see Theorem 7 below) that applies to unbounded function classes. To apply these results, we also obtain an upper bound on the covering entropy of Gaussian-smoothed bounded function class (Lemma 3) by upper bounding the Hölder norm of such functions and utilizing a refinement of the covering entropy estimate for the Hölder class, which specifies the dependence on dimension and smoothness parameter in the prefactor.
Theorem 1 provides deviation inequalities under a more general setting than covered by existing results. Note that the usual assumptions of compact support and bounded (away from zero) densities do not apply here. These regularity assumptions on the densities are now replaced effectively by the condition that the log-likelihood of its ratio belongs to an appropriate Orlicz class. It can be noted from the proofs that the condition $\sigma\leq 1$ is not a fundamental restriction, but only imposed since it simplifies the analysis in addition to being also the regime of interest for smoothed estimators.
Remark 1 (Deviation inequality from expectation).
Since
$\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\leq\mathbb{E}\left[\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right],$
by convexity of KL divergence (see e.g., [76, Theorem 11]), Lemma 2 also yields a deviation inequality for the upper tail, i.e., (8) holds with $\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|$ replaced by $\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathbb{E}\left[\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right]$. Similar claim also holds for (9). Further, an inspection of the proof shows that (8) also holds when $\hat{\mu}_{n}*\gamma_{\sigma}$ or $\hat{\nu}_{n}*\gamma_{\sigma}$ (or both) are set to $\mu*\gamma_{\sigma}$ and $\nu*\gamma_{\sigma}$, respectively.
The proof of Theorem 1 is based on upper bounding expected empirical process suprema in terms of entropy integrals (see Lemma 2). While this approach results in easy to compute deviation inequalities, it is not always tight. A more precise deviation inequality can be obtained by replacing the expected suprema by its symmetrized version w.r.t. Rademacher variables, albeit by sacrificing computational simplicity. To state this result, let $\varepsilon:=\{\varepsilon_{i}\}_{i=1}^{n}$ denote i.i.d. Rademacher variables independent of $(X^{n},Y^{n})$, i.e., $\mathbb{P}(\varepsilon_{i}=1)=\mathbb{P}(\varepsilon_{i}=-1)=0.5$. Denote by $\mathbb{E}_{\varepsilon}[\cdot]$, the expectation w.r.t. $\varepsilon$. We have the following deviation inequality for compactly supported distributions (see Lemma 5 in Section 4.2 for a more general statement).
Theorem 2 (Deviation inequality in terms of Rademacher expectation).
Let $r\geq 0$. Set
[…]
where $b_{r,\sigma}:=\frac{r^{2}+4r}{\sigma^{2}}$,
$\mathcal{F}_{\sigma}(b)$ $:=\{f*\varphi_{\sigma}:\left|f(x)\right|\leq 0.5b(1+\left\|x\right\|),~\forall~x\in\mathbb{R}^{d}\},$
$\mathcal{E}_{\sigma}(b)$ $:=\{e^{f}*\varphi_{\sigma}:\left|f(x)\right|\leq 0.5b(1+\left\|x\right\|),~\forall~x\in\mathbb{R}^{d}\}.$
Then
$\sup_{\begin{subarray}{c}\mu,\nu\in\mathcal{P}\left(\mathbb{B}_{d}(r)\right)\end{subarray}}\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\geq 7n^{-1}\big{(}Z^{(\varepsilon)}_{n,\sigma}+\bar{Z}^{(\varepsilon)}_{n,\sigma}\big{)}+44\xi_{r,d,\sigma}z\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\leq 2e^{-\big{(}nz^{2}\wedge nz\big{)}},$
where $\xi_{r,d,\sigma}$ is given in (35).
The proof of Theorem 2 is given in Section 4.2 and follows by replacing Talagrand’s inequality in the proof of Lemma 2 by a statistical version involving symmetrized Rademacher expectations given in [6, Theorem 2.1]. In fact, a more general claim can be proven as stated below, which could be of independent interest.
Lemma 1 (Statistical Fuk-Nagaev inequality).
[…]
$Z_{n}$ $:=\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\mu}[f]\big{)}\right|,$ (11a)
$\mbox{and }Z^{(\varepsilon)}_{n}$ $:=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|\right].$ (11b)
Then
[…]
where $c_{q}$ is as specified in (83). Further, (12) holds with $Z^{(\varepsilon)}_{n}$ replaced by
[…]
Lemma 1 can be considered as a generalization of [6, Theorem 2.1] to an unbounded function class whose envelope is integrable in a suitable Orlicz sense. The proof is given in Section 5.3. The essential idea is based on truncation of $\mathcal{F}_{n}$ to a bounded class of functions $\mathcal{F}_{n}(T_{n})=\left\{f\mathbbm{1}_{F_{n}\leq T_{n}}:f\in\mathcal{F}_{n}\right\}$, applying [6, Theorem 2.1] to the truncated class, invoking the Hoffmann-Jørgensen inequality (see e.g. [41, Proposition 6.8]) to control the remainder by selecting an appropriate $T_{n}$, and showing that $\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}_{n}(T_{n})}\left|\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|\right]\leq Z^{(\varepsilon)}_{n}$.
Remark 2 (Statistical Fuk-Nagaev with Lebesgue integrable envelope).
Under the weaker condition $F_{n}\in L^{p}(\mu)$ for some $p\geq 2$, it is possible to derive the inequality
[…]
This follows by applying Markov’s inequality instead of (91) in the proof of Lemma 1, which yields
$\mathbb{P}\left(Z_{n}^{\prime\prime}\geq\frac{z}{4}\right)\leq 4^{p}\frac{\mathbb{E}\big{[}Z_{n}^{\prime\prime p}\big{]}}{z^{p}}\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{2(16)^{p}\mathbb{E}\left[M_{n}^{p}\right]}{z^{p}},$
where $(a)$ is via an application of Hoffmann-Jørgensen inequality given in [41, Proposition 6.8]:
$\mathbb{E}\big{[}Z_{n}^{\prime\prime p}\big{]}\leq 2(4^{p})\mathbb{E}\left[\max_{1\leq i\leq n}F_{n}^{p}\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right]\leq 2(4^{p})\mathbb{E}\left[M_{n}^{p}\right].$
Remark 3 (Confidence interval with unknown distribution).
[…]
Next, we state the analogue of Theorem 1 and Theorem 2 for Rényi divergences of order $\alpha\in(0,1)\cup(1,\infty)$.
Theorem 3 (Compactly supported distributions).
Let $r\geq 0$ and $0<\sigma\leq 1$. Then
[…]
where $\tilde{\xi}_{r,d,\sigma}$ and $\lambda_{\alpha,d,r,\sigma}$ are defined in (36) and (48), respectively.
The proof of Theorem 3 (see Section 4.3) is similar to that of Theorem 1, and utilizes the variational expression given in (2) to upper bound the desired deviation in terms of empirical process suprema. Note that existing exponential concentration inequalities such as those given in [59] do not apply here since the underlying densities are neither compactly supported nor bounded away from zero.
Remark 4 (Deviation inequality from expectation).
By joint convexity of Rényi divergences in both its arguments for $\alpha\in[0,1]$ [76, Theorem 11], Theorem 3 also holds with $|\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)|$ replaced by $\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathbb{E}\left[\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right]$ for $\alpha\in(0,1]$. Similarly, for all $\alpha\in(0,1)\cup(1,\infty)$, $|\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)|$ can be replaced by $\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathbb{E}\left[\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right]$ due to convexity of Rényi divergences in the second argument for all orders [76, Theorem 12]. Further, (13) also applies when $\hat{\mu}_{n}*\gamma_{\sigma}$ or $\hat{\nu}_{n}*\gamma_{\sigma}$ (or both) are set to $\mu*\gamma_{\sigma}$ and $\nu*\gamma_{\sigma}$, respectively.
Thus far, we considered deviation inequalities for smoothed plug-in estimators of Rényi divergences. Under additional assumptions uniformly bounding the ratio of densities of $\mu$ and $\nu$, it is possible to obtain probabilistic inequalities bounding the deviation of the estimate from the unsmoothed value. For stating the result, we require the concept of a Lipschitz class of functions given in Definition 2. For $\rho$ such that $\mu,\nu\ll\rho$, set
$\mathcal{P}_{M,\rho}$ $:=\left\{(\mu,\nu):\mu\ll\gg\nu\ll\rho,\left\|\frac{p_{\mu}}{p_{\nu}}\right\|_{\infty,\rho}\vee\left\|\frac{p_{\nu}}{p_{\mu}}\right\|_{\infty,\rho}\leq M\right\},$
$\tilde{\mathcal{P}}_{M,r,s,\rho}$ $:=\left\{(\mu,\nu)\in\mathcal{P}\big{(}\mathbb{B}_{d}(r)\big{)}\cap\mathcal{P}_{M,\rho}:p_{\mu},p_{\nu}\in\mathsf{Lip}_{s,1,M}(\rho)\right\}.$
With this, we have the following corollary to Theorem 1.
Corollary 1 (Deviation from unsmoothed KL divergence).
Let $s\in(0,1]$, $r\geq 0$, and $M\geq 1$. Then, with $\sigma_{n}=n^{-\frac{1}{d+2s+4}}$, we have
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\\ \tilde{\mathcal{P}}_{M,r,s,\rho}\end{subarray}}\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right)\right|\gtrsim\left(\bar{c}_{d,r}+c_{d,s}M^{2}+(r^{2}+4r)(1+r\sqrt{d})z\right)n^{\frac{-s}{d+2s+4}}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\leq e^{-\left(n^{\frac{d}{d+2s+4}}z^{2}\wedge n^{\frac{d+s+2}{d+2s+4}}z\right)},$ (14)
where $\bar{c}_{d,r}:=(3ed)^{3(d+2)}\sqrt{d}r(r+1)^{\frac{d}{2}+3}$ and $c_{d,s}:=\int_{\mathbb{R}^{d}}\left\|z\right\|^{s}\varphi_{1}(z)dz$.
The proof of Corollary 1 (see Section 4.4) relies on the fact that the difference between smoothed Rényi divergence and its unsmoothed value can be upper bounded given the densities $p_{\mu}$ and $p_{\nu}$ belong to an appropriate Lipschitz class with bounded ratios. The corresponding approximation error is quantified in the next proposition which is a generalization of [67, Lemma 3].
Proposition 1 (Approximation of Rényi divergences).
Let $\sigma\geq 0$, $s\in(0,1]$ and $M\geq 1$. Suppose $(\mu,\nu)\in\mathcal{P}_{M,\rho}$ are such that $p_{\mu},p_{\nu}\in\mathsf{Lip}_{s,1,M}(\rho)$. Then
$\left|\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\leq c_{d,s}M\left(M+1+\log M\right)\sigma^{s},$ (15a)
$\left|\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\leq c_{d,s}\left(M^{2\alpha+1}+\frac{\alpha}{\left|\alpha-1\right|}M^{2(1\vee\alpha)}\right)\sigma^{s},~\alpha\in(0,1)\cup(1,\infty),$ (15b)
where $c_{d,s}:=\int_{\mathbb{R}^{d}}\left\|z\right\|^{s}\varphi_{1}(z)dz$.
The proof of Proposition 1 is given in Section 4.5 and relies on a multi-variate Taylor’s expansion of a certain functional of the densities. The Lipschitz and regularity assumptions on the densities enable to control the remainder term in this expansion.
Figure 1. Random coding over a Gaussian channel where each codeword $X_{i}$ of a random codebook $\{X_{i}\}_{i=1}^{n}$ of size $n$ is a vector of length $d$ generated i.i.d. according to $\mu$. A power constraint on each input codeword naturally translates to a bound on its Euclidean norm.
Before we end this subsection, we briefly mention that the deviation inequalities given in Theorem 1 and Theorem 3 could be useful in other contexts in information theory such as physical layer security or covert communications, where the secrecy metric of interest is often expressed in terms of Rényi divergences. Exponential deviation inequalities then provide high confidence bounds in random coding- or expurgation-based arguments to show existence of codebooks satisfying the security criteria approximately. For instance, consider the case of information-theoretically secure coding over a Gaussian wiretap channel with a per-codeword power constraint. In this context, $\{X_{i}\}_{i=1}^{n}$ would denote a random codebook of size $n$ with each codeword of length $d$ sampled independently according to $\mu$ and $\hat{\mu}_{n}*\gamma_{\sigma}$ would correspond to the output distribution of a discrete-time additive Gaussian noise channel of variance $\sigma^{2}$ with input uniformly sampled from the codebook. Then, it is of interest to show that $\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\leq\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)+o(1)$ with overwhelming probability, where $\nu$ is an auxiliary distribution of interest. To the best of our knowledge, existing deviation inequalities relevant to this setting are applicable only to bounded metrics or finite (output-) alphabet channels (see e.g., [44, Theorem 31] and [25, Lemma 4]). Hence, we expect Theorem 1 and 3 to have potential new applications in wiretap coding or covert communications over Gaussian channels.
3.2. Deviation Inequalities for Neural Estimators of Rényi Divergences
Let $\mathcal{G}$ be a class of real-valued functions formed by the output of a neural network, referred to as neural net class. The variational forms (1b) and (2) leads to neural estimators of KL and Rényi divergence of order $\alpha\in(0,1)\cup(1,\infty)$ given by
$\hat{\mathsf{D}}_{\mathsf{KL},\mathcal{G}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)$ $:=\sup_{g\in\mathcal{G}}\mathbb{E}_{\hat{\mu}_{n}}[g]-\mathbb{E}_{\hat{\nu}_{n}}[e^{g}]+1,$ (16a)
and
$\hat{\mathsf{D}}_{\alpha,\mathcal{G}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)$ $:=\sup_{g\in\mathcal{G}}\frac{\alpha}{\alpha-1}\log\mathbb{E}_{\hat{\mu}_{n}}\big{[}e^{(\alpha-1)g}\big{]}-\log\mathbb{E}_{\hat{\nu}_{n}}\big{[}e^{\alpha g}\big{]},$ (16b)
respectively. We will assume that $\mathcal{G}$ satisfies the following mild assumption.
Assumption 2 (Neural class).
$\mathcal{G}$ is a class of measurable functions on $\mathbb{R}^{d}$ which is pointwise separable (i.e., separable w.r.t. the topology of pointwise convergence on the support of $\rho$) such that $\left\|g\right\|_{\infty,\rho}\leq b$, $\forall~g\in\mathcal{G}$, and
$\sup_{\gamma}\log N(\epsilon,\mathcal{G},\left\|\cdot\right\|_{2,\gamma})$ $\leq\hat{c}_{d,\beta}(b\epsilon^{-1})^{\frac{d}{\beta}},$ (17)
for some $\hat{c}_{d,\beta}\geq 0$, where the supremum is taken over all finitely supported discrete probability measures $\gamma$ with $\mathrm{spt}(\gamma)\subseteq\mathrm{spt}(\rho)$.
It is easy to see that Assumption 2 is satisfied by neural nets with finite width, finite depth, bounded parameters and continuous activation when $\rho$ has compact support for an appropriately chosen $b$.
We next describe the class of probability measures of interest for neural estimators. Let
$\bar{\mathcal{P}}_{M,\rho}(\mathcal{G},\delta)=\left\{(\mu,\nu)\in\mathcal{P}_{M,\rho}:\exists~g\in\mathcal{G}\mbox{ s.t. }\left\|f_{\mu,\nu}^{\star}-g\right\|_{\infty,\rho}\leq\delta\right\}.$
In words, $\bar{\mathcal{P}}_{M,\rho}(\mathcal{G},\delta)$ is the subclass of $\mathcal{P}_{M,\rho}$ such that the common optimizer $f_{\mu,\nu}^{\star}$ of the variational form in (1) and (2) can be approximated $\rho$-a.e. by some function in $\mathcal{G}$ up to an error $\delta$. Since quantifying the approximation error $\delta$ is not our focus, we henceforth leave it as an open parameter. With this, we are ready to state a uniform deviation inequality for neural estimators of Rényi divergences.
Theorem 4 (Uniform deviation inequality for neural estimator).
Let $\beta>d/2$, $M\geq 1$, $a:=\frac{d}{2\beta}$, $b:=\log M$, $\mathcal{G}_{b,\beta,\rho}$ be a neural class satisfying Assumption 2, and $c_{d,\beta}$ be such that $1\leq c_{d,\beta}\lesssim\frac{\sqrt{\hat{c}_{d,\beta}}}{1-a}$, where $\hat{c}_{d,\beta}$ is as given in (17). Then, the following hold:
  1. (i)
    KL divergence:
[…]
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\leq 2e^{-(nz^{2}\wedge nz)}.$ (18)
  2. (ii)
    Rényi divergence $\big{(}\alpha\in(0,1)\cup(1,\infty)\big{)}$:
    $\sup_{(\mu,\nu)\in\bar{\mathcal{P}}_{M,\rho}(\mathcal{G}_{b,\beta,\rho},\delta)}\mathbb{P}\left(\left|\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{b,\beta,\rho}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)\right|\geq\big{(}M^{2\alpha}+M^{2\left|\alpha-1\right|}\big{)}\delta+c_{d,\beta}v_{M,\alpha,a}\big{(}n^{-\frac{1}{2}}+z\big{)}\right)$
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\leq 2e^{-(nz^{2}\wedge nz)},$ (19)
    where $v_{M,\alpha,a}=\frac{\alpha M^{2|\alpha-1|}}{\left|\alpha-1\right|}\left(1+(\left|\alpha-1\right|\log M)^{\frac{a}{2}}\right)+M^{2\alpha}\big{(}1+(\alpha\log M)^{a}\big{)}$.
The proof of Theorem 4 is given in Section 4.6 and relies on an application of Talagrand’s concentration inequality for uniformly bounded function class.
[…]
3.3. Application: Auditing Rényi DP
Here, we consider an application of Theorem 4 to auditing DP, which was introduced in [21] as an approach for quantifying privacy leakage of privatization mechanisms. We recall some DP notions that are relevant to our setting. Consider a set $\mathfrak{U}$ with a relation $\sim$ such that $u\sim v$, for $u,v\in\mathfrak{U}$, denotes that $u$ and $v$ are adjacent. In the DP context, $\mathfrak{U}$ is a set of databases and $u\sim v$ denotes that $u$ and $v$ are adjacent databases, differing on a single entry.
Definition 6 (DP mechanisms[21, 20, 79, 47]).
Let $\epsilon,\delta\geq 0$. A randomized (measurable) mechanism $f:\mathfrak{U}\rightarrow\mathbb{R}^{d}$ is
  1. (i)
    $\epsilon$-differentially private if $\mathbb{P}\left(f(u)\in\mathcal{T}\right)\leq e^{\epsilon}\,\mathbb{P}\left(f(v)\in\mathcal{T}\right)$ for every $u\sim v$ and $\mathcal{T}\in\mathcal{B}(\mathbb{R}^{d})$;
  2. (ii)
    $(\epsilon,\delta)$-differentially private if $\mathbb{P}\left(f(u)\in\mathcal{T}\right)\leq e^{\epsilon}\,\mathbb{P}\left(f(v)\in\mathcal{T}\right)+\delta$ for every $u\sim v$ and $\mathcal{T}\in\mathcal{B}(\mathbb{R}^{d})$;
  3. (iii)
    $\epsilon$- Rényi differentially private of order $\alpha$ if $\mathsf{D}_{\alpha}\left(\mu_{u}\middle\|\mu_{v}\right)\leq\epsilon$ for every $u\sim v$, where $\mu_{u}\in\mathcal{P}(\mathbb{R}^{d})$ is the distribution of $f(u)$.
As $\epsilon$-DP is equivalent to $\sup_{u\sim v}\mathsf{D}_{\max}(\mu_{u}\|\mu_{v})\leq\epsilon$, where $\mathsf{D}_{\max}$ is the $\infty$-order Rényi divergence and $\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)$ is non-decreasing in $\alpha$ [76, Theorem 3], it is clear that Rényi DP of any finite order $\alpha$ is a relaxation of DP. However, the two notions become equivalent at $\alpha=\infty$, and hence $\epsilon$- Rényi DP is a good approximation of $\epsilon$-DP for sufficiently large $\alpha$.
In practice, given output samples from a privacy mechanism, one encounters the problem of ascertaining whether the mechanism is differentially private or not, referred to as auditing DP. This can be framed as a binary hypothesis testing problem where the null $H_{0}$ and the alternative $H_{1}$ corresponds to when privacy holds and does not hold, respectively. Note that the privacy mechanisms to be audited, hence also the output distributions $\mu_{u},\mu_{v}$, are usually unknown. Hence, auditing DP cannot be reduced to the problem of hypothesis testing between probability distributions and computing Rényi divergence to ascertain whether the mechanism is differentially private or not.
We next formally state the relevant hypothesis testing problem. Let $f:\mathfrak{U}\to\mathbb{R}^{d}$ be a privacy mechanism and denote a pair of adjacent databases [^5] Note that the current model subsumes the case of deterministic $(U,V)$ by taking $\tilde{\pi}$ to be a point mass on a pair of adjacent databases. In this case, the randomness in $\big{(}f(U),f(V))$ only comes from the mechanism. by $(U,V)\sim\tilde{\pi}\in\mathcal{P}(\mathfrak{U}\times\mathfrak{U})$. Let $\pi:=(f,f)_{\#}\tilde{\pi}$ be the joint distribution of $\big{(}f(U),f(V)\big{)}$, where #\# is the pushforward operation. Let $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ $\sim\pi$ be pairwise i.i.d. samples of the privacy mechanism’s output when acting on i.i.d. pairs of adjacent databases, where $\pi=\pi_{0}$ under $H_{0}$ and $\pi=\pi_{1}$ under $H_{1}$. With the marginals of $\pi_{0}$ and $\pi_{1}$ denoted by $(\mu_{0},\nu_{0})$ and $(\mu_{1},\nu_{1})$, respectively, the relevant hypothesis test for auditing Rényi DP of order $\alpha$ is
$\begin{split}&H_{0}:\mathsf{D}_{\alpha}\left(\mu_{0}\middle\|\nu_{0}\right)\leq\epsilon,\\ &H_{1}:\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)>\epsilon.\end{split}$ (20)
Denote the empirical measures of $(X_{1},\ldots,X_{n})$ and $(Y_{1},\ldots,Y_{n})$ by $\hat{\mu}_{n}$ and $\hat{\nu}_{n}$, respectively. For a test statistic $T_{n}=T_{n}(X_{1},\ldots,X_{n},Y_{1},\ldots,$ $Y_{n})$, a standard class of tests rejects $H_{0}$ if $T_{n}>t_{n}$ for some $t_{n}\in\mathbb{R}$. The type I and type II error probabilities achieved by the test $T_{n}$ are $e_{1,n}(T_{n},t_{n})=\mathbb{P}(T_{n}>t_{n}|H=H_{0})$ and $e_{2,n}(T_{n},t_{n})=\mathbb{P}(T_{n}\leq t_{n}|H=H_{1})$, respectively, where $H$ denotes the true hypothesis.
In [18], a hypothesis test for auditing DP using regularized kernel Rényi divergence is proposed, where the null hypothesis is that the mechanism satisfies $(\epsilon,\delta)$-DP. A decision rule achieving any type I error probability is proposed, leaving the characterization of the type II error probability open. In [67], a hypothesis testing based framework for auditing DP using KL divergence as the metric was proposed along with full asymptotic performance guarantees that relied on the knowledge of limit distribution for smoothed KL divergence. Here, we propose a neural estimator based hypothesis testing framework for auditing Rényi DP along with finite sample performance guarantees. We will assume that the neural net class $\mathcal{G}_{n}$ used for the audit and the marginal distributions $(\mu_{i},\nu_{i})$, $i=0,1$, satisfy the following assumption.
Assumption 3 (Assumptions on the neural class and distributions).
Let $\alpha\geq 1$, $M\geq 1$, $\beta>d/2$, and $b=\log M$. The neural net class $\mathcal{G}_{n}$ and $(\mu_{i},\nu_{i})$, $i=0,1$, satisfy $\mathcal{G}_{n}\subseteq\mathcal{G}_{b,\beta,\rho}$ and $(\mu_{i},\nu_{i})\in\mathcal{P}_{M,\beta}(\mathcal{G}_{n},\delta_{n})$, for some $\delta_{n}=o(1)$. Further, $\mathsf{D}_{\alpha}\left(\mu_{0}\middle\|\nu_{0}\right)\leq\epsilon$ and $\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)>\epsilon$.
Assumption 3 is not much restrictive in practice. Indeed, the definition of DP itself necessitates that $\left\|p_{\mu_{u}}/p_{\mu_{v}}\right\|_{\infty,\rho}$ is bounded [^6] Here, $\rho$ can be taken to be the Lebesgue measure in the continuous distribution setting or average of probability mass functions in the discrete setting. uniformly for all $u,v\in\mathfrak{U}$ with $u\sim v$. Moreover, as alluded to earlier, it is known that even restricting to shallow neural nets, the class of distributions $\mathcal{P}_{M,\beta}(\mathcal{G}_{n},\delta_{n})$ with $\delta_{n}=O(n^{-1/2})$ is sufficiently rich. In particular, for $\beta>d/2$ and $M$ sufficiently large, such a class includes discrete distributions that are mutually absolutely continuous or compactly supported continuous distributions with bounded ratios of densities. Hence, Assumption 3 allows for audit of most privacy mechanisms of practical interest.
The next result proposes a test statistic for Rényi DP audit and characterizes the associated error probabilities (see Section 4.7 for proof).
Proposition 2 (DP audit guarantees).
[…]
$t_{n}=\epsilon+\big{(}M^{2\alpha}+M^{2\left|\alpha-1\right|}\big{)}\delta_{n}+c_{d,\beta}v_{M,\alpha,a}\big{(}n^{-\frac{1}{2}}+n^{\frac{\tau-1}{2}}\big{)},$
satisfies
$e_{1,n}(T_{n},t_{n})$ $\leq 2e^{-n^{\tau}},$
$e_{2,n}(T_{n},t_{n})$ $\leq 2e^{-\vartheta_{n,\alpha,\delta_{n}}(\tau)},$
where $\vartheta_{n,\alpha,\delta_{n}}(\tau)$ is given in (61) below.
We note that the performance guarantees given above, though derived for auditing DP, apply to any hypothesis test of a similar kind using Rényi divergence as the metric to differentiate the hypotheses.
4. Proofs
We first state a few definitions that will be used throughout this section. For $\alpha\in(0,\infty)$, let
$\mathcal{F}_{\sigma}(\mathcal{F}):=\big{\{}f*\varphi_{\sigma}:f\in\mathcal{F}\big{\}}\quad\mbox{and }\quad\mathcal{E}_{\sigma,\alpha}(\mathcal{F}):=\big{\{}e^{\alpha f}*\varphi_{\sigma}:f\in\mathcal{F}\big{\}},$ (21)
be the Gaussian-smoothed function classes obtained by convolving the functions in $\mathcal{F}$ and $e^{\alpha\mathcal{F}}:=\{e^{\alpha f}:f\in\mathcal{F}\}$ by $\varphi_{\sigma}$, respectively. We will denote $\mathcal{E}_{\sigma,1}(\mathcal{F})$ by $\mathcal{E}_{\sigma}(\mathcal{F})$ henceforth.
For $\delta\geq 0$ and $\mathcal{F}$ with envelope $F$, define the entropy integral (see (3) for definition of covering entropy):
$J(\delta,F,\mathcal{F},\mu):=\int_{0}^{\delta}\sup_{\gamma}\sqrt{\log N(\epsilon\left\|F\right\|_{2,\gamma},F,\left\|\cdot\right\|_{2,\gamma})}d\epsilon.$ (22)
Here, the supremum is over all probability measures $\gamma$ with discrete finite support contained in that of $\mu$.
4.1. Proof of Theorem 1
We require the following lemma which provides a general deviation inequality for Gaussian-smoothed KL divergence.
Lemma 2 (Deviation inequality for KL divergence).
Let $\sigma>0$, $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$, $X^{n}\sim\mu^{\otimes n}$ and $Y^{n}\sim\nu^{\otimes n}$. Suppose $\mathcal{F}_{n}$ with envelope $F_{n}$ is such that $\mathcal{F}_{n,\sigma}:=\mathcal{F}_{\sigma}(\mathcal{F}_{n})$ and $\mathcal{E}_{n,\sigma}:=\mathcal{E}_{\sigma}(\mathcal{F}_{n})$ with envelopes [^7] By Young’s convolution inequality, a sufficient condition for existence of a measurable envelope $F_{n}*\varphi_{\sigma}$ for $\mathcal{F}_{n,\sigma}$ is $F_{n}\in L^{p}(\mathbb{R}^{d})$ for some $p\geq 1$. $F_{n}*\varphi_{\sigma}$ and $e^{F_{n}}*\varphi_{\sigma}$, respectively, satisfy Assumption 1. Further, assume that $f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star},f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}\in\mathcal{F}_{n}$ a.s., and $F_{n}$ satisfies
$\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}\vee\left\|e^{F_{n}}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}}<\infty,$ (23)
for some $q\in(0,1]$ and every $n\in\mathbb{N}$, where $X\sim\mu$ and $Y\sim\nu$. Then
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\gtrsim_{q}n^{-\frac{1}{2}}t_{n,\sigma,\mu,\nu}+z\right)$
$\qquad\qquad\qquad\qquad\leq e^{-R_{n,q,\sigma,\mu}(z)}+e^{-\tilde{R}_{n,q,\sigma,\nu}(z)}+3e^{-O_{n,q,\sigma,\mu}(z)}+3e^{-\tilde{O}_{n,q,\sigma,\nu}(z)},$ (24)
where
$t_{n,\sigma,\mu,\nu}:=J(1,F_{n}*\varphi_{\sigma},\mathcal{F}_{n,\sigma},\mu)\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}+J(1,e^{F_{n}}*\varphi_{\sigma},\mathcal{E}_{n,\sigma},\nu)\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{2,\nu},$ (25a)
[…]
$O_{n,q,\sigma,\mu}(z):=\frac{n^{q}z^{q}}{\left\|M_{n,\mathcal{F}_{n,\sigma}}(X^{n})\right\|_{\psi_{q}}^{q}}\quad\mbox{ and }\quad\tilde{O}_{n,q,\sigma,\nu}(z):=\frac{n^{q}z^{q}}{\left\|\tilde{M}_{n,\mathcal{E}_{n,\sigma}}(Y^{n})\right\|_{\psi_{q}}^{q}},$ (25c)
$M_{n,\mathcal{F}_{n,\sigma}}(X^{n}):=\max_{1\leq i\leq n}F_{n}*\varphi_{\sigma}(X_{i})\quad\mbox{ and }\quad\tilde{M}_{n,\mathcal{E}_{n,\sigma}}(Y^{n}):=\max_{1\leq i\leq n}e^{F_{n}}*\varphi_{\sigma}(Y_{i}),$ (25d)
[…]
$\check{M}_{n,\mathcal{E}_{n,\sigma}}\vee\left\|\tilde{M}_{n,\mathcal{E}_{n,\sigma}}(Y^{n})\right\|_{\psi_{q}}\lesssim_{q}\left(\log(1+n)\right)^{\frac{1}{q}}\left\|e^{F}_{n}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}}<\infty.$ (25g)
When $\mu,\nu$ are compactly supported and (23) is replaced by the condition
$\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}\vee\left\|F_{n}*\varphi_{\sigma}\right\|_{\infty,\mu}\vee\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{2,\nu}\vee\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{\infty,\nu}<\infty,$
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\gtrsim n^{-\frac{1}{2}}t_{n,\sigma,\mu,\nu}+z\right)$
$\qquad\qquad\qquad\qquad\leq e^{-\left(\frac{nz^{2}}{\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}^{2}}\wedge\frac{nz}{\left\|F_{n}*\varphi_{\sigma}\right\|_{\infty,\mu}}\right)}+e^{-\left(\frac{nz^{2}}{\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{2,\mu}^{2}}\wedge\frac{nz}{\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{\infty,\nu}}\right)}.$ (26)
The proof of Lemma 2 is given in Section 5.1. The key idea in its proof is to use the variational expression (1b) to upper bound the deviation $|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)|$ in terms of sum of suprema of empirical processes indexed by $\mathcal{F}_{n,\sigma}$ and $\mathcal{E}_{n,\sigma}$. Then, (24) and (26) follow by invoking a Fuk-Nagaev type inequality given in [2, Theorem 4] (see Theorem 7 below) and Talagrand’s concentration inequality given in [12, Theorem 2.3] (see Theorem 6 below), respectively.
Remark 5 (One-sample case).
From the proof of Lemma 2, it is easy to see that (24) and (26) holds with $\nu*\gamma_{\sigma}$ and $\hat{\nu}_{n}*\gamma_{\sigma}$ replaced with $\nu$ such that $\mu*\gamma_{\sigma},\hat{\mu}_{n}*\gamma_{\sigma}\ll\nu$ if the integrability conditions on only $F_{n}$ hold, i.e., $\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}<\infty$ and $\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}\vee\left\|F_{n}*\varphi_{\sigma}\right\|_{\infty,\mu}<\infty$, respectively. For instance, (24) simplifies to
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu\right)\right|\gtrsim_{q}n^{-\frac{1}{2}}J(1,F_{n}*\varphi_{\sigma},\mathcal{F}_{n,\sigma},\mu)\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}+z\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq e^{-R_{n,q,\sigma,\mu}(z)}+3e^{-O_{n,q,\sigma,\mu}(z)}.$ (27)
Also, by considering the corresponding simplication for (26), the following improvement of (8) holds in the one-sample setting where $\nu$ is known and only $\mu$ is estimated as $\hat{\mu}_{n}$:
[…]
where $\tilde{\xi}_{r,d,\sigma}$ is given in (36) below, and
$\bar{\xi}_{r,d,\sigma}:=(r^{2}+4r)(1+r\sigma\sqrt{d})\sigma^{-2}.$
To prove Part $(i)$, fix $\mu,\nu\in\mathcal{P}\big{(}\mathbb{B}_{d}(r)\big{)}$. We may assume that $r>0$ since (8) trivially holds for $r=0$. Since $\mathrm{spt}(\mu),\mathrm{spt}(\nu)\subseteq\mathbb{B}_{d}(r)$, $\max_{1\leq i\leq n}\left\|X_{i}\right\|\vee\left\|Y_{i}\right\|\leq r$. Then
$\frac{\hat{\mu}_{n}*\varphi_{\sigma}(x)}{\hat{\nu}_{n}*\varphi_{\sigma}(x)}=\frac{\sum_{i=1}^{n}e^{-\frac{\left\|x-X_{i}\right\|^{2}}{2\sigma^{2}}}}{\sum_{i=1}^{n}e^{-\frac{\left\|x-Y_{i}\right\|^{2}}{2\sigma^{2}}}}\leq\frac{e^{-\min_{i}\frac{\left\|x-X_{i}\right\|^{2}}{2\sigma^{2}}}}{e^{-\max_{i}\frac{\left\|x-Y_{i}\right\|^{2}}{2\sigma^{2}}}}\leq\frac{e^{-\min_{i}\frac{\left\|x\right\|^{2}+\left\|X_{i}\right\|^{2}-2\left\|x\right\|\left\|X_{i}\right\|}{2\sigma^{2}}}}{e^{-\max_{i}\frac{\left\|x\right\|^{2}+\left\|Y_{i}\right\|^{2}+2\left\|x\right\|\left\|Y_{i}\right\|}{2\sigma^{2}}}}\leq e^{\frac{r^{2}+4r\left\|x\right\|}{2\sigma^{2}}}.$
Similarly,
$\frac{\mu*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}=\frac{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-X\right\|^{2}}{2\sigma^{2}}}d\mu(x)}{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-Y\right\|^{2}}{2\sigma^{2}}}d\nu(y)}\leq\frac{e^{-\min_{\left\|x^{\prime}\right\|\leq r}\frac{\left\|x-x^{\prime}\right\|^{2}}{2\sigma^{2}}}}{e^{-\max_{\left\|y\right\|\leq r}\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}}\leq\frac{e^{-\frac{(\left\|x\right\|^{2}-2r\left\|x\right\|)}{2\sigma^{2}}}}{e^{-\frac{(\left\|x\right\|^{2}+r^{2}+2\left\|x\right\|r)}{2\sigma^{2}}}}\leq e^{\frac{r^{2}+4r\left\|x\right\|}{2\sigma^{2}}}.$
By similar steps, the same bounds also hold for the ratios $\nu*\varphi_{\sigma}(x)/\mu*\varphi_{\sigma}(x)$ and $\hat{\nu}_{n}*\varphi_{\sigma}(x)/\hat{\mu}_{n}*\varphi_{\sigma}(x)$. Consequently, we have
$\left|f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star}(x)\right|\vee\left|f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}(x)\right|\leq\frac{r^{2}+4r\left\|x\right\|}{2\sigma^{2}}.$
Hence
$f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star},f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}\in\mathcal{F}\big{(}b_{r,\sigma},1\big{)}\mbox{ with }b_{r,\sigma}:=\frac{r^{2}+4r}{\sigma^{2}},$ (29a)
where $\mathcal{F}(b,q):=\left\{f:\mathbb{R}^{d}\rightarrow\mathbb{R},\left|f(x)\right|\leq 0.5b(1+\left\|x\right\|^{q})\right\}.$ (29b)
For applying Lemma 2, define
$\mathcal{F}_{\sigma}(b,q)$ $:=\{f*\varphi_{\sigma}:f\in\mathcal{F}(b,q)\},$ (30a)
$\mathcal{E}_{\sigma,\alpha}(b,q)$ $:=\{e^{\alpha f}*\varphi_{\sigma}:f\in\mathcal{F}(b,q)\},$ (30b)
$F_{b,q}(x)$ $:=0.5b\big{(}1+\left\|x\right\|^{q}\big{)}.$ (30c)
Set $\mathcal{F}_{\sigma}(b):=\mathcal{F}_{\sigma}(b,1)$, $\mathcal{E}_{\sigma}(b):=\mathcal{E}_{\sigma,1}(b,1)$, $F_{b}(x)=0.5b(1+\left\|x\right\|)$, and for $\mathcal{X}\subseteq\mathbb{R}^{d}$
$v_{b,q}(\mathcal{X}):=\sup_{x\in\mathcal{X}}0.5b\big{(}1+\left\|x\right\|^{q}\big{)}.$ (31)
We will also require a bound on the entropy integrals, $J(1,F_{b_{r,\sigma}}*\varphi_{\sigma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\mu)$ and $J(1,e^{F_{b_{r,\sigma}}}*\varphi_{\sigma},\mathcal{E}_{\sigma}(b_{r,\sigma}),\nu)$. To this end, we first obtain a bound on the covering entropy of $\mathcal{F}_{\sigma}(b)$.
Lemma 3 (Entropy bounds for smoothed function class).
Let $0<\sigma\leq 1$, $b\geq 0$, $0<\beta<\infty$, and $q\in\{0\}\cup[1,\infty)$. Then, for all $f\in\mathcal{F}(b,q)$ and $\mathcal{X}\subseteq\mathbb{R}^{d}$ with non-empty interior,
[…]
where $\check{c}_{d,\beta,q}$ is specified in (79) below. Further, for any $p\geq 1$ and $\gamma\in\mathcal{P}\big{(}\mathbb{B}_{d}(r)\big{)}$,
[…]
where $c_{d,\beta,q}$ is given in (82) below. The same upper bounds in (33a)-(33b) also holds with $\mathcal{F}_{\sigma}(b,q)$ and $F_{b,q}$ replaced by $\mathcal{E}_{\sigma,\alpha}(b,q)$ and $e^{\alpha F_{b,q}}$, respectively. Moreover, the above upper bounds with $q=1$ also applies to all $q\in[0,1]$ up to a universal multiplicative factor, and $c_{d,\frac{d}{2}+1,q}=O\big{(}d^{6d+9}(9e^{2})^{d}\big{)}$ for all $q\in[0,1]$.
The proof of Lemma 3 is given in Section 5.2 and relies on a systematic analysis of the difference of derivatives of a function belonging to $\mathcal{F}_{\sigma}(b,q)$.
Let $\beta=\frac{d}{2}+1$. It follows from (33b) that for all probability measures $\gamma$ such that $\mathrm{spt}(\gamma)\subseteq\mathbb{B}_{d}(r)$, we have
$\log N\Big{(}\epsilon\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\left\|\cdot\right\|_{2,\gamma}\Big{)}\lesssim d^{6d+9}(9e^{2})^{d}(r+1)^{d+2}\sigma^{-d}\epsilon^{-\frac{d}{\beta}}.$
Hence
$J(1,F_{b_{r,\sigma}}*\varphi_{\sigma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\mu)$ $:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon$
$\lesssim(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}},$
where we used that $\int_{0}^{1}\epsilon^{-\frac{d}{2\beta}}d\epsilon\lesssim d$ for $\beta=\frac{d}{2}+1$. Similarly,
$J(1,e^{F_{b_{r,\sigma}}}*\varphi_{\sigma},\mathcal{E}_{\sigma}(b_{r,\sigma}),\nu)\lesssim(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}}.$ (34)
We next bound $\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\mu}$ and $\left\|e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}\right\|_{2,\nu}$. For this purpose, we will use the following relation between sub-Gaussian and norm sub-Gaussian distributions. $\mu\in\mathcal{P}\big{(}\mathbb{R}^{d}\big{)}$ is $\sigma^{2}$-norm sub-Gaussian for $\sigma>0$ if $X\sim\mu$ satisfies
$\mu\big{(}\left\|X-\mathbb{E}[X]\right\|>t\big{)}\leq 2e^{\frac{-t^{2}}{2\sigma^{2}}},\quad\forall~t\in\mathbb{R}.$
Lemma 4.
[35, Lemma 1, Lemma 2] If $\mu\in\mathcal{P}\big{(}\mathbb{R}^{d}\big{)}$ is $\sigma^{2}$-sub-Gaussian, then it is $8d\sigma^{2}$-norm sub-Gaussian. In particular, if $X\sim\mu$ and $\mathbb{E}[X]=0$, then $\left(\mathbb{E}[\left\|X\right\|^{p}]\right)^{\frac{1}{p}}\lesssim\sqrt{dp}\sigma$ and $\mathbb{E}\big{[}e^{t\left\|X\right\|}\big{]}\leq e^{4t^{2}d\sigma^{2}}$.
We have
$F_{b_{r,\sigma}}*\varphi_{\sigma}(x)$ $=\frac{1}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}F_{b_{r,\sigma}}(y)e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
$=\frac{r^{2}+4r}{2(2\pi)^{\frac{d}{2}}\sigma^{d+2}}\int_{\mathbb{R}^{d}}(1+\left\|y\right\|)e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
$\lesssim\frac{(r^{2}+4r)(1+\left\|x\right\|\sigma\sqrt{d})}{\sigma^{2}},$
where to obtain the last inequality, we used $\left\|y\right\|\leq\left\|x\right\|+\left\|y-x\right\|$ and the moment bound in Lemma 4 for $p=1$. Hence,
$\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\mu}\vee\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{\infty,\mu}\lesssim\frac{(r^{2}+4r)\big{(}1+r\sigma\sqrt{d}\big{)}}{\sigma^{2}}.$
Similarly,
$e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}(x)$ $=\frac{1}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}e^{F_{b_{r,\sigma}}(y)}e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
$\leq\frac{e^{\frac{(r^{2}+4r)(1+\left\|x\right\|)}{2\sigma^{2}}}}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}e^{\frac{(r^{2}+4r)\left\|y-x\right\|-\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
$\leq e^{\frac{(r^{2}+4r)(1+\left\|x\right\|)+2d(r^{2}+4r)^{2}}{2\sigma^{2}}},$
where we used the exponential moment bound in Lemma 4. Hence
$\left\|e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}\right\|_{2,\nu}\vee\left\|e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}\right\|_{\infty,\nu}\leq e^{\frac{(r^{2}+4r)\left(2dr^{2}+(8d+1)r+1\right)}{2\sigma^{2}}}=:\xi_{r,d,\sigma}.$ (35)
Substituting the above bounds, we obtain for $\sigma\leq 1$ that
$t_{n,\sigma,\mu,\nu}$
$\lesssim J(1,F_{b_{r,\sigma}}*\varphi_{\sigma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\mu)\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\mu}+J(1,e^{F_{b_{r,\sigma}}}*\varphi_{\sigma},\mathcal{E}_{\sigma}(b_{r,\sigma}),\nu)\left\|e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}\right\|_{2,\nu}$
$\lesssim(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}}\left(\frac{(r^{2}+4r)(1+r\sigma\sqrt{d})}{\sigma^{2}}+e^{\frac{(r^{2}+4r)(2dr^{2}+(8d+1)r+1)}{2\sigma^{2}}}\right)$
$\lesssim\tilde{\xi}_{r,d,\sigma}\xi_{r,d,\sigma},$
where
$\tilde{\xi}_{r,d,\sigma}:=(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}}.$ (36)
Applying (26) and taking supremum over $\mu,\nu\in\mathcal{P}\big{(}\mathbb{B}_{d}(r)\big{)}$, we obtain (8).
Next, consider Part $(ii)$. Fix $\mu,\nu\in\mathsf{SG}(L)$. We have
[…]
where $(a)$ follows by upper bounding average by the maximum and Cauchy-Schwarz, and $(b)$ is via an application of Jensens inequality to convex function $e^{x}$. Hence, we have
$\log\left(\frac{\hat{\mu}_{n}*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}\right)$ $\leq\frac{\left\|x\right\|\max_{i}\left\|X_{i}\right\|}{\sigma^{2}}+\frac{\mathbb{E}_{\nu}[\left\|Y\right\|^{2}]+2x\cdot\mathbb{E}[Y]}{2\sigma^{2}}$
$\leq\frac{\left\|x\right\|\max_{i}\left\|X_{i}\right\|+(L^{2}+16dL)+\left\|x\right\|L}{\sigma^{2}},$
where we used that since
$\mathbb{E}_{\nu}[\left\|Y\right\|^{2}]\leq 2\left(\left\|\mathbb{E}_{\nu}[Y]\right\|^{2}+\mathbb{E}_{\nu}[\left\|Y-\mathbb{E}_{\nu}[Y]\right\|^{2}]\right)\leq 2(L^{2}+16dL).$
On the other hand,
$\frac{\nu*\varphi_{\sigma}(x)}{\hat{\mu}_{n}*\varphi_{\sigma}(x)}$ $=\frac{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\nu(y)}{\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{\left\|x-X_{i}\right\|^{2}}{2\sigma^{2}}}}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{\mathbb{E}_{\nu}\left[e^{-\frac{(\left\|x\right\|^{2}+\left\|Y\right\|^{2}-2x\cdot Y)}{2\sigma^{2}}}\right]}{e^{-\max_{i}\frac{\left\|x\right\|^{2}+\left\|X_{i}\right\|^{2}-2x\cdot X_{i}}{2\sigma^{2}}}}$
$\leq e^{\max_{i}\frac{(\left\|X_{i}\right\|^{2}-2x\cdot X_{i})}{2\sigma^{2}}}\mathbb{E}_{\nu}\left[e^{\frac{x\cdot Y}{\sigma^{2}}}\right]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}e^{\max_{i}\frac{(\left\|X_{i}\right\|^{2}-2x\cdot X_{i})}{2\sigma^{2}}}e^{\frac{L\left\|x\right\|^{2}}{2\sigma^{4}}+\frac{L\left\|x\right\|}{\sigma^{2}}},$ (38)
where $(a)$ follows by lower bounding the average by the minimum, and $(b)$ is because $\nu\in\mathsf{SG}(L)$ implies
$\mathbb{E}_{\nu}\left[e^{\frac{x\cdot Y}{\sigma^{2}}}\right]\leq e^{\frac{\left\|x\right\|^{2}L}{\sigma^{4}}+\frac{L\left\|x\right\|}{\sigma^{2}}}.$
From (37) and (38),
$\left|\log\left(\frac{\hat{\mu}_{n}*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}\right)\right|\leq\frac{L\left\|x\right\|^{2}}{2\sigma^{4}}+\left\|x\right\|\left(\frac{L+\max_{i}\left\|X_{i}\right\|}{\sigma^{2}}\right)+\max_{i}\frac{\left\|X_{i}\right\|^{2}}{2\sigma^{2}}+\frac{L^{2}+16dL}{\sigma^{2}}.$ (39)
We will apply (27) to arrive at the desired conclusion. However, for this purpose, we require the RHS of (39) be bounded almost surely. Since this does not hold for $\mu\in\mathsf{SG}(L)$ in general due to the presence of the term $\max_{i}\left\|X_{i}\right\|^{2}$, we will use a truncation argument by considering the event
$E_{n}(t):=\left\{\max_{1\leq i\leq n}\left\|X_{i}\right\|\leq t\right\},$
under which (27) can be applied. The parameter $t$ is then scaled appropriately w.r.t. $n$ such that the probability of $\bar{E}_{n}(t)$ (the complement event of $E_{n}(t)$) decays exponentially.
Since $X\sim\mu$ satisfies $\left\|X\right\|_{\psi_{p}}\leq L$ for some $p\geq 2$ and $L\geq 1$, we have from [75, Lemma 2.2.2] that
$\left\|\max_{1\leq i\leq n}\left\|X_{i}\right\|\right\|_{\psi_{p}}\leq 2^{\frac{1}{p}}\left(\log(1+n)\right)^{\frac{1}{p}}\left\|X\right\|_{\psi_{p}}.$
By Markov’s inequality and definition of Orlicz norm $\psi_{p}$,
$\mathbb{P}\left(\max_{1\leq i\leq n}\left\|X_{i}\right\|\geq t\right)\leq 2e^{-\frac{t^{p}}{\left\|\max_{1\leq i\leq n}\left\|X_{i}\right\|\right\|_{\psi_{p}}^{p}}}.$ (40)
Setting $t=T_{n,p,\kappa,L}:=2^{\frac{1}{p}}Ln^{\kappa}$ for some $\kappa\geq 0$, we obtain from (40) that
$\mathbb{P}\big{(}\bar{E}_{n}(T_{n,p,\kappa,L})\big{)}\leq 2e^{-\frac{n^{p\kappa}}{\log(1+n)}}.$
For any $\bar{t}\geq 0$, we obtain
$\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\geq\bar{t}~\right)$
$\leq\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\geq\bar{t}~\Big{|}E_{n}\big{(}T_{n,p,\kappa,L}\big{)}\right)+\mathbb{P}\left(\bar{E}_{n}\big{(}T_{n,p,\kappa,L}\big{)}\right)$
$\leq\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\geq\bar{t}~\Big{|}E_{n}\big{(}T_{n,p,\kappa,L}\big{)}\right)+2e^{-\frac{n^{p\kappa}}{\log(1+n)}}.$ (41)
Next, we bound the first term in the RHS above. Note that given $E_{n}\big{(}T_{n,p,\kappa,L}\big{)}$ holds, we have $X^{n}\sim\tilde{\mu}^{\otimes n}$, where
$\tilde{\mu}(\mathcal{A})=\frac{\mu\left(\mathcal{A}\cap\mathbb{B}_{d}(T_{n,p,\kappa,L})\right)}{\mu\left(\mathbb{B}_{d}(T_{n,p,\kappa,L})\right)},~\forall~\mathcal{A}\in\mathcal{B}(\mathbb{R}^{d}).$
Then, (39) yields that under this event
$\left|f^{\star}_{\hat{\mu}_{n}*\varphi_{\sigma},\hat{\nu}_{n}*\varphi_{\sigma}}(x)\right|\leq\frac{L\left\|x\right\|^{2}}{2\sigma^{4}}+\left\|x\right\|\left(\frac{L+T_{n,p,\kappa,L}}{\sigma^{2}}\right)+\frac{T_{n,p,\kappa,L}^{2}}{2\sigma^{2}}+\frac{L^{2}+16dL}{\sigma^{2}}.$
Further
$\frac{\mu*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}$ $=\frac{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\mu(y)}{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\nu(y)}$
$\leq\frac{\mathbb{E}_{\mu}\left[e^{\frac{x\cdot Y}{\sigma^{2}}}\right]}{\mathbb{E}_{\nu}\left[e^{\frac{2x\cdot Y-\left\|Y\right\|^{2}}{2\sigma^{2}}}\right]}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{e^{\frac{\left\|x\right\|^{2}L}{2\sigma^{4}}+\frac{\left\|x\right\|L}{\sigma^{2}}}}{e^{\frac{2x\cdot\mathbb{E}_{\nu}[Y]-\mathbb{E}_{\nu}\left[\left\|Y\right\|^{2}\right]}{2\sigma^{2}}}}$
$=e^{\frac{\left\|x\right\|^{2}L}{2\sigma^{4}}+\frac{\left\|x\right\|L}{\sigma^{2}}+\frac{\mathbb{E}_{\nu}\left[\left\|Y\right\|^{2}\right]-2x\cdot\mathbb{E}_{\nu}[Y]}{2\sigma^{2}}}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}e^{\frac{\left\|x\right\|^{2}L}{2\sigma^{4}}+\frac{(L^{2}+16dL)+2\left\|x\right\|L}{\sigma^{2}}},$
where $(a)$ used that $\mu\in\mathsf{SG}(L)$ as well as Jensen’s inequality applied to convex function $e^{x}$, while $(b)$ is due to $\nu\in\mathsf{SG}(L)$ . Similar bound also holds for $\nu*\varphi_{\sigma}(x)/\mu*\varphi_{\sigma}(x)$. Hence, given $E_{n}\big{(}T_{n,p,\kappa,L}\big{)}$ holds, (39) and above equation implies
$\left|f^{\star}_{\mu*\varphi_{\sigma},\nu*\varphi_{\sigma}}(x)\right|$ $\leq\frac{L\left\|x\right\|^{2}}{2\sigma^{4}}+\frac{2\left\|x\right\|(L+T_{n,p,\kappa,L})}{\sigma^{2}}+\frac{2L^{2}+32dL+T_{n,p,\kappa,L}^{2}}{2\sigma^{2}}$
$\leq 2\left(\frac{T_{n,p,\kappa,L}+8dL}{\sigma^{2}}\right)^{2}\big{(}1+\left\|x\right\|^{2}\big{)}.$
It follows that
$f^{\star}_{\mu*\varphi_{\sigma},\nu*\varphi_{\sigma}},f^{\star}_{\mu*\varphi_{\sigma},\nu*\varphi_{\sigma}}\in\mathcal{F}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)}\mbox{ with }\bar{b}_{n,d,p,\kappa,\sigma,L}=4\left(\frac{T_{n,p,\kappa,L}+8dL}{\sigma^{2}}\right)^{2}.$
[…]
$\log N\Big{(}\epsilon\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\sigma}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\left\|\cdot\right\|_{2,\gamma}\Big{)}\lesssim d^{6d+9}(9e^{2})^{d}\big{(}T_{n,p,\kappa,L}+1\big{)}^{d+4}\sigma^{-d}\epsilon^{-\frac{d}{\beta}}.$
Hence
[…]
$:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\sigma}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon$
$\lesssim(3ed)^{3(d+2)}\big{(}T_{n,p,\kappa,L}+1\big{)}^{\frac{d}{2}+2}\sigma^{-\frac{d}{2}}$
$\leq(3ed)^{3(d+2)}2^{\frac{d}{2}+2}\Big{(}2^{\frac{1}{p}}L\Big{)}^{\frac{d}{2}+2}n^{\kappa\left(\frac{d}{2}+2\right)}\sigma^{-\frac{d}{2}}.$
By using $\left\|y\right\|^{2}\leq 2\big{(}\left\|x\right\|^{2}+\left\|x-y\right\|^{2}\big{)}$, we have
$\left|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}(x)\right|$ $\leq\int_{\mathbb{R}^{d}}0.5\bar{b}_{n,d,p,\kappa,\sigma,L}\big{(}1+2\left\|x\right\|^{2}+2\left\|x-y\right\|^{2}\big{)}\varphi_{\sigma}(x-y)dy$
$=\bar{b}_{n,d,p,\kappa,\sigma,L}\left(0.5+\left\|x\right\|^{2}+d\sigma^{2}\right).$
This implies that for $X^{n}\sim\tilde{\mu}^{\otimes n}$ with $\mathrm{spt}(\tilde{\mu})\subseteq\mathbb{B}_{d}(T_{n,p,\kappa,L})$ and any constant $C>0$,
$\mathbb{E}_{\tilde{\mu}}\left[e^{\left(\frac{\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}(X)}{C}\right)^{q}}\right]\leq e^{\frac{\bar{b}_{n,d,p,\kappa,\sigma,L}^{q}\big{(}0.5+d\sigma^{2}+T_{n,p,\kappa,L}^{2}\big{)}^{q}}{C^{q}}}.$
Hence
$\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}$ $\leq\frac{\bar{b}_{n,d,p,\kappa,\sigma,L}\big{(}0.5+d\sigma^{2}+T_{n,p,\kappa,L}^{2}\big{)}}{(\log 2)^{\frac{1}{q}}}$
$\leq 4\left(\frac{T_{n,p,\kappa,L}+8dL}{\sigma^{2}}\right)^{2}\frac{\big{(}0.5+d\sigma^{2}+T_{n,p,\kappa,L}^{2}\big{)}}{(\log 2)^{\frac{1}{q}}}$
$\lesssim\left(\frac{3}{2}\right)^{\frac{1}{q}}L^{4}d^{3}\sigma^{-4}n^{4\kappa}.$
Further, since
$\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}\right\|_{p,\tilde{\mu}*\gamma_{\sigma}}^{p}$
[…]
$\leq 2^{2p-1}\bar{b}_{n,d,p,\kappa,\sigma,L}^{p}T_{n,p,\kappa,L}^{2p}+2^{2p-1}\bar{b}_{n,d,p,\kappa,\sigma,L}^{p}(dp)^{p}\sigma^{2p},$
$\lesssim\big{(}2^{12}L^{4}d^{3}pn^{4\kappa}\sigma^{-4}\big{)}^{p},$
we have
[…]
$J\big{(}1,\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma},\mathcal{F}_{\sigma}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\tilde{\mu}\big{)}\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\tilde{\mu}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\lesssim(3ed)^{3(d+2)}4^{\frac{d}{2}+2}L^{\frac{d}{2}+6}d^{3}\sigma^{-\left(\frac{d}{2}+4\right)}n^{\kappa\left(\frac{d}{2}+6\right)}.$
Substituting the above in (25), we obtain
$O_{n,1,\sigma,\tilde{\mu}}(z)$ $\gtrsim\frac{\sigma^{4}zn^{1-4\kappa}}{L^{4}d^{3}\log(n+1)},$
$R_{n,1,\sigma,\tilde{\mu}}(z)$ $\gtrsim\frac{nz^{2}}{\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\tilde{\mu}}^{2}}\wedge\frac{nz}{\log(n+1)\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}(X)\right\|_{\psi_{1}}}$
$\gtrsim\frac{n^{1-8\kappa}z^{2}\sigma^{8}}{L^{8}d^{6}}\wedge\frac{n^{1-4\kappa}z\sigma^{4}}{L^{4}d^{3}\log(n+1)}.$
Let $0\leq\tau\leq 1$. Set
$\kappa_{d,\tau}^{\star}$ $:=\frac{\tau}{d+12}$ (42a)
$\hat{\xi}_{d,L,\sigma}$ $:=(3ed)^{3(d+2)}(4L)^{\frac{d}{2}+2}\sigma^{-\frac{d}{2}},$ (42b)
$\check{\xi}_{d,L,\sigma}$ $:=L^{4}d^{3}\sigma^{-4}.$ (42c)
Then, substituting $\kappa=\kappa_{d,\tau}^{\star}$ in the above equations yields
$J\big{(}1,\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma},\mathcal{F}_{\sigma}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\tilde{\mu}\big{)}\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\tilde{\mu}}\lesssim\hat{\xi}_{d,L,\sigma}\check{\xi}_{d,L,\sigma}n^{\frac{\tau}{2}},$
$R_{n,1,\sigma,\tilde{\mu}}(z)\gtrsim\frac{n^{\frac{d+12-8\tau}{d+12}}z^{2}}{\check{\xi}_{d,L,\sigma}^{2}}\wedge\frac{n^{\frac{d+12-4\tau}{d+12}}z}{\check{\xi}_{d,L,\sigma}\log(n+1)},$
$O_{n,1,\sigma,\tilde{\mu}}(z)\gtrsim\frac{n^{\frac{d+12-8\tau}{d+12}}z}{\check{\xi}_{d,L,\sigma}\log(n+1)}.$
Applying (27) in Remark 5 and simplifying, we obtain
$\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\gtrsim\check{\xi}_{d,L,\sigma}\left(\hat{\xi}_{d,L,\sigma}n^{-\frac{1}{2}(1-\tau)}+z\right)\Big{|}E_{n}\big{(}T_{n,p,\kappa_{d,\tau}^{\star},L}\big{)}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq 4e^{-\check{\zeta}_{n,d,\tau}(z)},$
with
$\check{\zeta}_{n,d,\tau}(z):=n^{\frac{d+12-8\tau}{d+12}}z^{2}\wedge\frac{n^{\frac{d+12-4\tau}{d+12}}z}{\log(n+1)}.$
[…]
$\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\gtrsim\check{\xi}_{d,L,\sigma}\big{(}\hat{\xi}_{d,L,\sigma}n^{-\frac{1}{2}(1-\tau)}+z\big{)}\right)\leq 6e^{-\hat{\zeta}_{n,d,p,\tau}(z)},$
where
$\hat{\zeta}_{n,d,p,\tau}(z):=n^{\frac{d+12-8\tau}{d+12}}z^{2}\wedge\frac{n^{\frac{d+12-4\tau}{d+12}}z}{\log(n+1)}\wedge\frac{n^{\frac{p\tau}{d+12}}}{\log(n+1)}.$ (43)
Taking supremum over all $\mu,\nu\in\mathsf{SG}(L)$ such that $\left\|X\right\|_{\psi_{p}}\leq L$ yields (9), thus completing the proof.
4.2. Proof of Theorem 2
Lemma 5 (Deviation inequality in terms of Rademacher expectation).
Consider the setting of Lemma 2. Let
[…]
Then
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\geq 7n^{-1}\big{(}Z^{(\varepsilon)}_{n,\sigma}+\bar{Z}^{(\varepsilon)}_{n,\sigma}\big{)}+2z\right)$
[…]
In particular, for compactly supported $\mu,\nu$,
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\geq 7n^{-1}\big{(}Z^{(\varepsilon)}_{n,\sigma}+\bar{Z}^{(\varepsilon)}_{n,\sigma}\big{)}+2z\right)$
$\qquad\qquad\qquad\qquad\leq e^{-\left(\frac{nz^{2}}{8\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}^{2}}\wedge\frac{nz}{22\left\|F_{n}*\varphi_{\sigma}\right\|_{\infty,\mu}}\right)}+e^{-\left(\frac{nz^{2}}{8\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{2,\mu}^{2}}\wedge\frac{nz}{22\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{\infty,\nu}}\right)}.$ (45)
The proof of Lemma 5 follows similar to Lemma 2 up to (66), and then applies (12) along with union bound to obtain (44), and (85) with $\eta=1+\sqrt{6}$ to derive (45).
Theorem 2 follows from (45) by noting that for $\mu,\nu\in\mathcal{P}\left(\mathbb{B}_{d}(r)\right)$, and function classes $\mathcal{F}_{\sigma}(b_{r,\sigma})$ and $\mathcal{E}_{\sigma}(b_{r,\sigma})$ with envelopes $F_{b_{r,\sigma}}*\varphi_{\sigma}$ and $e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}$, respectively, we have
$\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\mu}\vee\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{\infty,\mu}\vee\left\|e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}\right\|_{2,\nu}\vee\left\|e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}\right\|_{\infty,\nu}\leq\xi_{r,d,\sigma},$
where $\xi_{r,d,\sigma}$ is given in (35).
4.3. Proof of Theorem 3
Fix $\mu,\nu\in\mathcal{P}\left(\mathbb{B}_{d}(r)\right)$. As shown in (29a), $f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star},f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}\in\mathcal{F}\big{(}b_{r,\sigma}\big{)}$, where $b_{r,\sigma}:=\frac{r^{2}+4r}{\sigma^{2}}$. Also, recall that $F_{b_{r,\sigma}}(x):=0.5b_{r,\sigma}\big{(}1+\left\|x\right\|\big{)}$ is an envelope for $\mathcal{F}\big{(}b_{r,\sigma}\big{)}$. Let
[…]
Note that for arbitrary functionals $A:\mathcal{F}\rightarrow\mathbb{R}$ and $B:\mathcal{F}\rightarrow\mathbb{R}$ such that at least one of $\sup_{f\in\mathcal{F}}A(f)$ and $\sup_{f\in\mathcal{F}}B(f)$ is finite, we have
$\left|\sup_{f\in\mathcal{F}}A(f)-\sup_{f\in\mathcal{F}}B(f)\right|\leq\sup_{f\in\mathcal{F}}\left|A(f)-B(f)\right|.$ (46)
Using this and observing that with probability one,
$0<\sup_{f\in\mathcal{F}(b_{r,\sigma})}\log L_{\alpha-1,\sigma,\hat{\mu}_{n}},\sup_{f\in\mathcal{F}(b_{r,\sigma})}\log L_{\alpha-1,\sigma,\mu},\sup_{f\in\mathcal{F}(b_{r,\sigma})}\log L_{\alpha,\sigma,\hat{\nu}_{n}},\sup_{f\in\mathcal{F}(b_{r,\sigma})}\log L_{\alpha,\sigma,\nu}<\infty,$
we obtain
$\left|\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|$
$\leq\sup_{f\in\mathcal{F}(b_{r,\sigma})}\frac{\alpha}{\left|\alpha-1\right|}\left|\log L_{\alpha-1,\sigma,\hat{\mu}_{n}}(f)-\log L_{\alpha-1,\sigma,\mu}(f)\right|+\sup_{f\in\mathcal{F}(b_{r,\sigma})}\left|\log L_{\alpha,\sigma,\hat{\nu}_{n}}(f)-\log L_{\alpha,\sigma,\nu}(f)\right|$
$\leq\frac{\alpha}{\left|\alpha-1\right|}\sup_{f\in\mathcal{F}(b_{r,\sigma})}\frac{\left|L_{\alpha-1,\sigma,\hat{\mu}_{n}}(f)-L_{\alpha-1,\sigma,\mu}(f)\right|}{L_{\alpha-1,\sigma,\hat{\mu}_{n}}(f)\wedge L_{\alpha-1,\sigma,\mu}(f)}+\sup_{f\in\mathcal{F}(b_{r,\sigma})}\frac{\left|L_{\alpha,\sigma,\hat{\nu}_{n}}(f)-L_{\alpha,\sigma,\nu}(f)\right|}{L_{\alpha,\sigma,\hat{\nu}_{n}}(f)\wedge L_{\alpha,\sigma,\nu}(f)}$
$\leq\frac{\alpha}{\left|\alpha-1\right|}\frac{\sup_{f\in\mathcal{F}(b_{r,\sigma})}\left|L_{\alpha-1,\sigma,\hat{\mu}_{n}}(f)-L_{\alpha-1,\sigma,\mu}(f)\right|}{L_{-|\alpha-1|,\sigma,\hat{\mu}_{n}}(F_{b_{r,\sigma}})\wedge L_{-|\alpha-1|,\sigma,\mu}(F_{b_{r,\sigma}})}+\frac{\sup_{f\in\mathcal{F}(b_{r,\sigma})}\left|L_{\alpha,\sigma,\hat{\nu}_{n}}(f)-L_{\alpha,\sigma,\nu}(f)\right|}{L_{-\alpha,\sigma,\hat{\nu}_{n}}(F_{b_{r,\sigma}})\wedge L_{-\alpha,\sigma,\nu}(F_{b_{r,\sigma}})},$ (47)
where the penultimate inequality used $\left|\log\frac{x}{y}\right|\leq\frac{|x-y|}{x\wedge y}$.
Note that since $\mathrm{spt}(\mu)\subseteq\mathbb{B}_{d}(r)$, we have
[…]
where we used that $\mathbb{E}[\left\|Z\right\|]=\sqrt{d}\sigma$ for $Z\sim N(0,\sigma^{2}I_{d})$ and $\kappa_{\alpha,r,d,\sigma}:=e^{\frac{\alpha b_{r,\sigma}(1+r+\sqrt{d}\sigma)}{2}}$. Similarly, using $\mathrm{spt}(\hat{\mu}_{n}),\mathrm{spt}(\hat{\nu}_{n}),\mathrm{spt}(\nu)\subseteq\mathbb{B}_{d}(r)$, we can lower bound the terms in the denominator of the expression in (47), which upon substitution therein yields
$\left|\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|$
$\leq\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}\sup_{f\in\mathcal{F}(b_{r,\sigma})}\left|L_{\alpha-1,\sigma,\hat{\mu}_{n}}(f)-L_{\alpha-1,\sigma,\mu}(f)\right|+\kappa_{\alpha,r,d,\sigma}\sup_{f\in\mathcal{F}(b_{r,\sigma})}\left|L_{\alpha,\sigma,\hat{\nu}_{n}}(f)-L_{\alpha,\sigma,\nu}(f)\right|$
$=\sup_{f\in\bar{\mathcal{E}}_{\sigma,\alpha,r,d}}\left|(\hat{\mu}_{n}-\mu)(f)\right|+\sup_{f\in\tilde{\mathcal{E}}_{\sigma,\alpha,r,d}}\left|(\hat{\nu}_{n}-\nu)(f)\right|.$
Here
$\bar{\mathcal{E}}_{\sigma,\alpha,r,d}$ $:=\left\{\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{|\alpha-1|f}*\varphi_{\sigma}:f\in\mathcal{F}(b_{r,\sigma})\right\},$
$\tilde{\mathcal{E}}_{\sigma,\alpha,r,d}$ $:=\left\{\kappa_{\alpha,r,d,\sigma}e^{\alpha f}*\varphi_{\sigma}:f\in\mathcal{F}(b_{r,\sigma})\right\},$
are Gaussian-smoothed function classes with envelopes $\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}$ and $\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}$, respectively, where
$\bar{E}_{\sigma,\alpha,r,d}$ $:=\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{|\alpha-1|F_{b_{r,\sigma}}},$
$\tilde{E}_{\sigma,\alpha,r,d}$ $:=\kappa_{\alpha,r,d,\sigma}e^{\alpha F_{b_{r,\sigma}}}.$
It follows analogous to (34) that
$J(1,\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\bar{\mathcal{E}}_{\sigma,\alpha,r,d},\mu)\vee J(1,\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\tilde{\mathcal{E}}_{\sigma,\alpha,r,d},\nu)$ $\leq(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}}=:\tilde{\xi}_{r,d,\sigma},$
where $\tilde{\xi}_{r,d,\sigma}$ is defined in (36).
Next, note that
$\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}(x)$ $=\frac{1}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{|\alpha-1|F_{b_{r,\sigma}}(y)}e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
$\leq\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{\frac{|\alpha-1|b_{r,\sigma}(1+\left\|x\right\|)}{2}}\mathbb{E}\left[e^{\frac{|\alpha-1|b_{r,\sigma}\left\|Z\right\|}{2}}\right]$
$\leq\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{\frac{|\alpha-1|b_{r,\sigma}(1+\left\|x\right\|)}{2}+d|\alpha-1|^{2}b_{r,\sigma}^{2}\sigma^{2}},$
where we used the exponential moment bound in Lemma 4. Hence
$\left\|\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{2,\mu}\vee\left\|\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{\infty,\mu}$ $\leq\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{\frac{|\alpha-1|b_{r,\sigma}(1+r)}{2}+d|\alpha-1|^{2}b_{r,\sigma}^{2}\sigma^{2}}$
$\leq\frac{\alpha}{\left|\alpha-1\right|}e^{|\alpha-1|b_{r,\sigma}\left(1+r+\frac{\sqrt{d}\sigma}{2}+d|\alpha-1|b_{r,\sigma}\sigma^{2}\right)}.$
Likewise, we have
$\left\|\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{2,\nu}\vee\left\|\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{\infty,\nu}$ $\leq\kappa_{\alpha,r,d,\sigma}e^{\frac{\alpha b_{r,\sigma}(1+r)}{2}+d\alpha^{2}b_{r,\sigma}^{2}\sigma^{2}}\leq e^{\alpha b_{r,\sigma}\left(1+r+\frac{\sqrt{d}\sigma}{2}+d\alpha b_{r,\sigma}\sigma^{2}\right)}.$
Substituting the above bounds, we obtain for $\sigma\leq 1$ that
$t_{n,\sigma,\mu,\nu}$
$:=J(1,\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\bar{\mathcal{E}}_{\sigma,\alpha,r,d},\mu)\left\|\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{2,\mu}+J(1,\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\tilde{\mathcal{E}}_{\sigma,\alpha,r,d},\nu)\left\|\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{2,\nu}$
$\lesssim(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}}\left(\frac{\alpha}{\left|\alpha-1\right|}e^{|\alpha-1|b_{r,\sigma}\left(1+r+\frac{\sqrt{d}\sigma}{2}+d|\alpha-1|b_{r,\sigma}\sigma^{2}\right)}+e^{\alpha b_{r,\sigma}\left(1+r+\frac{\sqrt{d}\sigma}{2}+d\alpha b_{r,\sigma}\sigma^{2}\right)}\right)$
$\lesssim\tilde{\xi}_{r,d,\sigma}\lambda_{\alpha,d,r,\sigma},$
where
$\lambda_{\alpha,d,r,\sigma}:=\left(\frac{\alpha}{\left|\alpha-1\right|}+1\right)e^{(\alpha+1)b_{r,\sigma}\left(1+r+\frac{\sqrt{d}\sigma}{2}+d(\alpha+1)b_{r,\sigma}\sigma^{2}\right)}.$ (48)
Since
$\left\|\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{2,\mu}\vee\left\|\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{\infty,\mu}\vee\left\|\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{2,\nu}\vee\left\|\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{\infty,\nu}\leq\lambda_{\alpha,d,r,\sigma},$
following similar steps as leading to (26) and substituting the above bounds in the resulting inequality, we obtain (13).
4.4. Proof of Corollary 1
From (28), we have for $0<\sigma\leq 1$ and every $\mu,\nu\in\mathcal{P}\left(\mathbb{B}_{d}(r)\right)$,
$\mathbb{P}\left(\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\gtrsim\bar{c}_{d,r}\sigma^{-\left(\frac{d}{2}+2\right)}n^{-\frac{1}{2}}+\frac{(r^{2}+4r)(1+r\sqrt{d})}{\sigma^{2}}z\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq e^{-\left(nz^{2}\wedge nz\right)},$
where $\bar{c}_{d,r}:=(3ed)^{3(d+2)}\sqrt{d}r(r+1)^{\frac{d}{2}+3}$. On the other hand, (15a) yields
$\left|\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\leq c_{d,s}M\left(M+1+\log M\right)\sigma^{s}\leq 2c_{d,s}M^{2}\sigma^{s}.$
Combining these, and setting $\sigma=n^{-\frac{1}{d+2s+4}}\leq 1$ and $z=n^{\frac{-(s+2)}{d+2s+4}}\bar{z}$, (14) follows via triangle inequality.
4.5. Proof of Proposition 1
The inequality (15a) follows via same steps as in [67, Lemma 3] up to replacing Lebesgue measure by $\rho$ and taking $p_{\mu}$, $p_{\nu}$ as densities w.r.t. $\rho$. We next show (15b) using the same techniques therein. For $\tau\in[0,1]$, let $z_{\mu,\tau}(x):=(1-\tau)p_{\mu}(x)+\tau\mu*\varphi_{\sigma}(x)$ and $z_{\nu,\tau}(x):=(1-\tau)p_{\nu}(x)+\tau\nu*\varphi_{\sigma}(x)$. By Taylor’s theorem,
$\left(\mu*\varphi_{\sigma}(x)\right)^{\alpha}\left(\nu*\varphi_{\sigma}(x)\right)^{1-\alpha}$ $=p_{\mu}(x)^{\alpha}p_{\nu}(x)^{1-\alpha}+\alpha\big{(}\mu*\varphi_{\sigma}(x)-p_{\mu}(x)\big{)}\int_{0}^{1}\bigg{(}\frac{z_{\mu,\tau}(x)}{z_{\nu,\tau}(x)}\bigg{)}^{\alpha-1}d\tau$
$\quad+(1-\alpha)\big{(}\nu*\varphi_{\sigma}(x)-p_{\nu}(x)\big{)}\int_{0}^{1}\bigg{(}\frac{z_{\mu,\tau}(x)}{z_{\nu,\tau}(x)}\bigg{)}^{\alpha}d\tau.$ (49)
Since $1/M\leq p_{\mu}(x)/p_{\nu}(x)\leq M$ $\rho$-a.e. by assumption, we have $1/M\leq\mu*\varphi_{\sigma}(x)/\nu*\varphi_{\sigma}(x)\leq M$ $\rho$-a.e. Hence,
$\frac{1}{M}$ $\leq\frac{p_{\mu}(x)}{p_{\nu}(x)}\wedge\frac{\mu*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}\leq\frac{z_{\mu,\tau}(x)}{z_{\nu,\tau}(x)}\leq\frac{p_{\mu}(x)}{p_{\nu}(x)}\vee\frac{\mu*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}\leq M,~~~\rho~\mbox{a.e.}$ (50)
Consider $\alpha\in(0,1)$. Integrating w.r.t. $\rho$ in (49), we obtain (note that $M\geq 1$)
[…]
$\leq c_{d,s}\big{(}\alpha M^{2-\alpha}+(1-\alpha)M^{\alpha+1}\big{)}\sigma^{s}.$ (51)
In the final inequality, we used that
$\left(\int_{\mathbb{R}^{d}}\left|p_{\mu}-\mu*\varphi_{\sigma}\right|d\rho\right)\vee\left(\int_{\mathbb{R}^{d}}\left|p_{\nu}-\nu*\varphi_{\sigma}\right|d\rho\right)\leq c_{d,s}M\sigma^{s},$
which follows under the Lipschitz assumption on the densities similarly to the proof of [67, Lemma 3]. On the other hand, for $\alpha\in(1,\infty)$, integrating w.r.t. $\rho$ in (49) leads to
[…]
$\leq c_{d,s}\big{(}\alpha M^{\alpha}+(\alpha-1)M^{\alpha+1}\big{)}\sigma^{s}.$ (52)
Next, observe that (50) implies
$e^{(\alpha-1)\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)}=\int p_{\mu}(x)^{\alpha}p_{\nu}(x)^{1-\alpha}d\rho\geq M^{-\alpha},$ (53a)
and similarly
$e^{(\alpha-1)\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)}\geq M^{-\alpha}.$ (53b)
Applying the inequality $\big{|}\log x-\log y\big{|}\leq\left|x-y\right|/(x\wedge y)$ for $x,y>0$ with $x=e^{(\alpha-1)\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)}$, $y=e^{(\alpha-1)\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)}$ and using (51)-(53) yields the desired claim.
4.6. Proof of Theorem 4
Part $(i)$: Using (16a) and (46),
$\left|\hat{\mathsf{D}}_{\mathsf{KL},\mathcal{G}_{b,\beta,\rho}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\hat{\mathsf{D}}_{\mathsf{KL},\mathcal{G}_{b,\beta,\rho}}\left(\mu,\nu\right)\right|$ $\leq\sup_{g\in\mathcal{G}_{b,\beta,\rho}}\left|(\hat{\mu}_{n}-\mu)(g)-(\hat{\nu}_{n}-\nu)(e^{g})\right|$
$\leq\sup_{g\in\mathcal{G}_{b,\beta,\rho}}\left|(\hat{\mu}_{n}-\mu)(g)\right|+\sup_{g\in\mathcal{G}_{b,\beta,\rho}}\left|(\hat{\nu}_{n}-\nu)(e^{g})\right|$
$=\frac{Z_{n,\mathcal{G}_{b,\beta,\rho}}+\tilde{Z}_{n,\mathcal{E}_{b,\beta,\rho}}}{n},$ (54)
where $\mathcal{G}_{b,\beta,\rho}$ is as given in Theorem 4, $\mathcal{E}_{b,\beta,\rho}=\{e^{g}:g\in\mathcal{G}_{b,\beta,\rho}\}$,
$Z_{n,\mathcal{G}_{b,\beta,\rho}}:=\sup_{g\in\mathcal{G}_{b,\beta,\rho}}\left|\sum_{i=1}^{n}\big{(}g(X_{i})-\mathbb{E}_{\mu}[g]\big{)}\right|\quad\mbox{ and }\quad\tilde{Z}_{n,\mathcal{E}_{b,\beta,\rho}}:=\sup_{g\in\mathcal{E}_{b,\beta,\rho}}\left|\sum_{i=1}^{n}\big{(}g(Y_{i})-\mathbb{E}_{\nu}[g]\big{)}\right|.$ (55)
To obtain the desired deviation inequality, we will use Theorem 6. To this end, we require an upper bound $\mathbb{E}[Z_{n,\mathcal{G}_{b,\beta,\rho}}]$ and $\mathbb{E}[\tilde{Z}_{n,\mathcal{E}_{b,\beta,\rho}}]$, for which we use the result stated next.
Theorem 5 (see e.g., Theorem 2.14.1 in [75]).
Suppose $\mathcal{F}$ satisfies Assumption 1 with envelope $F\in L^{p}(\mu)$ for some $p\geq 2$. For $X^{n}\sim\mu^{\otimes n}$, let $Z_{n}$ be as given in (11a) (with $\mathcal{F}_{n}=\mathcal{F}$ for all $n\in\mathbb{N}$). Then
$\mathbb{E}\left[Z_{n}\right]\leq\left(\mathbb{E}[Z_{n}^{p}]\right)^{\frac{1}{p}}\leq c_{p}\sqrt{n}J(1,F,\mathcal{F},\mu)\left\|F\right\|_{p,\mu},$ (56)
where $c_{p}>0$ is a constant.
Recall that $\mu,\nu\ll\rho$, $b=\log M$, $\beta>d/2$ and $a:=\frac{d}{2\beta}<1$. From the definition of $J(1,F,\mathcal{F},\mu)$ (see (22)) with $\mathcal{F}=\mathcal{G}_{b,\beta,\rho}$ and envelope $F$ such that $F=b$ $\rho$-a.e., we have using (17) that
$J(1,b,\mathcal{G}_{b,\beta,\rho},\mu)$ $:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon b,\mathcal{G}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon\leq\sqrt{\hat{c}_{d,\beta}}\int_{0}^{1}\epsilon^{-\frac{d}{2\beta}}d\epsilon\leq\frac{\sqrt{\hat{c}_{d,\beta}}}{1-a}.$
It follows from (56) that
[…]
Observe that for $z_{1},z_{2}\in\mathbb{R}$ such that $\left|z_{1}\right|\vee\left|z_{2}\right|\leq b$, we have $\left|e^{z_{1}}-e^{z_{2}}\right|\leq e^{b}\left|z_{1}-z_{2}\right|$. Since this implies that for $g_{1},g_{2}\in\mathcal{G}_{b,\beta,\rho}$, $\left\|e^{g_{1}}-e^{g_{2}}\right\|_{2,\gamma}\leq e^{b}\left\|g_{1}-g_{2}\right\|_{2,\gamma}$ for any probability measure $\gamma$ with $\mathrm{spt}(\gamma)\subseteq\mathrm{spt}(\rho)$, we obtain
$N\big{(}\epsilon e^{b},\mathcal{E}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\big{)}\leq N\big{(}\epsilon,\mathcal{G}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\big{)}.$
Consequently
$J(1,e^{b},\mathcal{E}_{b,\beta,\rho},\nu)$ $:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon e^{b},\mathcal{E}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon$
$\leq\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon,\mathcal{G}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon$
$\leq\frac{\sqrt{\hat{c}_{d,\beta}}b^{a}}{1-a},$
where we again used (17) to obtain the last inequality. Then, (56) yields
$\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{b,\beta,\rho}}\right]\lesssim\sqrt{n}J\big{(}1,e^{b},\mathcal{E}_{b,\beta,\rho},\nu\big{)}e^{b}\leq\frac{\sqrt{n\hat{c}_{d,\beta}}b^{a}e^{b}}{1-a}.$
In order to apply (62b), we also use the following bounds:
$\sqrt{V_{n}}$ $:=\sqrt{n\theta^{2}+2b\mathbb{E}\left[Z_{n,\mathcal{G}_{b,\beta,\rho}}\right]}\leq\sqrt{n}\theta+\sqrt{2b\mathbb{E}\left[Z_{n,\mathcal{G}_{b,\beta,\rho}}\right]}\lesssim\sqrt{n}b+\sqrt{2}b(\hat{c}_{d,\beta})^{\frac{1}{4}}(1-a)^{-\frac{1}{2}}n^{\frac{1}{4}},$
$\sqrt{\tilde{V}_{n}}$ $:=\sqrt{n\tilde{\theta}^{2}+2e^{b}\mathbb{E}[\tilde{Z}_{n,\mathcal{E}_{b,\beta,\rho}}]}\leq\sqrt{n}\tilde{\theta}+\sqrt{2e^{b}\mathbb{E}[\tilde{Z}_{n,\mathcal{E}_{b,\beta,\rho}}]}\leq\sqrt{n}e^{b}+\sqrt{2}e^{b}b^{\frac{a}{2}}(\hat{c}_{d,\beta})^{\frac{1}{4}}(1-a)^{-\frac{1}{2}}n^{\frac{1}{4}}.$
Setting $c_{d,\beta}:=c\frac{\sqrt{\hat{c}_{d,\beta}}}{1-a}$ for an appropriate constant $c\geq 1$ and noting that the above bounds yields $\sqrt{V_{n}}\leq\sqrt{c_{d,\beta}n}b$ and $\sqrt{\tilde{V}_{n}}\leq\sqrt{c_{d,\beta}n}e^{b}(1+b^{\frac{a}{2}})$,
$\mathbb{P}\left(\left|\hat{\mathsf{D}}_{\mathsf{KL},\mathcal{G}_{b,\beta,\rho}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\hat{\mathsf{D}}_{\mathsf{KL},\mathcal{G}_{b,\beta,\rho}}\left(\mu,\nu\right)\right|\geq c_{d,\beta}(b+b^{a}e^{b})n^{-\frac{1}{2}}+4\sqrt{2c_{d,\beta}}e^{b}\big{(}1+b^{\frac{a}{2}}\big{)}z\right)$
$\leq\mathbb{P}\left(Z_{n,\mathcal{G}_{b,\beta,\rho}}+\tilde{Z}_{n,\mathcal{E}_{b,\beta,\rho}}\geq c_{d,\beta}(b+b^{a}e^{b})\sqrt{n}+4\sqrt{2c_{d,\beta}}e^{b}\big{(}1+b^{\frac{a}{2}}\big{)}nz\right)$
$\leq\mathbb{P}\left(Z_{n,\mathcal{G}_{b,\beta,\rho}}-\mathbb{E}\big{[}Z_{n,\mathcal{G}_{b,\beta,\rho}}\big{]}\geq 2\sqrt{2c_{d,\beta}}bnz\right)+\mathbb{P}\left(\tilde{Z}_{n,\mathcal{E}_{b,\beta,\rho}}-\mathbb{E}\big{[}\tilde{Z}_{n,\mathcal{E}_{b,\beta,\rho}}\big{]}\geq 2\sqrt{2c_{d,\beta}}e^{b}\big{(}1+b^{\frac{a}{2}}\big{)}nz\right)$
$\leq 2e^{-(nz^{2}\wedge nz)}.$ (57)
Here, the first inequality follows from (54) while the final step is via an application of (62b).
Next, observe that for $(\mu,\nu)\in\bar{\mathcal{P}}_{M,\rho}(\mathcal{G}_{b,\beta,\rho},\delta)$, there exists $f\in\mathcal{G}_{b,\beta,\rho}$ (by definition) such that $\left\|f_{\mu,\nu}^{\star}-f\right\|_{\infty,\rho}\leq\delta$. This implies
$\left|\hat{\mathsf{D}}_{\mathsf{KL},\mathcal{G}_{b,\beta,\rho}}\left(\mu,\nu\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right)\right|$ $=\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right)-\hat{\mathsf{D}}_{\mathsf{KL},\mathcal{G}_{b,\beta,\rho}}\left(\mu,\nu\right)$
$=\mathbb{E}_{\mu}[f_{\mu,\nu}^{\star}]-\mathbb{E}_{\nu}[e^{f_{\mu,\nu}^{\star}}]+1-\hat{\mathsf{D}}_{\mathsf{KL},\mathcal{G}_{b,\beta,\rho}}\left(\mu,\nu\right)$
$\leq\mathbb{E}_{\mu}[f_{\mu,\nu}^{\star}-f]-\mathbb{E}_{\nu}[e^{f_{\mu,\nu}^{\star}}-e^{f}]$
$\leq\mathbb{E}_{\mu}\big{[}\left|f_{\mu,\nu}^{\star}-f\right|\big{]}+M\mathbb{E}_{\nu}\left[\left|f_{\mu,\nu}^{\star}-f\right|\right]$
$\leq(M+1)\delta,$
where in the penultimate inequality, we used $\left\|f\right\|_{\infty,\rho}\leq b=\log M$ and the fact that the Lipschitz constant of $e^{x}$ is bounded by $M$ for $\left|x\right|\leq\log M$. Hence, using triangle inequality and adjusting the constant $c_{d,\beta}$, we have
$\mathbb{P}\left(\left|\hat{\mathsf{D}}_{\mathsf{KL},\mathcal{G}_{b,\beta,\rho}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu\middle\|\nu\right)\right|\geq(M+1)\delta+c_{d,\beta}M\left(1+(\log M)^{a}\right)\big{(}n^{-\frac{1}{2}}+z\big{)}\right)\leq 2e^{-(nz^{2}\wedge nz)}.$
Taking supremum w.r.t. $(\mu,\nu)\in\bar{\mathcal{P}}_{M,\rho}(\mathcal{G}_{b,\beta,\rho},\delta)$ completes the proof of (18).
Part $(ii)$: We have
$\left|\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{b,\beta,\rho}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{b,\beta,\rho}}\left(\mu,\nu\right)\right|$
$\leq\sup_{g\in\mathcal{G}_{b,\beta,\rho}}\frac{\alpha}{\left|\alpha-1\right|}\left|\log\left(\frac{\mathbb{E}_{\hat{\mu}_{n}}\big{[}e^{(\alpha-1)g}\big{]}}{\mathbb{E}_{\mu}\big{[}e^{(\alpha-1)g}\big{]}}\right)\right|+\sup_{g\in\mathcal{G}_{b,\beta,\rho}}\left|\log\left(\frac{\mathbb{E}_{\hat{\nu}_{n}}\big{[}e^{\alpha g}\big{]}}{\mathbb{E}_{\nu}\big{[}e^{\alpha g}\big{]}}\right)\right|$
$\leq\sup_{g\in\mathcal{G}_{b,\beta,\rho}}\frac{\alpha e^{\left|\alpha-1\right|b}}{\left|\alpha-1\right|}\left|\mathbb{E}_{\hat{\mu}_{n}}\big{[}e^{(\alpha-1)g}\big{]}-\mathbb{E}_{\mu}\big{[}e^{(\alpha-1)g}\big{]}\right|+e^{\alpha b}\sup_{g\in\mathcal{G}_{b,\beta,\rho}}\left|\mathbb{E}_{\hat{\nu}_{n}}\big{[}e^{\alpha g}\big{]}-\mathbb{E}_{\nu}\big{[}e^{\alpha g}\big{]}\right|$
$=\sup_{g\in\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}}\left|(\hat{\mu}_{n}-\mu)(g)\right|+\sup_{g\in\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho}}\left|(\hat{\nu}_{n}-\nu)(g)\right|$
$=\frac{Z_{n,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}}+\tilde{Z}_{n,\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho}}}{n},$
where $Z_{n,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}}$ and $\tilde{Z}_{n,\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho}}$ are as defined in (55) with
$\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}$ $:=\left\{\frac{\alpha e^{\left|\alpha-1\right|b}}{\left|\alpha-1\right|}e^{(\alpha-1)g}:g\in\mathcal{G}_{b,\beta,\rho}\right\},$ (58a)
$\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho}$ $:=\big{\{}e^{\alpha(b+g)}:g\in\mathcal{G}_{b,\beta,\rho}\big{\}}.$ (58b)
In the first inequality above, we used (16b) and (46). In the second inequality, we used $\left|\log\frac{x}{y}\right|\leq\frac{|x-y|}{x\wedge y}$ and that $\left\|g\right\|_{\infty,\rho}\leq b=\log M$ for $g\in\mathcal{G}_{b,\beta,\rho}$. Note that for any $c\in\mathbb{R}$ and $x,y\in\mathbb{R}$ such that $\left|x\right|\vee\left|y\right|\leq b$, $\left|e^{cx}-e^{cy}\right|\leq\left|c\right|e^{\left|c\right|b}\left|x-y\right|$. Hence, for any probability measure $\gamma$ such that $\mathrm{spt}(\gamma)\subseteq\mathrm{spt}(\rho)$,
$N\Big{(}\alpha e^{2\left|\alpha-1\right|b}\epsilon,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}\leq N\Big{(}\epsilon,\mathcal{G}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)},$ (59a)
$N\Big{(}\alpha e^{2\alpha b}\epsilon,\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}\leq N\Big{(}\epsilon,\mathcal{G}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}.$ (59b)
Let $F$ be an envelope for $\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}$ such that $F=(\alpha/\left|\alpha-1\right|)e^{2\left|\alpha-1\right|b}$ $\rho$-a.e. Then, we have using (17) that for $\beta>d/2$,
$J(1,F,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho},\mu)$ $:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon F,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon\leq\frac{\sqrt{\hat{c}_{d,\beta}}}{1-a}\big{(}b\left|\alpha-1\right|\big{)}^{a}.$
Similarly, considering $F$ such that $F=e^{2\alpha b}$ $\rho$-a.e. to be an envelope for $\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho}$, we obtain for $\beta>d/2$ that
$J(1,F,\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho},\nu)$ $:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon F,\tilde{\mathcal{G}}_{b,\alpha,\beta},\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon\leq\frac{\sqrt{\hat{c}_{d,\beta}}}{1-a}(b\alpha)^{a}.$
Consequently, (56) yields (with respective envelopes $F$) that
$\mathbb{E}\left[Z_{n,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}}\right]$ $\lesssim\sqrt{n}J\big{(}1,F,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho},\mu\big{)}\left\|F\right\|_{2,\mu}\leq\frac{\sqrt{\hat{c}_{d,\beta}}}{1-a}\sqrt{n}\alpha b^{a}\left|\alpha-1\right|^{a-1}e^{2\left|\alpha-1\right|b},$
$\mathbb{E}\left[\tilde{Z}_{n,\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho}}\right]$ $\lesssim\sqrt{n}J\big{(}1,F,\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho},\nu\big{)}\left\|F\right\|_{2,\mu}\leq\frac{\sqrt{\hat{c}_{d,\beta}}}{1-a}\sqrt{n}(b\alpha)^{a}e^{2\alpha b}.$
Moreover
$\sqrt{V_{n}}$ $\leq\sqrt{n}\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}+\sqrt{2\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}\mathbb{E}\big{[}Z_{n,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}}\big{]}}$
$\leq\sqrt{n}\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}+\frac{\sqrt{2c_{d,\beta}}\alpha e^{2\left|\alpha-1\right|b}b^{\frac{a}{2}}n^{\frac{1}{4}}}{\left|\alpha-1\right|^{1-\frac{a}{2}}}$
$\leq\frac{\sqrt{nc_{d,\beta}}\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}\left(1+(b\left|\alpha-1\right|)^{\frac{a}{2}}\right),$
and
$\sqrt{\tilde{V}_{n}}\leq\sqrt{n}e^{2\alpha b}+\sqrt{2b\mathbb{E}\big{[}\tilde{Z}_{n,\tilde{\mathcal{G}}_{b,\alpha,\beta}}\big{]}}\leq\sqrt{n}e^{2\alpha b}+\sqrt{2c_{d,\beta}}e^{2\alpha b}(b\alpha)^{\frac{a}{2}}n^{\frac{1}{4}}\leq\sqrt{nc_{d,\beta}}e^{2\alpha b}\big{(}1+(b\alpha)^{\frac{a}{2}}\big{)}.$
Applying (62b), it follows similar to (57) that for $z\geq 0$:
$\mathbb{P}\left(\left|\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{b,\beta,\rho}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{b,\beta,\rho}}\left(\mu,\nu\right)\right|\geq c_{d,\beta}\big{(}n^{-\frac{1}{2}}+z\big{)}v_{e^{b},\alpha,a}\right)\leq 2e^{-(nz^{2}\wedge nz)},$ (60)
where $v_{e^{b},\alpha,a}:=\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}\left(1+(b\left|\alpha-1\right|)^{\frac{a}{2}}\right)+e^{2\alpha b}\big{(}1+(b\alpha)^{a}\big{)}$.
Next, for $(\mu,\nu)\in\bar{\mathcal{P}}_{M,\rho}(\mathcal{G}_{b,\beta,\rho},\delta)$, there exists an $f\in\mathcal{G}_{b,\beta,\rho}$ such that $\left\|f_{\mu,\nu}^{\star}-f\right\|_{\infty,\rho}\leq\delta$. This implies
$\left|\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{b,\beta,\rho}}\left(\mu,\nu\right)-\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)\right|$
$=\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)-\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{b,\beta,\rho}}\left(\mu,\nu\right)$
$\leq\frac{\alpha}{\alpha-1}\left(\log\mathbb{E}_{\hat{\mu}}\big{[}e^{(\alpha-1)f_{\mu,\nu}^{\star}}\big{]}-\log\mathbb{E}_{\hat{\mu}}\big{[}e^{(\alpha-1)f}\big{]}\right)-\left(\log\mathbb{E}_{\nu}\big{[}e^{\alpha f_{\mu,\nu}^{\star}}\big{]}-\log\mathbb{E}_{\nu}\big{[}e^{\alpha f}\big{]}\right)$
$\leq\alpha e^{2\left|\alpha-1\right|b}\mathbb{E}_{\mu}\big{[}\left|f_{\mu,\nu}^{\star}-f\right|\big{]}+\alpha e^{2\alpha b}\mathbb{E}_{\nu}\left[\left|f_{\mu,\nu}^{\star}-f\right|\right]$
$\leq\alpha(M^{2\left|\alpha-1\right|}+M^{2\alpha})\delta,$
where in the penultimate inequality, we used that the Lipschitz constant of $\log x$ is bounded by $1/\left|c\right|$ for $x\geq c>0$ and that of $e^{x}$ is bounded by $e^{c}$ for $x\in[-c,c]$, and in the last equation, we used that $\left\|f_{\mu,\nu}^{\star}-f\right\|_{\infty,\rho}\leq\delta$ and $b=\log M$. The claim in (19) follows via triangle inequality from (60) and the above equation after simplification.
4.7. Proof of Proposition 2
Let $0\leq\tau\leq 1$. Consider the test statistic $T_{n}=\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{n}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)$ with critical value
$t_{n}=\epsilon+\big{(}M^{2\alpha}+M^{2\left|\alpha-1\right|}\big{)}\delta_{n}+c_{d,\beta}v_{M,\alpha,a}\big{(}n^{-\frac{1}{2}}+n^{0.5(\tau-1)}\big{)}.$
Define the event $E_{n}:=\left\{\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{n}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)>t_{n}\right\}$ and let $\bar{E}_{n}$ denote its compliment. Since $\mathsf{D}_{\alpha}\left(\mu_{0}\middle\|\nu_{0}\right)\leq\epsilon$ by assumption of the hypotheses, we have
$e_{1,n}(T_{n},t_{n})$
$=\mathbb{P}\big{(}E_{n}|H_{0}\big{)}$
$=\mathbb{P}\left(\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{n}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)>\epsilon+\big{(}M^{2\alpha}+M^{2\left|\alpha-1\right|}\big{)}\delta_{n}+c_{d,\beta}v_{M,\alpha,a}\big{(}n^{-\frac{1}{2}}+n^{0.5(\tau-1)}\big{)}\big{|}H_{0}\right)$
$\leq\mathbb{P}\left(\left|\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{n}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\mathsf{D}_{\alpha}\left(\mu_{0}\middle\|\nu_{0}\right)\right|\geq\big{(}M^{2\alpha}+M^{2\left|\alpha-1\right|}\big{)}\delta_{n}+c_{d,\beta}v_{M,\alpha,a}\big{(}n^{-\frac{1}{2}}+n^{0.5(\tau-1)}\big{)}\big{|}H_{0}\right)$
$\leq 2e^{-\left(n^{\tau}\wedge n^{0.5(\tau+1)}\right)},$
where the last inequality follows using (19). Note that $n^{\tau}\wedge n^{0.5(\tau+1)}=n^{\tau}$ for $0\leq\tau\leq 1$.
On the other hand, since $\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)>\epsilon$ by assumption of the hypotheses,
$e_{2,n}(T_{n},t_{n})$
$=\mathbb{P}\left(\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{n}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)\leq t_{n}\big{|}H_{1}\right)$
$\leq\mathbb{P}\left(\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{n}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)\leq-(\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)-t_{n})\big{|}H_{1}\right)$
$\leq\mathbb{P}\left(\left|\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{n}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)\right|\geq\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)-t_{n}\big{|}H_{1}\right)$
$=\mathbb{P}\left(\left|\hat{\mathsf{D}}_{\alpha,\mathcal{G}_{n}}\left(\hat{\mu}_{n},\hat{\nu}_{n}\right)-\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)\right|\geq\big{(}M^{2\alpha}+M^{2\left|\alpha-1\right|}\big{)}\delta_{n}+c_{d,\beta}v_{M,\alpha,a}\big{(}n^{-\frac{1}{2}}+\bar{\vartheta}_{n,\alpha,\delta_{n}}(\tau)\big{)}\big{|}H_{1}\right)$
$\leq 2e^{-\vartheta_{n,\alpha,\delta_{n}}(\tau)},$
where in the final inequality, we applied (19) with $z=\bar{\vartheta}_{n,\alpha,\delta_{n}}(\tau)$, and
$\vartheta_{n,\alpha,\delta_{n}}(\tau)$ $:=n\big{(}\bar{\vartheta}_{n,\alpha,\delta_{n}}(\tau)\vee 0\big{)}^{2}\wedge n\big{(}\bar{\vartheta}_{n,\alpha,\delta_{n}}(\tau)\vee 0),$ (61a)
$\bar{\vartheta}_{n,\alpha,\delta_{n}}(\tau)$ $:=\frac{1}{c_{d,\beta}v_{M,\alpha,a}}\left(\mathsf{D}_{\alpha}\left(\mu_{1}\middle\|\nu_{1}\right)-\epsilon-2\big{(}M^{2\alpha}+M^{2\left|\alpha-1\right|}\big{)}\delta_{n}-2c_{d,\beta}v_{M,\alpha,a}n^{-\frac{1}{2}}\right)-n^{0.5(\tau-1)}.$ (61b)
5. Proofs of Lemmas
5.1. Proof of Lemma 2
To handle the case of compactly supported distributions, we will use a version of Talagrand’s concentration inequality given in [12, Theorem 2.3].
Theorem 6 (Theorem 2.3 in [12]).
Suppose that $\mathcal{F}$ satisfies Assumption 1 and $\left\|f\right\|_{\infty,\mu}\leq b$ for all $f\in\mathcal{F}$. For $X^{n}\sim\mu^{\otimes n}$, let $Z_{n}$ be as given in (11a) (with $\mathcal{F}_{n}=\mathcal{F}$ for all $n\in\mathbb{N}$). Let $\theta^{2}:=\sup_{f\in\mathcal{F}}\mathbb{E}_{\mu}\big{[}\big{(}f-\mathbb{E}_{\mu}[f]\big{)}^{2}\big{]}$ and $V_{n}:=n\theta^{2}+2b\mathbb{E}[Z_{n}]$. Then, for all $z\geq 0$,
$\mathbb{P}\left(Z_{n}\geq\mathbb{E}[Z_{n}]+\sqrt{2zV_{n}}+\frac{bz}{3}\right)\leq e^{-z}.$ (62a)
In particular, for every $z\geq 0$,
$\mathbb{P}\left(Z_{n}\geq\mathbb{E}[Z_{n}]+z\right)\leq e^{-\big{(}\frac{z^{2}}{8V_{n}}\wedge\frac{3z}{2b}\big{)}},$ (62b)
$\mathbb{P}\left(Z_{n}\geq 2\mathbb{E}[Z_{n}]+z\right)\leq e^{-\big{(}\frac{z^{2}}{8n\theta^{2}}\wedge\frac{3z}{8b}\big{)}}.$ (62c)
Equation (62b) follows from (62a) by considering the cases $\sqrt{2zV_{n}}\geq bz/3$ and its opposite. Equation (62c) follows from (62a) by noting that $\sqrt{2zV_{n}}\leq\theta\sqrt{2nz}+2\sqrt{bz\mathbb{E}[Z_{n}]}\leq\theta\sqrt{2nz}+bz+\mathbb{E}[Z_{n}]$, where we used $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ for $x,y\geq 0$ and $2xy\leq x^{2}+y^{2}$ for $x,y\in\mathbb{R}$, and considering the cases $\sqrt{2zn\theta^{2}}\geq 4bz/3$ and its opposite.
To show (24), we will use the following Fuk-Nagaev type inequality which is applicable for a sequence of function classes $(\mathcal{F}_{n})_{n\in\mathbb{N}}$ with unbounded envelope. It is a restatement of [2, Theorem 4] in a form with explicit constants which is useful for our purposes.
Theorem 7 (Theorem 4 in [2]).
In the setting of Lemma 1, for $c_{q}$ as specified in (83),
$\mathbb{P}\Big{(}Z_{n}\geq 2\mathbb{E}[Z_{n}]+z\Big{)}\leq e^{-\big{(}\frac{z^{2}}{128n\left\|F_{n}\right\|_{2,\mu}^{2}}\wedge\frac{3z}{256\bar{M}_{n}}\big{)}}+3e^{-\left(\frac{z}{c_{q}\left\|M_{n}\right\|_{\psi_{q}}}\right)^{q}}.$ (63)
The constants in Theorem 7 are determined similar to that in Lemma 1.
Proceeding with the proof of (24), let $X^{n}\sim\mu^{\otimes n}$, $Y^{n}\sim\nu^{\otimes n}$, and $\mathcal{F}_{n}$ be the function class having the properties specified in Lemma 2. Recall that $f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}:=\log\left(\mu*\varphi_{\sigma}/\nu*\varphi_{\sigma}\right)$. Since $f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star},f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}\in\mathcal{F}_{n}$ by assumption,
$\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\hat{\mu}_{n}*\gamma_{\sigma}}[f]-\mathbb{E}_{\hat{\nu}_{n}*\gamma_{\sigma}}[e^{f}]+1$ $\leq\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)$
$=\mathbb{E}_{\hat{\mu}_{n}*\gamma_{\sigma}}\big{[}f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star}\big{]}-\mathbb{E}_{\hat{\nu}_{n}*\gamma_{\sigma}}\big{[}e^{f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star}}\big{]}+1,$
$\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\mu*\gamma_{\sigma}}[f]-\mathbb{E}_{\nu*\gamma_{\sigma}}[e^{f}]+1$ $\leq\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)$
$=\mathbb{E}_{\mu*\gamma_{\sigma}}\big{[}f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}\big{]}-\mathbb{E}_{\nu*\gamma_{\sigma}}\big{[}e^{f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}}\big{]}+1.$
Hence
$\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)=\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\hat{\mu}_{n}*\gamma_{\sigma}}[f]-\mathbb{E}_{\hat{\nu}_{n}*\gamma_{\sigma}}[e^{f}]+1,$ (64a)
$\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)=\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\mu*\gamma_{\sigma}}[f]-\mathbb{E}_{\nu*\gamma_{\sigma}}[e^{f}]+1.$ (64b)
Moreover, since $\left|f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}(x)\right|\leq F_{n}(x)$ for all $x\in\mathbb{R}^{d}$ and $\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}<\infty$ for some $q\in(0,1]$ implies that $\left\|F_{n}\right\|_{1,\mu*\gamma_{\sigma}}=\left\|F_{n}*\varphi_{\sigma}\right\|_{1,\mu}<\infty$ by (103) given below, $\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)<\infty$. Then, from (46) and (64), we obtain
$\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}$
$\leq\sup_{f\in\mathcal{F}_{n}}\left|(\hat{\mu}_{n}*\varphi_{\sigma}-\mu*\varphi_{\sigma})(f)-(\hat{\nu}_{n}*\varphi_{\sigma}-\nu*\varphi_{\sigma})(e^{f})\right|$
$\leq\sup_{f\in\mathcal{F}_{n}}\left|(\hat{\mu}_{n}*\varphi_{\sigma}-\mu*\varphi_{\sigma})(f)\right|+\sup_{f\in\mathcal{F}_{n}}\left|(\hat{\nu}_{n}*\varphi_{\sigma}-\nu*\varphi_{\sigma})(e^{f})\right|$
$=\frac{Z_{n,\mathcal{F}_{n,\sigma}}+\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}}{n},$
where
$Z_{n,\mathcal{F}_{n,\sigma}}:=\sup_{f\in\mathcal{F}_{n,\sigma}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\mu}[f]\big{)}\right|\quad\mbox{ and }\quad\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}:=\sup_{f\in\mathcal{E}_{n,\sigma}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\nu}[f]\big{)}\right|,$
and $\mathcal{F}_{n,\sigma}$ and $\mathcal{E}_{n,\sigma}$ are as defined in Lemma 2. By an application of Theorem 5, we obtain
$\mathbb{E}[Z_{n,\mathcal{F}_{n,\sigma}}]$ $\leq c\sqrt{n}J(1,F_{n}*\varphi_{\sigma},\mathcal{F}_{n,\sigma},\mu)\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu},$ (65a)
$\mathbb{E}[\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}]$ $\leq c\sqrt{n}J(1,e^{F_{n}}*\varphi_{\sigma},\mathcal{E}_{n,\sigma},\nu)\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{2,\nu}.$ (65b)
Now, we can upper bound the probability of deviation as follows:
$\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\geq 2n^{-\frac{1}{2}}t_{n,\sigma,\mu,\nu}+z\right)$
$\leq\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\geq 2n^{-1}\mathbb{E}\left[Z_{n,\mathcal{F}_{n,\sigma}}\right]+2n^{-1}\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\right]+z\right)$
$\leq\mathbb{P}\left(Z_{n,\mathcal{F}_{n,\sigma}}+\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\geq 2\mathbb{E}\left[Z_{n,\mathcal{F}_{n,\sigma}}\right]+2\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\right]+nz\right)$
$\leq\mathbb{P}\left(Z_{n,\mathcal{F}_{n,\sigma}}\geq 2\mathbb{E}\left[Z_{n,\mathcal{F}_{n,\sigma}}\right]+\frac{nz}{2}\right)+\mathbb{P}\left(\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\geq 2\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\right]+\frac{nz}{2}\right).$ (66)
Applying (63) to each term above and scaling $z$ by appropriate constants (depending on $q$), we obtain (24).
To show (25f)-(25g) and thus complete the proof, we will use the following maximal inequality upper bounding the Orlicz norm of maximum of finitely many random variables.
Lemma 6 (Orlicz norm of maxima).
For any $0<q<\infty$ and $\psi_{q}(x)=e^{x^{q}}-1$, we have
$\left\|\max_{1\leq i\leq n}X_{i}\right\|_{\psi_{q}}\leq\big{(}5\log(1+2n)\big{)}^{\frac{1}{q}}\max_{1\leq i\leq n}\left\|X_{i}\right\|_{\psi_{q}}.$ (67)
The proof of (67) is given in Appendix C.2, which is similar in spirit to the proof of [75, Lemma 2.2.2.], albeit without invoking convexity of $\psi_{q}$ (required here since $\psi_{q}=e^{x^{q}}-1$ is not convex near origin for $0<q<1$).
Continuing with the proof of Lemma 2, it follows from (67) that
$\left\|M_{n,\mathcal{F}_{n,\sigma}}\right\|_{\psi_{q}}$ $\leq\big{(}5\log(2n+1)\big{)}^{\frac{1}{q}}\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}},$
$\left\|\tilde{M}_{n,\mathcal{E}_{n,\sigma}}\right\|_{\psi_{q}}$ $\leq\big{(}5\log(2n+1)\big{)}^{\frac{1}{q}}\left\|e^{F_{n}}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}},$
where RHS of the equations above are finite by the assumption $\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}\vee\left\|e^{F_{n}}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}}$ $\infty$ for some $q\in(0,1]$. Further, by using (103) followed by (67), we obtain
[…]
$\check{M}_{n,\mathcal{E}_{n,\sigma}}\leq\left\lceil\frac{1}{q}\right\rceil!\left\|\tilde{M}_{n,\mathcal{E}_{n,\sigma}}\right\|_{\psi_{q}}\leq\left\lceil\frac{1}{q}\right\rceil!\big{(}5\log(2n+1)\big{)}^{\frac{1}{q}}\left\|e^{F_{n}}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}}.$
The proof of (26) is analogous up to (66), and the conclusion then follows by applying (62c) instead and scaling $z$ appropriately.
5.2. Proof of Lemma 3
We may assume without loss of generality that $b>0$, since otherwise, the claims hold trivially. Consider $q=0$ or $q\in[1,\infty)$. For $f\in\mathcal{F}(b,q)$, set $f_{\sigma}:=f*\varphi_{\sigma}$. Let $H_{m}(z)$, $m\in\mathbb{Z}_{\geq 0}$, denote the Hermite polynomial of degree $m$ defined by
$H_{m}(z)=(-1)^{m}e^{\frac{z^{2}}{2}}\left[\frac{d^{m}}{dz^{m}}e^{-\frac{z^{2}}{2}}\right].$
We have
$D^{k}f_{\sigma}(x)=(-\sigma)^{-|k|}\int f(y)\varphi_{\sigma}(x-y)\underbrace{\left[\prod_{j=1}^{d}H_{k_{j}}\left(\frac{x_{j}-y_{j}}{\sigma}\right)\right]}_{B_{k,\sigma}(x-y)}dy.$ (68)
Hence
$\left|D^{k}f_{\sigma}(x)-D^{k}f_{\sigma}(z)\right|\leq\sigma^{-|k|}\int\left|f(y)\right|\left|\varphi_{\sigma}(x-y)B_{k,\sigma}(x-y)-\varphi_{\sigma}(z-y)B_{k,\sigma}(z-y)\right|dy.$ (69)
By triangle inequality for absolute value, we have
$\left|\varphi_{\sigma}(x-y)B_{k,\sigma}(x-y)-\varphi_{\sigma}(z-y)B_{k,\sigma}(z-y)\right|$
$\leq\varphi_{\sigma}(x-y)\left|B_{k,\sigma}(x-y)-B_{k,\sigma}(z-y)\right|+\left|B_{k,\sigma}(z-y)\right|\left|\varphi_{\sigma}(x-y)-\varphi_{\sigma}(z-y)\right|$
$\leq\left|B_{k,\sigma}(x-y)-B_{k,\sigma}(z-y)\right|\big{(}\varphi_{\sigma}(x-y)+\left|\varphi_{\sigma}(x-y)-\varphi_{\sigma}(z-y)\right|\big{)}$
$\qquad\qquad+\left|B_{k,\sigma}(x-y)\right|\left|\varphi_{\sigma}(x-y)-\varphi_{\sigma}(z-y)\right|.$ (70)
In the second inequality, we used $\left|B_{k,\sigma}(z-y)\right|\leq\left|B_{k,\sigma}(x-y)\right|+\left|B_{k,\sigma}(x-y)-B_{k,\sigma}(z-y)\right|$.
Next observe that
$\varphi_{\sigma}(x-y)-\varphi_{\sigma}(z-y)$ $=\varphi_{\sigma}(x-y)\left(1-e^{-\frac{\big{(}\left\|z-y\right\|^{2}-\left\|x-y\right\|^{2}\big{)}}{2\sigma^{2}}}\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\varphi_{\sigma}(x-y)\left(1-e^{-\frac{\big{(}\left\|z-y\right\|+\left\|x-y\right\|\big{)}\left\|x-z\right\|}{2\sigma^{2}}}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{\left\|x-z\right\|}{2\sigma^{2}}\varphi_{\sigma}(x-y)\big{(}\left\|z-y\right\|+\left\|x-y\right\|\big{)}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{\left\|x-z\right\|^{2}}{2\sigma^{2}}\varphi_{\sigma}(x-y)+\frac{\left\|x-z\right\|\left\|x-y\right\|}{\sigma^{2}}\varphi_{\sigma}(x-y),$ (71)
where
  1. (a)
    follows by using $a^{2}-b^{2}=(a+b)(a-b)$ for $a,b\in\mathbb{R}$ and $\left\|u\right\|-\left\|v\right\|\leq\left\|u-v\right\|$ for $u,v\in\mathbb{R}^{d}$;
  2. (b)
    is due to $1-e^{-x}\leq x$ for $x\in\mathbb{R}$;
  3. (c)
    is since $\left\|z-y\right\|+\left\|x-y\right\|\leq\left\|x-z\right\|+2\left\|x-y\right\|$ for $x,y,z\in\mathbb{R}^{d}$.
Substituting (71) in (70), we obtain
$\left|\varphi_{\sigma}(x-y)B_{k,\sigma}(x-y)-\varphi_{\sigma}(z-y)B_{k,\sigma}(z-y)\right|$
$\qquad\qquad\qquad\qquad\leq\varphi_{\sigma}(x-y)\left|B_{k,\sigma}(x-y)-B_{k,\sigma}(z-y)\right|\left(1+\frac{\left\|x-z\right\|^{2}}{2\sigma^{2}}+\frac{\left\|x-z\right\|\left\|x-y\right\|}{\sigma^{2}}\right)$
$\qquad\qquad\qquad\qquad\qquad+\varphi_{\sigma}(x-y)\left|B_{k,\sigma}(x-y)\right|\left\|x-z\right\|\left(\frac{\left\|x-z\right\|}{2\sigma^{2}}+\frac{\left\|x-y\right\|}{\sigma^{2}}\right).$
Consider the term $B_{k,\sigma}(x-y)$. We have
$B_{k,\sigma}(x-y)$ $:=\prod_{j=1}^{d}H_{k_{j}}\left(\frac{x_{j}-y_{j}}{\sigma}\right)=\sum_{l=1}^{|k|}\sigma^{-l}\sum_{\begin{subarray}{c}s\in\mathbb{Z}_{\geq 0}^{d}:\\ |v|=l\end{subarray}}\tilde{c}_{k,s}\prod_{j=1}^{d}(x_{j}-y_{j})^{v_{j}},$
for some constants $\tilde{c}_{k,s}$. A crude upper bound on these constants can be obtained by using the closed form expression for the coefficients of a Hermite polynomial as
$\max_{\left|s\right|\leq\left|k\right|}\left|\tilde{c}_{k,s}\right|\leq\max_{\left|s\right|\leq\left|k\right|}\frac{|k|!}{|s|!(|k|-2|s|)!2^{|s|}}\leq\frac{|k|!}{\left(\left\lfloor\frac{\left|k\right|}{3}\right\rfloor!\right)^{2}2^{\left\lfloor\frac{\left|k\right|}{3}\right\rfloor}}\leq 3\left|k\right|^{\frac{3}{2}}(2\left|k\right|)^{\frac{\left|k\right|}{3}}.$
Hence,
$\left|B_{k,\sigma}(x-y)-B_{k,\sigma}(z-y)\right|\leq\sum_{l=1}^{|k|}\sigma^{-l}\sum_{\begin{subarray}{c}s\in\mathbb{Z}_{\geq 0}^{d}:\\ |s|=l\end{subarray}}\left|\tilde{c}_{k,s}\right|\left|\prod_{j=1}^{d}(x_{j}-y_{j})^{s_{j}}-\prod_{j=1}^{d}(z_{j}-y_{j})^{s_{j}}\right|.$ (72)
Now, note that for any $u,v\in\mathbb{R}^{i}$, we have
$\prod_{j=1}^{i}u_{j}-\prod_{j=1}^{i}v_{j}$ $=\sum_{l=0}^{i-1}\left(\prod_{j=1}^{i-l}u_{j}\prod_{j=i-l+1}^{i}v_{j}-\prod_{j=1}^{i-l-1}u_{j}\prod_{j=i-l}^{i}v_{j}\right)$
$=\sum_{l=0}^{i-1}(u_{i-l}-v_{i-l})\left(\prod_{j=1}^{i-l-1}u_{j}\right)\left(\prod_{j=i-l+1}^{i}v_{j}\right)$
Hence, we have
$\left|\prod_{j=1}^{i}u_{j}-\prod_{j=1}^{i}v_{j}\right|\leq\sum_{l=0}^{i-1}\left\|u-v\right\|_{\infty}\left(\left\|u\right\|_{\infty}+\left\|v\right\|_{\infty}\right)^{i}\leq i2^{i-1}\left\|u-v\right\|_{\infty}\big{(}\left\|u\right\|_{\infty}^{i}+\left\|v\right\|_{\infty}^{i}\big{)}.$
Substituting in (72), we obtain
$\left|B_{k,\sigma}(x-y)-B_{k,\sigma}(z-y)\right|$ $\leq\sum_{l=1}^{|k|}\sigma^{-l}\sum_{\begin{subarray}{c}s\in\mathbb{Z}_{\geq 0}^{d}:|s|=l\end{subarray}}\left|\tilde{c}_{k,s}\right||s|2^{|s|-1}\left\|x-z\right\|\left(\left\|x-y\right\|^{|s|}+\left\|z-y\right\|^{|s|}\right)$
$\leq\hat{c}_{d,\left|k\right|}\sigma^{-|k|}\left\|x-z\right\|\left(1+\left\|x-y\right\|^{|k|}+\left\|z-y\right\|^{|k|}\right)$
$\leq\hat{c}_{d,\left|k\right|}\sigma^{-|k|}\left\|x-z\right\|\left(1+\left\|x-y\right\|^{|k|}+\left\|x-z\right\|^{|k|}\right),$
where
$\hat{c}_{d,\left|k\right|}:=\frac{|k|!(\left|k\right|+1)^{d}}{\left(\left\lfloor\frac{\left|k\right|}{3}\right\rfloor!\right)^{2}2^{\left\lfloor\frac{\left|k\right|}{3}\right\rfloor}}.$
Hence, since $\sigma^{-|k|}\geq\sigma^{{-|k^{\prime}|}}$ for all $\sigma\leq 1$ and $|k|\geq|k^{\prime}|$, we have for such $\sigma$ that
$\left|\varphi_{\sigma}(x-y)B_{k,\sigma}(x-y)-\varphi_{\sigma}(z-y)B_{k,\sigma}(z-y)\right|$
$\leq\hat{c}_{d,\left|k\right|}\sigma^{-|k|}\varphi_{\sigma}(x-y)\left\|x-z\right\|\left(1+\left\|x-y\right\|^{|k|}+\left\|x-z\right\|^{|k|}\right)\left(1+\frac{\left\|x-z\right\|^{2}}{2\sigma^{2}}+\frac{\left\|x-z\right\|\left\|x-y\right\|}{\sigma^{2}}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad+\varphi_{\sigma}(x-y)\left|B_{k,\sigma}(x-y)\right|\left\|x-z\right\|\left(\frac{\left\|x-z\right\|}{2\sigma^{2}}+\frac{\left\|x-y\right\|}{\sigma^{2}}\right).$
Next observe that for $q\geq 1$ or $q=0$ and $f\in\mathcal{F}(b,q)$,
$\left|f(y)\right|\leq 0.5b(1+\left\|y\right\|^{q})\leq 0.5b(1+2^{q-1}\left\|x-y\right\|^{q}+2^{q-1}\left\|x\right\|^{q}).$
Hence, for $x\in\mathcal{X}$, we obtain using the definition of $v_{b,q}(\mathcal{X})$ (see (31)) that
$\left|f(y)\right|\leq 2^{q-1}\left(v_{b,q}(\mathcal{X})+b\left\|x-y\right\|^{q}\right).$
Using this in the above equation yields for $x\in\mathcal{X}$ that
$\left|f(y)\right|\left|\varphi_{\sigma}(x-y)B_{k,\sigma}(x-y)-\varphi_{\sigma}(z-y)B_{k,\sigma}(z-y)\right|$
$\leq 2^{q-1}\hat{c}_{d,\left|k\right|}\big{(}v_{b,q}(\mathcal{X})+b\left\|x-y\right\|^{q}\big{)}\varphi_{\sigma}(x-y)\left\|x-z\right\|\Bigg{(}\sigma^{-|k|}\left(1+\left\|x-y\right\|^{|k|}+\left\|x-z\right\|^{|k|}\right)$
$\qquad\qquad\qquad\left(1+\frac{\left\|x-z\right\|^{2}}{2\sigma^{2}}+\frac{\left\|x-z\right\|\left\|x-y\right\|}{\sigma^{2}}\right)+\left|B_{k,\sigma}(x-y)\right|\left(\frac{\left\|x-z\right\|}{2\sigma^{2}}+\frac{\left\|x-y\right\|}{\sigma^{2}}\right)\Bigg{)},$
$\leq 2^{q-1}\hat{c}_{d,\left|k\right|}\big{(}v_{b,q}(\mathcal{X})+b\left\|x-y\right\|^{q}\big{)}\varphi_{\sigma}(x-y)\left\|x-z\right\|\bigg{(}\sigma^{-|k|}\Big{(}1+\left\|x-y\right\|^{|k|}+\left\|x-z\right\|^{|k|}\Big{)}$
$\qquad\qquad\qquad\qquad+\left|B_{k,\sigma}(x-y)\right|\frac{\left\|x-y\right\|}{\sigma^{2}}\bigg{)}+\left\|x-z\right\|^{2}R_{d,q,\sigma,|k|},$ (73)
where
$R_{d,q,\sigma,|k|}$ $:=2^{q-1}\hat{c}_{d,\left|k\right|}\big{(}v_{b,q}(\mathcal{X})+b\left\|x-y\right\|^{q}\big{)}\varphi_{\sigma}(x-y)\bigg{(}\left|B_{k,\sigma}(x-y)\right|+\sigma^{-(|k|+2)}$
$\qquad\qquad\qquad\qquad\qquad\quad\times\Big{(}1+\left\|x-y\right\|^{|k|}+\left\|x-z\right\|^{|k|}\Big{)}\left(\left\|x-z\right\|+\left\|x-y\right\|\right)\bigg{)}.$
For $U\sim N(0,\sigma^{2})$ and $V=N(0,\sigma^{2}I_{d})$, a straightforward computation yields
$\big{(}\mathbb{E}[H_{m}(U)])^{2}\leq\mathbb{E}[H_{m}^{2}(U)]=m!,$ (74a)
$\mathbb{E}[\left\|V\right\|^{q}]^{\frac{1}{q}}\leq\sqrt{qd}\sigma.$ (74b)
From these and $\prod_{j=1}^{d}k_{j}!\leq|k|!$, we obtain using the definition of $B_{k,\sigma}(x-y)$ (see (68)) that
$\int_{\mathbb{R}^{d}}\left\|x-y\right\|^{a}\varphi_{\sigma}(x-y)dy=\left(\sqrt{ad}\sigma\right)^{a},$ (75a)
$\int_{\mathbb{R}^{d}}\left|B_{k,\sigma}(x-y)\right|\varphi_{\sigma}(x-y)dy\leq\left(\prod_{j=1}^{d}\mathbb{E}\Big{[}H_{k_{j}}^{2}(U)\Big{]}\right)^{\frac{1}{2}}\leq\sqrt{|k|!},$ (75b)
and
$\int_{\mathbb{R}^{d}}\left\|x-y\right\|^{a}\left|B_{k,\sigma}(x-y)\right|\varphi_{\sigma}(x-y)dy\leq\left(\mathbb{E}\Big{[}\left\|V\right\|^{2a}\Big{]}\right)^{\frac{1}{2}}\left(\prod_{j=1}^{d}\mathbb{E}\Big{[}H_{k_{j}}^{2}(U)\Big{]}\right)^{\frac{1}{2}}\leq\big{(}\sqrt{2ad}\sigma\big{)}^{a}\sqrt{|k|!},$ (75c)
where the penultimate inequality follows from Cauchy-Schwarz inequality. Substituting (73) in (69), integrating w.r.t. $y$ and using (75), we obtain for $x,z\in\mathcal{X}$ and $\sigma\leq 1$ that
$\left|D^{k}f_{\sigma}(x)-D^{k}f_{\sigma}(z)\right|$ $\leq\bar{c}_{d,|k|,q}v_{b,q}(\mathcal{X})\sigma^{-(|k|\vee 1)}\left\|x-z\right\|+O_{d,\sigma,q,|k|}\left(\left\|x-z\right\|^{2}\right),$
where
$\bar{c}_{d,|k|,q}$ $:=2^{q-1}\hat{c}_{d,\left|k\right|}\left(\sqrt{(q+1)|k|!}\vee\sqrt{(q+|k|)^{q+|k|}}\right)\left(\sqrt{d}\vee\sqrt{d^{q+|k|}}\right).$
From this, we obtain for $|k|\geq 1$, $x\in\mathcal{X}$ and $\sigma\leq 1$ that
$\left|D^{k}f_{\sigma}(x)\right|$ $\leq\bar{c}_{d,|k|-1,q}v_{b,q}(\mathcal{X})\sigma^{-\big{(}(|k|-1)\vee 1\big{)}}.$
On the other hand, for $x\in\mathcal{X}$, we have for $q\geq 1$ and $\sigma\leq 1$ that
$\left|f_{\sigma}(x)\right|$ $\leq\int_{\mathbb{R}^{d}}\left|f(y)\right|\varphi_{\sigma}(x-y)dy$
$\leq\int_{\mathbb{R}^{d}}0.5b(1+\left\|y\right\|^{q})\varphi_{\sigma}(x-y)dy$
$\leq\int_{\mathbb{R}^{d}}0.5b(1+2^{q-1}\left\|x\right\|^{q}+2^{q-1}\left\|x-y\right\|^{q})\varphi_{\sigma}(x-y)dy$
$\leq 2^{q-1}v_{b,q}(\mathcal{X})\left(1+\big{(}\sqrt{qd}\big{)}^{q}\right),$ (76)
where to obtain (76), we used the definition of $v_{b,q}(\mathcal{X})$ and (74b). For $f\in\mathcal{F}(b)$ (corresponds to $q=0$) and any $x\in\mathbb{R}^{d}$,
$\left|f_{\sigma}(x)\right|$ $\leq\int_{\mathbb{R}^{d}}\left|f(y)\right|\varphi_{\sigma}(x-y)dy\leq b\int_{\mathbb{R}^{d}}\varphi_{\sigma}(x-y)dy\leq b.$
Note that the RHS of (76) simplifies to $b$ when $q=0$, and hence we may use this expression as a common upper bound for $q\in\{0\}\cup[1,\infty)$. Combining all the above, we obtain that for $|k|\in\mathbb{Z}_{\geq 0}$ and $\sigma\leq 1$,
$\sup_{x\in\mathcal{X}}\left|D^{k}f_{\sigma}(x)\right|$ $\leq\bar{c}_{d,|k|,q}v_{b,q}(\mathcal{X})\sigma^{-\big{(}(|k|-1)\vee 1\big{)}}.$ (77)
Finally, we have
$\max_{|k|=\underaccent{\bar}{\beta}}\sup_{x,z}\frac{\left|D^{k}f_{\sigma}(x)-D^{k}f_{\sigma}(z)\right|}{\left\|x-z\right\|^{\beta-\underaccent{\bar}{\beta}}}$
$\leq\max_{|k|=\underaccent{\bar}{\beta}}\sup_{x,z}\left(\frac{\left|D^{k}f_{\sigma}(x)-D^{k}f_{\sigma}(z)\right|}{\left\|x-z\right\|}\right)^{\beta-\underaccent{\bar}{\beta}}\max_{|k|=\underaccent{\bar}{\beta}}\sup_{x,z}\left|D^{k}f_{\sigma}(x)-D^{k}f_{\sigma}(z)\right|^{1-\beta+\underaccent{\bar}{\beta}}$
$\leq 2\max_{|k|=\underaccent{\bar}{\beta}}\sup_{x,z}\left(\frac{\left|D^{k}f_{\sigma}(x)-D^{k}f_{\sigma}(z)\right|}{\left\|x-z\right\|}\right)^{\beta-\underaccent{\bar}{\beta}}\max_{|k|=\underaccent{\bar}{\beta}}\sup_{x}\left|D^{k}f_{\sigma}(x)\right|^{1-\beta+\underaccent{\bar}{\beta}}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left(\sqrt{d}\bar{c}_{d,\underaccent{\bar}{\beta},q}v_{b,q}(\mathcal{X})\sigma^{-(\underaccent{\bar}{\beta}\vee 1)}\right)^{\beta-\underaccent{\bar}{\beta}}\left(\bar{c}_{d,\underaccent{\bar}{\beta},q}v_{b,q}(\mathcal{X})\sigma^{-\big{(}(\underaccent{\bar}{\beta}-1)\vee 1\big{)}}\right)^{1-\beta+\underaccent{\bar}{\beta}}$
$\leq\check{c}_{d,\beta,q}v_{b,q}(\mathcal{X})\sigma^{-(\underaccent{\bar}{\beta}\vee 1)},$ (78)
where the suprema above are over $x,z$ in the interior of $\mathcal{X}$, and
$\check{c}_{d,\beta,q}$ $:=2^{q-1}\hat{c}_{d,\underaccent{\bar}{\beta}}\left(\sqrt{(q+1)\underaccent{\bar}{\beta}!}\vee\sqrt{(q+\underaccent{\bar}{\beta})^{q+\underaccent{\bar}{\beta}}}\right)\left(d\vee\sqrt{d^{\underaccent{\bar}{\beta}+q+1}}\right)$
$\leq 2^{q-1}\frac{\underaccent{\bar}{\beta}!(\underaccent{\bar}{\beta}+1)^{d}}{\left(\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor!\right)^{2}2^{\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor}}\left(\sqrt{(q+1)\underaccent{\bar}{\beta}!}\vee\sqrt{(q+\underaccent{\bar}{\beta})^{q+\underaccent{\bar}{\beta}}}\right)\left(d\vee\sqrt{d^{\underaccent{\bar}{\beta}+q+1}}\right).$ (79)
In $(a)$, we used (77) and
$\sup_{x,z}\frac{\left|D^{k}f_{\sigma}(x)-D^{k}f_{\sigma}(z)\right|}{\left\|x-z\right\|}\leq\sup_{x}\left\|\nabla\big{(}D^{k}f_{\sigma}(x)\big{)}\right\|\leq\sqrt{d}\bar{c}_{d,|k|,q}v_{b,q}(\mathcal{X})\sigma^{-(|k|\vee 1)}.$
Here, the first inequality follows by Lagrange’s mean value theorem for multiple variables which applies due to differentiability of $D^{k}f_{\sigma}$ everywhere, and the second inequality uses (77) along with $\left\|x\right\|\leq\sqrt{d}\left\|x\right\|_{\infty}$ for $x\in\mathbb{R}^{d}$.
Combining (77), (78) and the definition of Hölder norm given in (6), we obtain (32). Since $\left\|x\right\|^{q}\leq 1+\left\|x\right\|$ for all $q\in[0,1]$, the upper bound for $q=1$ (up to a universal multiplicative constant) also applies to $q\in[0,1)$.
Equipped with (32), to prove (33a)-(33b), we will use the following result from [75] that provides covering entropy bounds for Hölder ball of functions with bounded support.
Theorem 8 (Theorem 2.7.1 and Corollary 2.7.2 in [75]).
Let $\mathcal{X}$ be a bounded Borel-measurable convex subset of $\mathbb{R}^{d}$ with non-empty interior. For $\beta,L>0$, let $C_{L}^{\beta}(\mathcal{X})$ denote the functions $f:\mathcal{X}\rightarrow\mathbb{R}$ such that $\left\|f\right\|_{\beta,\mathcal{X}}\leq L$. Then
[…]
where $\mathcal{B}(\mathcal{X})$ denotes the set of positive Borel measures on $\mathcal{X}$. For $\mathcal{X}=\mathbb{B}_{d}(r)$, the above inequalities hold with $c_{d,\beta,\mathbb{B}_{d}(r)}$ as given in (98) below.
The proof of Theorem 8 given in [75] establishes (80), albeit without specifying the dependence of $d,\beta$ in the pre factor $c_{d,\beta,\mathcal{X}}$. In Appendix B, we provide an upper bound on this constant when $\mathcal{X}=\mathbb{B}_{d}(r)$, which will be used below to complete the proof of Lemma 3.
Let $\mathcal{F}_{\sigma}(b,q,\mathcal{X})$ denote the class of functions obtained by restricting the domain of functions in $\mathcal{F}_{\sigma}(b,q)$ to $\mathcal{X}$. It follows from (32) that $\mathcal{F}_{\sigma}\big{(}b,q,\mathbb{B}_{d}(r)\big{)}\subseteq C_{L}^{\beta}\big{(}\mathbb{B}_{d}(r)\big{)}$ for any $\beta>0$ with $L:=\check{c}_{d,\beta,q}b(1+r^{q})\sigma^{-(\underaccent{\bar}{\beta}\vee 1)}$, where $\check{c}_{d,\beta,q}$ is given in (79). Then, applying (80) with $L$ as above and noting that $\left\|F_{b,q}*\varphi_{\sigma}\right\|_{\infty,\mathbb{B}_{d}(r)}\geq 0.5b$ yields
[…]
where $c_{d,\beta}$ is as given in (98), and
$c_{d,\beta,q}$ $:=c_{d,\beta}\big{(}2\check{c}_{d,\beta,q}\big{)}^{\frac{d}{\beta}}$
$=(\underaccent{\bar}{\beta}+1)^{d}2^{\frac{d}{\beta}}d^{\frac{d}{2}}(e^{d}+1)^{\frac{d}{\beta}}\left(\frac{\beta 2^{\frac{d}{\beta}}}{d}\vee 3^{d}\log\Big{(}2^{\beta+1}d^{\frac{\beta}{2}}(e^{d}+1)+1\Big{)}\right)$
$\qquad\qquad\qquad\left(2^{q}\frac{\underaccent{\bar}{\beta}!(\underaccent{\bar}{\beta}+1)^{d}}{\left(\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor!\right)^{2}2^{\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor}}\left(\sqrt{(q+1)\underaccent{\bar}{\beta}!}\vee\sqrt{(q+\underaccent{\bar}{\beta})^{q+\underaccent{\bar}{\beta}}}\right)\left(d\vee\sqrt{d^{\underaccent{\bar}{\beta}+q+1}}\right)\right)^{\frac{d}{\beta}}.$ (82)
For $\beta=\frac{d}{2}+1$ and $q\in[0,1]$, simplifying the above expression yields that $c_{d,\frac{d}{2}+1,q}=O\big{(}d^{6d+9}(9e^{2})^{d}\big{)}$, where we used $\log x\leq x-1$ for $x>0$ and Stirling’s approximation for factorial given in [55]:
$\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}e^{\frac{1}{12n+1}}\leq n!\leq\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}e^{\frac{1}{12n}}.$
This proves (33a). Similarly, (81) yields (33b) by using that $\left\|F_{b,q}*\varphi_{\sigma}\right\|_{p,\gamma}\geq 0.5b$ for $p\geq 1$. Finally, the proof of the penultimate claim for $\mathcal{E}_{\sigma,\alpha}(b,q)$ follows straightforwardly via the same proof by replacing $v_{b,q}(\mathcal{X})$ by $e^{\alpha v_{b,q}(\mathcal{X})}$.
5.3. Proof of Lemma 1
We will show (12) with
[…]
where $\tilde{c}_{q}$ is the maximum of the constant appearing in [41, Theorem 6.21] and $1$. We will use the following statistical version of Talagrand’s inequality given in [6].
Theorem 9 (Theorem 2.1 in [6]).
Consider the setting of Theorem 6. Let $\{\varepsilon_{i}\}_{i=1}^{n}$ be i.i.d. Rademacher variables independent of $X^{n}$, and set $Z^{(\varepsilon)}_{n}:=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|\right]$. Then, for all $z\geq 0$,
$\mathbb{P}\left(Z_{n}\geq\inf_{\eta\in(1,\infty)}\left\{2\eta Z^{(\varepsilon)}_{n}+\theta\sqrt{2nz}+\frac{(3\eta^{2}+11\eta+4)bz}{3(\eta-1)}\right\}\right)\leq 2e^{-z}.$ (84)
In particular, for every $z\geq 0$ and $\eta\in(1,\infty)$,
$\mathbb{P}\left(Z_{n}\geq 2\eta Z^{(\varepsilon)}_{n}+z\right)\leq 2e^{-\left(\frac{z^{2}}{8n\hat{\theta}_{n}^{2}}\wedge\frac{3z(\eta-1)}{2(11\eta+3\eta^{2}+4)b}\right)}.$ (85)
Equation (84) follows from [6, Theorem 2.1] by setting $\eta=\frac{1+\alpha}{1-\alpha}$.
Let $T_{n}$ be a truncation level (to be specified later) for the truncated function class defined in Lemma 1. We have
$Z_{n}\leq Z_{n}^{\prime}+Z_{n}^{\prime\prime},$ (86)
where
$Z_{n}$ $:=\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\mu}[f]\big{)}\right|,$
$Z_{n}^{\prime}$ $:=\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\big{(}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})-\mathbb{E}_{\mu}\left[f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}\right]\big{)}\right|=\sup_{f\in\mathcal{F}_{n}(T_{n})}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\mu}[f]\big{)}\right|,$
$Z_{n}^{\prime\prime}$ $:=\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\Big{(}f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})-\mathbb{E}_{\mu}\left[f\mathbbm{1}_{\{F_{n}>T_{n}\}}\right]\Big{)}\right|.$
Set
$Z_{n}^{{}^{\prime}\varepsilon}:=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}_{n}(T_{n})}\left|\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|\right].$
Let $\eta\in(1,\infty)$. From the above and (86), it follows that
$\mathbb{P}\left(Z_{n}\geq 2\eta Z_{n}^{{}^{\prime}\varepsilon}+z\right)$ $\leq\mathbb{P}\left(Z_{n}^{\prime}+Z_{n}^{\prime\prime}\geq 2\eta Z_{n}^{{}^{\prime}\varepsilon}+z\right)\leq\mathbb{P}\left(Z_{n}^{\prime}\geq 2\eta Z_{n}^{{}^{\prime}\varepsilon}+\frac{3z}{4}\right)+\mathbb{P}\left(Z_{n}^{\prime\prime}\geq\frac{z}{4}\right).$ (87)
By symmetrization inequality (see e.g. [75, Lemma 2.3.1]), we have
$\mathbb{E}\big{[}Z_{n}^{\prime\prime}\big{]}\leq 2\mathbb{E}\left[\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\right].$ (88)
Then, applying the Hoffmann-Jørgensen inequality, we obtain
$\mathbb{E}\left[\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\right]$ $\leq 8\mathbb{E}\left[\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\right]+8t_{0}(T_{n})$
$\leq 8\mathbb{E}\left[M_{n}\right]+8t_{0}(T_{n}),$
where $M_{n}:=\max_{1\leq i\leq n}F_{n}(X_{i})$ and
$t_{0}(T_{n}):=\inf\left\{t>0:\mathbb{P}\left(\max_{1\leq k\leq n}\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{k}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|>t\right)\leq 1/8\right\}.$
For $T_{n}=T_{n}^{\star}:=8\mathbb{E}[M_{n}]$, Markov’s inequality yields
$\mathbb{P}\left(\max_{1\leq k\leq n}\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{k}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|>0\right)\leq\mathbb{P}\left(M_{n}>T_{n}\right)\leq\frac{\mathbb{E}[M_{n}]}{T_{n}}=\frac{1}{8}.$
Hence, $t_{0}\big{(}T_{n}^{\star}\big{)}=0$. Combining the above and substituting in (88), we obtain
$\left\|Z_{n}^{\prime\prime}\right\|_{1}=\mathbb{E}\big{[}Z_{n}^{\prime\prime}\big{]}\leq 16\mathbb{E}\left[M_{n}\right]\stackrel{{\scriptstyle(a)}}{{\leq}}16\left\lceil\frac{1}{q}\right\rceil!\left\|M_{n}\right\|_{\psi_{q}},$ (89)
where $(a)$ used (103) given below.
Next, applying (85) to the bounded function class $\mathcal{F}_{n}(T_{n}^{\star})$ with $b=8\mathbb{E}[M_{n}]$, we have for every $\eta\geq 1$ that
$\mathbb{P}\left(Z_{n}^{\prime}\geq 2\eta Z_{n}^{{}^{\prime}\varepsilon}+\frac{3z}{4}\right)\leq e^{-\big{(}\frac{9z^{2}}{128n\left\|F_{n}\right\|_{2,\mu}^{2}}\wedge\frac{9z(\eta-1)}{64(11\eta+3\eta^{2}+4)\mathbb{E}[M_{n}]}\big{)}},$
where we used
$\hat{\theta}_{n}^{2}:=\sup_{f\in\mathcal{F}_{n}(T_{n}^{\star})}\mathbb{E}_{\mu}\big{[}\big{(}f-\mathbb{E}_{\mu}[f]\big{)}^{2}\big{]}\leq\sup_{f\in\mathcal{F}_{n}(T_{n}^{\star})}\mathbb{E}_{\mu}\big{[}f^{2}\big{]}\leq\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\mu}\big{[}f^{2}\big{]}\leq\left\|F_{n}\right\|_{2,\mu}^{2}.$
Choosing $\eta=1+\sqrt{6}$ to maximize the second exponent, we obtain
[…]
For controlling the second term in (87), note that for $q\in(0,1]$,
$\mathbb{P}\left(Z_{n}^{\prime\prime}\geq\frac{z}{4}\right)\leq 2e^{\frac{-z^{q}}{4^{q}\left\|Z_{n}^{\prime\prime}\right\|_{\psi_{q}}^{q}}}.$ (91)
From [41, Theorem 6.21], we obtain
$\left\|Z_{n}^{\prime\prime}\right\|_{\psi_{q}}\leq\tilde{c}_{q}\left(\left\|Z_{n}^{\prime\prime}\right\|_{1}+\left\|\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})-\mathbb{E}_{\mu}\left[f\mathbbm{1}_{\{F_{n}>T_{n}\}}\right]\right|\right\|_{\psi_{q}}\right).$ (92)
We can upper bound the last term in the above equation as follows:
$\left\|\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})-\mathbb{E}_{\mu}\big{[}f\mathbbm{1}_{\{F_{n}>T_{n}\}}\big{]}\right|\right\|_{\psi_{q}}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}2^{\frac{1}{q}-1}\left\|\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\right\|_{\psi_{q}}+2^{\frac{1}{q}-1}\left\|\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|\mathbb{E}_{\mu}\big{[}f\mathbbm{1}_{\{F_{n}>T_{n}\}}\big{]}\right|\right\|_{\psi_{q}}$
[…]
where
  1. (a)
    follows from $\left\|W_{1}+W_{2}\right\|_{\psi_{q}}\leq 2^{\frac{1}{q}-1}\big{(}\left\|W_{1}\right\|_{\psi_{q}}+\left\|W_{2}\right\|_{\psi_{q}}\big{)}$ for $0<q\leq 1$ as shown in (99) below;
  2. (b)
    is due to $\left\|\mathbb{E}_{\mu}[W]\right\|_{\psi_{q}}\leq\left\lceil\frac{1}{q}\right\rceil!(\log 2)^{-\frac{1}{q}}\left\|W\right\|_{\psi_{q}}$ for $W\sim\mu$ as shown in (102).
Using the above bound and (89) in (92) and substituting the resulting bound in (91) yields
$\mathbb{P}\left(Z_{n}^{\prime\prime}\geq\frac{z}{4}\right)\leq 2e^{\frac{-z^{q}}{(c_{q}\left\|M_{n}\right\|_{\psi_{q}})^{q}}},$
with $c_{q}$ as given in (83). By substituting the above inequality to control the last term in (87) and using (90) to upper bound the first term therein, we obtain
[…]
To complete the proof, we will show that $Z^{{}^{\prime}(\varepsilon)}_{n}\leq Z^{(\varepsilon)}_{n}$ for $X^{n}$ surely. Fix $X^{n}$ and set $\mathcal{I}(X^{n}):=\{i\in[1:n],F_{n}(X_{i})\leq T_{n}\}$, $\bar{\mathcal{I}}(X^{n}):=\{1,\ldots,n\}\setminus\mathcal{I}(X^{n})$, and $\varepsilon_{\mathcal{I}(X^{n})}:=\{\varepsilon_{i},~i\in\mathcal{I}(X^{n})\}$. Then
$Z^{{}^{\prime}(\varepsilon)}_{n}$ $=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i=1}^{n}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})\right|\right]$
$=\mathbb{E}_{\varepsilon_{\mathcal{I}(X^{n})}}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})\right|\right]$
$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{\varepsilon_{\mathcal{I}(X^{n})}}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})+\sum_{i\in\bar{\mathcal{I}}(X^{n})}\mathbb{E}_{\varepsilon_{i}}\left[\varepsilon_{i}f(X_{i})\right]\right|\right]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})+\sum_{i\in\bar{\mathcal{I}}(X^{n})}\varepsilon_{i}f(X_{i})\right|\right]$
$=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f(X_{i})+\sum_{i\in\bar{\mathcal{I}}(X^{n})}\varepsilon_{i}f(X_{i})\right|\right]$
$=Z^{(\varepsilon)}_{n}$,
where $(a)$ is because $\mathbb{E}_{\varepsilon_{i}}\left[\varepsilon_{i}f(X_{i})\right]=\mathbb{E}_{\varepsilon_{i}}\left[\varepsilon_{i}\right]f(X_{i})=0$ by independence of $\varepsilon_{i}$ and $X_{i}$, and $\mathbb{E}_{\varepsilon_{i}}\left[\varepsilon_{i}\right]=0$, and $(b)$ is via Jensen’s inequality. This completes the proof.
6. Concluding Remarks
We established exponential deviation inequalities for smoothed plug-in estimators and neural estimators of Rényi divergences. The derived inequalities apply to smoothed KL divergence estimate when the unsmoothed distributions are compactly supported or sub-Gaussian, and to smoothed Rényi divergence estimate of order $\alpha\neq 1$ when the distributions are compactly supported. It is worth emphasizing that the dependence of data dimension and distributional parameters are specified in our bounds up to universal constants. However, these dependencies are possibly sub-optimal, and a tighter analysis quantifying their impact maybe warranted in scenarios where parameters are not fixed. For instance, such a scenario naturally arises in physical layer security, where the dimension (codeword length) scales as a function of sample size (codebook size). Also of interest is to determine whether the deviation inequalities are optimal in terms of dependence on $n$, $d$, and other distributional parameters, and otherwise, ascertain the gap from optimality. We hope that our results will be a starting point for further investigations along these lines.
Appendix A Variational expression for Rényi divergences
The following variational expression holds for Rényi divergences similar in spirit to the expression (1b) for KL divergence, which has the advantage of linearizing the second term in (2).
Lemma 7 (Variational expression for Rényi divergence).
Let $\alpha\in(0,1)\cup(1,\infty)$. Then
$\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)=\sup_{f}\frac{\alpha}{\alpha-1}\log\left(\int_{\mathbb{R}^{d}}e^{(\alpha-1)f}d\mu\right)-\int_{\mathbb{R}^{d}}e^{\alpha f}d\nu+1,$ (93)
where the supremum is over all measurable functions such that the second integral is finite. When $\mu\ll\nu$ and $\left(\frac{d\mu}{d\nu}\right)^{\alpha}\in L_{1}(\nu)$, the supremum is achieved by $f_{\alpha}^{\star}=\log\frac{d\mu}{d\nu}+\frac{1-\alpha}{\alpha}\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)$.
Proof.
We start with the variational expression given in [10]:
$\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)=\sup_{f}\frac{\alpha}{\alpha-1}\log\left(\int_{\mathbb{R}^{d}}e^{(\alpha-1)f}d\mu\right)-\log\left(\int_{\mathbb{R}^{d}}e^{\alpha f}d\nu\right),$ (94)
where the supremum is taken over the set of all bounded continuous functions. Substituting $g=e^{f}$, we obtain
$\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)=\sup_{g>0}\frac{\alpha}{\alpha-1}\log\left(\int_{\mathbb{R}^{d}}g^{(\alpha-1)}d\mu\right)-\log\left(\int_{\mathbb{R}^{d}}g^{\alpha}d\nu\right),$
where the supremum is over bounded continuous functions $g>0$. Note that the expression on the RHS is invariant to scaling the function $g$ by a positive constant $c$. Hence, we have
$\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)=\sup_{g>0}\frac{\alpha}{\alpha-1}\log\left(\int_{\mathbb{R}^{d}}(cg)^{(\alpha-1)}d\mu\right)-\log\left(\int_{\mathbb{R}^{d}}(cg)^{\alpha}d\nu\right).$
Since $\int_{\mathbb{R}^{d}}g^{\alpha}d\nu>0$ for any $g>0$, we may choose $c>0$ such that $\int_{\mathbb{R}^{d}}(cg)^{\alpha}d\nu=1$. Then
$\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)$ $=\sup_{\begin{subarray}{c}g>0:\\ \int_{\mathbb{R}^{d}}g^{\alpha}d\nu=1\end{subarray}}\frac{\alpha}{\alpha-1}\log\left(\int_{\mathbb{R}^{d}}g^{(\alpha-1)}d\mu\right)-\log\left(\int_{\mathbb{R}^{d}}g^{\alpha}d\nu\right)$
$=\sup_{\begin{subarray}{c}g>0:\\ \int_{\mathbb{R}^{d}}g^{\alpha}d\nu=1\end{subarray}}\frac{\alpha}{\alpha-1}\log\left(\int_{\mathbb{R}^{d}}g^{(\alpha-1)}d\mu\right)-\int_{\mathbb{R}^{d}}g^{\alpha}d\nu+1$
$\leq\sup_{g>0}\frac{\alpha}{\alpha-1}\log\left(\int_{\mathbb{R}^{d}}g^{(\alpha-1)}d\mu\right)-\int_{\mathbb{R}^{d}}g^{\alpha}d\nu+1$
$\leq\sup_{f}\frac{\alpha}{\alpha-1}\log\left(\int_{\mathbb{R}^{d}}e^{(\alpha-1)f}d\mu\right)-\int_{\mathbb{R}^{d}}e^{\alpha f}d\nu+1,$
where the supremum is over all bounded continuous functions $f$. On the other hand, by applying $\log x\leq x-1$ for $x\geq 0$ to last term in (94), we obtain the reverse inequality, thus proving (93). When $\mu\ll\nu$ and $\left(\frac{d\mu}{d\nu}\right)^{\alpha}\in L_{1}(\nu)$, $\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)<\infty$ and the integrals in (93) are finite. Moreover, it can be seen by substitution that the supremum in (93) is achieved by $f_{\alpha}^{\star}=\log\frac{d\mu}{d\nu}+\frac{1-\alpha}{\alpha}\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)$. ∎
Appendix B Determination of Prefactor in Theorem 8
The claim in (80) follows by determining the constants appearing in the proof of [75, Theorem 2.7.1]. It is sufficient to consider the case $L=1$ since the claim for general $L$ follows by scaling. Let $\delta=\epsilon^{\frac{1}{\beta}}$, $\left\|\cdot\right\|$ denote the Euclidean norm on $\mathbb{R}^{d}$, and $\{x_{(i)}\}_{i=1}^{m}$, with $m=N(\delta,\mathbb{B}_{d}(r),\left\|\cdot\right\|)$, denote the set of points in the interior of $\mathbb{B}_{d}(r)$ such that balls of radius $\delta$ centered at these points cover $\mathbb{B}_{d}(r)$. We will use the fact that $m\leq\left(1+\frac{2r}{\delta}\right)^{d}$ for $\delta<r$ and $m=1$ for $\delta\geq r$, which implies that $m\leq 1\vee\left(\frac{3r}{\delta}\right)^{d}$. Recall the differential operator $D^{k}$ for a $d$-dimensional vector $k$ with non-negative integers as elements (see Definition 1). Set $\frac{x^{k}}{k!}:=\prod_{i=1}^{d}\frac{x_{i}^{k_{i}}}{k_{i}!}$. For $k$ with $\left|k\right|\leq\underaccent{\bar}{\beta}$, and for each $f$, consider the vector
$A_{k}f=\left(\left\lfloor\frac{D^{k}f\big{(}x_{(1)}\big{)}}{\delta^{\beta-|k|}}\right\rfloor,\cdots,\left\lfloor\frac{D^{k}f\big{(}x_{(m)}\big{)}}{\delta^{\beta-|k|}}\right\rfloor\right).$
For any $x\in\mathbb{B}_{d}(r)$, there exists an $x_{(i)}$ for some $1\leq i\leq m$ such that $\left\|x-x_{(i)}\right\|\leq\delta$. By multivariate Taylor’s theorem with remainder estimates given in [23, Equation 8], for $f,g\in C_{1}^{\beta}\big{(}\mathbb{B}_{d}(r)\big{)}$ and $x\in\mathrm{int}\big{(}\mathbb{B}_{d}(r)\big{)}$, we have
$(f-g)(x)=\sum_{|k|\leq\underaccent{\bar}{\beta}}D^{k}(f-g)\big{(}x_{(i)}\big{)}\frac{(x-x_{(i)})^{k}}{k!}+R_{\beta}\big{(}f-g,x_{(i)},x-x_{(i)}\big{)},$ (95)
where
$R_{\beta}(f,x_{i},x-x_{(i)})=\underaccent{\bar}{\beta}\sum_{|k|=\underaccent{\bar}{\beta}}\frac{\big{(}x-x_{(i)}\big{)}^{k}}{k!}\int_{0}^{1}(1-t)^{\underaccent{\bar}{\beta}-1}\Big{(}D^{k}f\big{(}(1-t)x_{(i)}+tx\big{)}-D^{k}f\big{(}x_{(i)}\big{)}\Big{)}dt.$ (96)
By definition of $C_{1}^{\beta}\big{(}\mathbb{B}_{d}(r)\big{)}$, for $k$ such that $\left|k\right|=\underaccent{\bar}{\beta}$,
$\left|D^{k}f\big{(}(1-t)x_{(i)}+tx\big{)}-D^{k}f\big{(}x_{(i)}\big{)}\right|\leq\left\|x-x_{(i)}\right\|^{\beta-\underaccent{\bar}{\beta}}.$
Then, we obtain via triangle inequality that
$\left|D^{k}f\big{(}(1-t)x_{(i)}+tx\big{)}-D^{k}f\big{(}x_{(i)}\big{)}-D^{k}g\big{(}(1-t)x_{(i)}+tx\big{)}+D^{k}g\big{(}x_{(i)}\big{)}\right|\leq 2\left\|x-x_{(i)}\right\|^{\beta-\underaccent{\bar}{\beta}}.$
This yields as shown in the last equation in [23] that
$\left|R_{\beta}(f-g,x_{(i)},x-x_{(i)})\right|\leq 2\frac{\left\|x-x_{(i)}\right\|_{1}^{\underaccent{\bar}{\beta}}\left\|x-x_{(i)}\right\|^{\beta-\underaccent{\bar}{\beta}}}{\underaccent{\bar}{\beta}!}\leq 2d^{\frac{\underaccent{\bar}{\beta}}{2}}\frac{\left\|x-x_{(i)}\right\|^{\beta}}{\underaccent{\bar}{\beta}!}\leq\frac{2d^{\frac{\underaccent{\bar}{\beta}}{2}}\delta^{\beta}}{\underaccent{\bar}{\beta}!}.$
Thus, if $A_{k}f=A_{k}g$ for all $k$ such that $\left|k\right|\leq\underaccent{\bar}{\beta}$, we obtain from (95)-(96) that
$\left|f(x)-g(x)\right|\leq 2\sum_{|k|\leq\underaccent{\bar}{\beta}}d^{\frac{|k|}{2}}\left(\delta^{\beta-|k|}\frac{\delta^{|k|}}{k!}\right)+\frac{2d^{\frac{\underaccent{\bar}{\beta}}{2}}\delta^{\beta}}{\underaccent{\bar}{\beta}!}\leq 2d^{\frac{\underaccent{\bar}{\beta}}{2}}\delta^{\beta}\left(e^{d}+\frac{1}{\underaccent{\bar}{\beta}!}\right)\leq 2d^{\frac{\underaccent{\bar}{\beta}}{2}}\left(e^{d}+\frac{1}{\underaccent{\bar}{\beta}!}\right)\epsilon.$
[…]
Next, assume without loss of generality that $\{x_{(i)}\}_{i=1}^{m}$ is enumerated such that for each $j>1$, there is an index $i<j$ such that $\left\|x_{(i)}-x_{(j)}\right\|\leq 2\delta$. This follows by considering a sorting algorithm that iteratively builds the desired enumerated list by choosing a point (from the remaining unlisted points) and adding them at the rightmost possible location in the list such that the aforementioned property holds at each iteration. The feasibility of this procedure at each step is guaranteed. Otherwise the union of $\delta$ balls centered at points in the list would be disjoint with a strictly positive gap from the union of $\delta$ balls centered at points outside the list, thus contradicting the assumption that $\{x_{(i)}\}_{i=1}^{m}$ forms a $\delta$-covering of $\mathbb{B}_{d}(r)$.
Fixing the first column of $Af(x_{(1)})$, the smoothness properties of $f$ enforces limitations on the values the other columns may take. To see this, for each column indexed by $j>1$, choose an $i<j$ such that $\left\|x_{(i)}-x_{(j)}\right\|\leq 2\delta$. By Taylor’s theorem,
$D^{k}f\big{(}x_{(j)}\big{)}=\sum_{\left|l\right|\leq\underaccent{\bar}{\beta}-|k|}D^{k+l}f\big{(}x_{(i)}\big{)}\frac{\big{(}x_{(j)}-x_{(i)}\big{)}^{l}}{l!}+R_{\beta-\left|k\right|}\big{(}f,x_{(i)},x_{(j)}-x_{(i)}\big{)},$ (97)
with
$R_{\beta-\left|k\right|}\big{(}f,x_{(i)},x_{(j)}-x_{(i)}\big{)}$
$=(\underaccent{\bar}{\beta}-\left|k\right|)\sum_{|l|=\underaccent{\bar}{\beta}-\left|k\right|}\frac{\big{(}x_{(j)}-x_{(i)}\big{)}^{l}}{l!}\int_{0}^{1}(1-t)^{\underaccent{\bar}{\beta}-\left|k\right|-1}\Big{(}D^{k+l}f\big{(}(1-t)x_{(i)}+tx_{(j)}\big{)}-D^{k+l}f\big{(}x_{(i)}\big{)}\Big{)}dt.$
Let $B_{k}f:=\delta^{\beta-|k|}A_{k}f$ and $Bf$ be the matrix obtained by stacking $B_{k}f$ as rows. Suppose $Af\big{(}x_{(i)}\big{)}=Ag\big{(}x_{(i)}\big{)}$. Then $Bf\big{(}x_{(i)}\big{)}=Bg\big{(}x_{(i)}\big{)}$. Hence
$\left|D^{k}f\big{(}x_{(j)}\big{)}-D^{k}g\big{(}x_{(j)}\big{)}\right|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left|D^{k}f\big{(}x_{(j)}\big{)}-\sum_{\left|l\right|\leq\underaccent{\bar}{\beta}-|k|}B^{k+l}f\big{(}x_{(i)}\big{)}\frac{\big{(}x_{(j)}-x_{(i)}\big{)}^{l}}{l!}\right|+\left|D^{k}g\big{(}x_{(j)}\big{)}-\sum_{\left|l\right|\leq\underaccent{\bar}{\beta}-|k|}B^{k+l}g\big{(}x_{(i)}\big{)}\frac{\big{(}x_{(j)}-x_{(i)}\big{)}^{l}}{l!}\right|$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{\left|l\right|\leq\underaccent{\bar}{\beta}-|k|}\left|D^{k+l}f\big{(}x_{(i)}\big{)}-B^{k+l}f\big{(}x_{(i)}\big{)}\right|\frac{\left|\big{(}x_{(j)}-x_{(i)}\big{)}^{l}\right|}{l!}+\left|D^{k+l}g\big{(}x_{(i)}\big{)}-B^{k+l}g\big{(}x_{(i)}\big{)}\right|\frac{\left|\big{(}x_{(j)}-x_{(i)}\big{)}^{l}\right|}{l!}$
$\leq 2^{\beta-|k|+1}\sum_{\left|l\right|\leq\underaccent{\bar}{\beta}-|k|}d^{\frac{|l|}{2}}\delta^{\beta-|k|-|l|}\frac{\delta^{|l|}}{l!}+R_{\beta-\left|k\right|}\big{(}f,x_{(i)},x_{(j)}-x_{(i)}\big{)}+R_{\beta-\left|k\right|}\big{(}g,x_{(i)},x_{(j)}-x_{(i)}\big{)}$
$\leq 2^{\beta-|k|+1}d^{\frac{\underaccent{\bar}{\beta}-|k|}{2}}\delta^{\beta-|k|}e^{d}+2^{\beta-|k|+1}d^{\frac{\underaccent{\bar}{\beta}-|k|}{2}}\frac{\delta^{\beta-|k|}}{(\underaccent{\bar}{\beta}-|k|)!}$
$\leq 2^{\beta-|k|}\hat{c}_{d,\beta-|k|}\delta^{\beta-|k|},$
where $(a)$ follows via triangle inequality and $(b)$ is by substitution of (97). This means that fixing the values of the $i^{th}$ column of $Af$, each value in the $j^{th}$ column of $Af$ ranges over integers in an interval of length at most $2^{\beta-|k|}\hat{c}_{d,\beta-|k|}+1$. Using the upper bound $1\vee(2\delta^{-\beta})^{(\underaccent{\bar}{\beta}+1)^{d}}$ for the number of different values in the first column of $Af$, it follows that the number of possible matrices $Af$ and hence the covering number is upper bounded as
[…]
Taking logarithms, we obtain
[…]
$\leq(\underaccent{\bar}{\beta}+1)^{d}\frac{\beta}{d}2^{\frac{d}{\beta}}\epsilon^{-\frac{d}{\beta}}+\left(\frac{3r}{\delta}\right)^{d}(\underaccent{\bar}{\beta}+1)^{d}\log\big{(}2^{\beta}\hat{c}_{d,\beta}+1\big{)}$
$=:(\underaccent{\bar}{\beta}+1)^{d}\left(\frac{\beta 2^{\frac{d}{\beta}}}{d}+(3r)^{d}\log\big{(}2^{\beta}\hat{c}_{d,\beta}+1\big{)}\right)\epsilon^{-\frac{d}{\beta}}.$
Hence, by scaling $\epsilon$, (80) follows with
$c_{d,\beta,\mathbb{B}_{d}(r)}$ $:=(\underaccent{\bar}{\beta}+1)^{d}\hat{c}_{d,\beta}^{\frac{d}{\beta}}\left(\frac{\beta 2^{\frac{d}{\beta}}}{d}+(3r)^{d}\log\big{(}2^{\beta}\hat{c}_{d,\beta}+1\big{)}\right)$
$=(\underaccent{\bar}{\beta}+1)^{d}2^{\frac{d}{\beta}}d^{\frac{d}{2}}(e^{d}+1)^{\frac{d}{\beta}}\left(\frac{\beta 2^{\frac{d}{\beta}}}{d}+(3r)^{d}\log\Big{(}2^{\beta+1}d^{\frac{\beta}{2}}(e^{d}+1)+1\Big{)}\right)$
$=\underbrace{(\underaccent{\bar}{\beta}+1)^{d}2^{\frac{d}{\beta}}d^{\frac{d}{2}}(e^{d}+1)^{\frac{d}{\beta}}\left(\frac{\beta 2^{\frac{d}{\beta}}}{d}\vee 3^{d}\log\Big{(}2^{\beta+1}d^{\frac{\beta}{2}}(e^{d}+1)+1\Big{)}\right)}_{c_{d,\beta}}(1+r^{d})$ (98)
Given (80), the upper bound in (81) follows via similar steps given in [75, Corollary 2.7.2], which deals with bracketing entropy in place of covering entropy.
Appendix C Auxiliary Results related to Orlicz norm
C.1. Elementary inequalities involving Orlicz norm
Let $0<q\leq 1$ and $W_{1},W_{2}$ be Banach space valued random variables. We will show that
$\left\|W_{1}+W_{2}\right\|_{\psi_{q}}\leq 2^{\frac{1}{q}-1}\big{(}\left\|W_{1}\right\|_{\psi_{q}}+\left\|W_{2}\right\|_{\psi_{q}}\big{)}.$ (99)
Note that for $a,b\geq 0$ and $0\leq q\leq 1$, we have
$2^{q-1}(a^{q}+b^{q})\leq(a+b)^{q}=\frac{a+b}{(a+b)^{1-q}}\leq\frac{a}{(a+b)^{1-q}}+\frac{b}{(a+b)^{1-q}}\leq a^{q}+b^{q},$
where the first inequality follows from Jensen’s inequality applied to the concave map $x\mapsto x^{q}$. Denoting by $\left\|\cdot\right\|_{\mathfrak{B}}$ the relevant Banach space norm and setting $a=\left\|W_{1}\right\|_{\mathfrak{B}}$, $b=\left\|W_{2}\right\|_{\mathfrak{B}}$, $\bar{a}=\left\|W_{1}\right\|_{\psi_{q}}$, $\bar{b}=\left\|W_{2}\right\|_{\psi_{q}}$ and $\bar{c}_{q}=2^{\frac{1}{q}-1}$, we obtain by using the above inequalities that
$\mathbb{E}\left[e^{\left(\frac{\left\|W_{1}+W_{2}\right\|_{\mathfrak{B}}}{\bar{c}_{q}(\bar{a}+\bar{b})}\right)^{q}}\right]\leq\mathbb{E}\left[e^{\frac{(a+b)^{q}}{\bar{c}_{q}^{q}(\bar{a}+\bar{b})^{q}}}\right]\leq\mathbb{E}\left[e^{\frac{a^{q}+b^{q}}{\bar{c}_{q}^{q}(\bar{a}+\bar{b})^{q}}}\right]\leq\mathbb{E}\left[e^{\frac{a^{q}+b^{q}}{\bar{c}_{q}^{q}2^{q-1}(\bar{a}^{q}+\bar{b}^{q})}}\right]=\mathbb{E}\left[e^{\frac{a^{q}+b^{q}}{\bar{a}^{q}+\bar{b}^{q}}}\right].$ (100)
Continuing, we have
$\mathbb{E}\left[e^{\frac{a^{q}+b^{q}}{\bar{a}^{q}+\bar{b}^{q}}}\right]\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{\bar{a}^{q}}{\bar{a}^{q}+\bar{b}^{q}}\mathbb{E}\left[e^{\left(\frac{\left\|W_{1}\right\|_{\mathfrak{B}}}{\bar{a}}\right)^{q}}\right]+\frac{\bar{b}^{q}}{\bar{a}^{q}+\bar{b}^{q}}\mathbb{E}\left[e^{\left(\frac{\left\|W_{2}\right\|_{\mathfrak{B}}}{\bar{b}}\right)^{q}}\right]\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{2\bar{a}^{q}}{\bar{a}^{q}+\bar{b}^{q}}+\frac{2\bar{b}^{q}}{\bar{a}^{q}+\bar{b}^{q}}\leq 2,$ (101)
where $(a)$ is via Jensen’s inequality applied to the convex map $x\mapsto e^{x}$ and $(b)$ follows from the definition of Orlicz norm $\left\|\cdot\right\|_{\psi_{q}}$ which implies
$\mathbb{E}\left[e^{\left(\frac{\left\|W_{1}\right\|_{\mathfrak{B}}}{\bar{a}}\right)^{q}}\right]\vee\mathbb{E}\left[e^{\left(\frac{\left\|W_{2}\right\|_{\mathfrak{B}}}{\bar{b}}\right)^{q}}\right]\leq 2.$
Using the definition of Orlicz norm $\left\|\cdot\right\|_{\psi_{q}}$ again, we have $\left\|W_{1}+W_{2}\right\|_{\psi_{q}}\leq c_{q}(\bar{a}+\bar{b})$ since combining (100)-(101) yields
$\mathbb{E}\left[e^{\left(\frac{\left\|W_{1}+W_{2}\right\|_{\mathfrak{B}}}{c_{q}(\bar{a}+\bar{b})}\right)^{q}}\right]\leq 2.$
Next, we show that for a Banach space valued random variable $W\sim\mu$ such that $\mathbb{E}_{\mu}[W]$ exists (as a Bochner integral),
$\left\|\mathbb{E}_{\mu}[W]\right\|_{\psi_{q}}\leq\left\lceil\frac{1}{q}\right\rceil!~(\log 2)^{-\frac{1}{q}}\left\|W\right\|_{\psi_{q}}.$ (102)
We may assume that $\left\|\mathbb{E}_{\mu}[W]\right\|_{\psi_{q}}>0$ since otherwise (102) holds trivially. In this case, by continuity of $\psi_{q}$ and the definition of Orlicz norm $\left\|\cdot\right\|_{\psi_{q}}$, we have
$e^{\left(\frac{\left\|\mathbb{E}_{\mu}[W]\right\|_{\mathfrak{B}}}{\left\|\mathbb{E}_{\mu}[W]\right\|_{\psi_{q}}}\right)^{q}}=2.$
This implies that
$\left\|\mathbb{E}_{\mu}[W]\right\|_{\psi_{q}}=(\log 2)^{-\frac{1}{q}}\left\|\mathbb{E}_{\mu}[W]\right\|_{\mathfrak{B}}\leq\left\lceil\frac{1}{q}\right\rceil!(\log 2)^{-\frac{1}{q}}\left\|W\right\|_{\psi_{q}},$
where in the last inequality, we used
$\left\|\mathbb{E}_{\mu}[W]\right\|_{\mathfrak{B}}\leq\mathbb{E}_{\mu}\big{[}\left\|W\right\|_{\mathfrak{B}}\big{]}\leq\left\lceil\frac{1}{q}\right\rceil!\left\|W\right\|_{\psi_{q}}.$ (103)
The first inequality here, in turn, follows from properties of Bochner integral and the second inequality is a consequence of the fact that for a non-negative valued random variable $X$,
[…]
C.2. Proof of Lemma 6
We will follow the proof given in [75, Lemma 2.2.2.], by avoiding arguments that require convexity of $\psi$. For $\psi_{q}(x)=e^{x^{q}}-1$, let $x^{\star}_{q}=\big{(}\log(3/2)\big{)}^{1/q}$. Observe that for $x,y\geq x^{\star}_{q}$ and $c_{q}^{\prime}\geq\big{(}2/\log(3/2)\big{)}^{1/q}$, we have $\psi_{q}(x)\psi_{q}(y)\leq\psi_{q}(c_{q}^{\prime}xy)$. Consequently, for any $y\geq x^{\star}_{q}$ and $x\geq x^{\star}_{q}y$,
$\psi_{q}\left(\frac{x}{y}\right)\leq\frac{\psi_{q}(c_{q}^{\prime}x)}{\psi_{q}(y)}.$ (104)
Then, for any $y\geq x^{\star}_{q}$ and $C>0$, we have
$\max_{i}\psi_{q}\left(\frac{\left|X_{i}\right|}{Cy}\right)$ $=\max_{i}\left(\psi_{q}\left(\frac{\left|X_{i}\right|}{Cy}\right)\mathbbm{1}_{\big{\{}\left|X_{i}\right|\geq Cyx^{\star}_{q}\big{\}}}+\psi_{q}\left(\frac{\left|X_{i}\right|}{Cy}\right)\mathbbm{1}_{\big{\{}\left|X_{i}\right|<Cyx^{\star}_{q}\big{\}}}\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\max_{i}\left(\frac{\psi_{q}\left(\frac{c_{q}^{\prime}\left|X_{i}\right|}{C}\right)}{\psi_{q}(y)}+\psi_{q}\left(\frac{\left|X_{i}\right|}{Cy}\right)\mathbbm{1}_{\big{\{}\left|X_{i}\right|<Cyx^{\star}_{q}\big{\}}}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\max_{i}\frac{\psi_{q}\left(\frac{c_{q}^{\prime}\left|X_{i}\right|}{C}\right)}{\psi_{q}(y)}+\psi_{q}(x^{\star}_{q})$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\sum_{i}^{n}\frac{\psi_{q}\left(\frac{c_{q}^{\prime}\left|X_{i}\right|}{C}\right)}{\psi_{q}(y)}+\frac{1}{2},$
where in $(a)$, we applied (104) with $x=\left|X_{i}\right|/C$ under the event that $\left|X_{i}\right|\geq Cyx^{\star}_{q}$, $(b)$ used that $\psi_{q}$ is monotone increasing, and in $(c)$ follows by upper bounding $\max$ by sum, and $\psi_{q}(x^{\star}_{q})=0.5$. Setting $C=c_{q}^{\prime}\max_{i}\left\|X_{i}\right\|_{\psi_{q}}$ and taking expectation, we obtain for $y\geq x^{\star}_{q}$
$\mathbb{E}\left[\psi_{q}\left(\frac{\max_{i}\left|X_{i}\right|}{c_{q}^{\prime}\max_{i}\left\|X_{i}\right\|_{\psi_{q}}y}\right)\right]\leq\frac{n}{\psi_{q}(y)}+\frac{1}{2}.$
Choosing $y=\psi_{q}^{-1}(2n)\geq x^{\star}_{q}$, the RHS is less than 1, thus implying that (by choosing $c_{q}^{\prime}=\big{(}2/\log(3/2)\big{)}^{1/q}\leq 5^{\frac{1}{q}}$)
$\left\|\max_{i}\left|X_{i}\right|\right\|_{\psi_{q}}\leq 5^{\frac{1}{q}}\psi_{q}^{-1}(2n)\max_{i}\left\|X_{i}\right\|_{\psi_{q}}\leq\big{(}5\log(1+2n)\big{)}^{\frac{1}{q}}\max_{i}\left\|X_{i}\right\|_{\psi_{q}}.$